### v1

# Ali Dasdan KD Consulting Saratoga, CA, USA

Abstract
In 1930, John Maynard Keynes predicted that technological progress would reduce the working week to 15 hours by century’s end. Despite productivity increasing eightfold since then, workers globally still work roughly double these hours. Separately, leading AI researchers warn of existential risks from artificial intelligence, yet development accelerates rather than slows. We demonstrate these failures share identical game-theoretic structure: coordination failures where individually rational choices produce collectively suboptimal outcomes.
We synthesize from established game theory five necessary and sufficient conditions that characterize such coordination failures as structural tragedies: N-player structure with multiple independent actors, binary choices generating negative externalities, dominance property where defection yields higher payoffs, Pareto-inefficiency where cooperation dominates mutual defection, and enforcement difficulty from structural barriers. We extend this framework to quantify tragedy severity through condition intensities, introducing a Tragedy Index that reveals AI governance faces orders-of-magnitude greater coordination difficulty than historical cases including climate change or nuclear weapons. We validate this framework across canonical cases including commons tragedies, security dilemmas, bank runs, and space debris.
Applied to productivity competition, we prove firms face a coordination failure preventing productivity gains from translating fully to worker welfare. We demonstrate this formally for working hours; the same structural logic extends to wages, benefits, and job security. European evidence demonstrates that even under maximally favorable conditions, such as wealthy democracies, strong institutions, and supranational governance, productivity welfare decoupling persists, manifesting differently across institutional contexts.
Applied to AI governance, we show development faces the same structure but with amplified intensity across eight dimensions compared to successful arms control cases, making coordination structurally more difficult than for nuclear weapons. The Russia-Ukraine drone war provides empirical validation: despite international dialogue on weapons governance, both sides escalated from zero to thousands of drones monthly within two years, driven by competitive necessity that overwhelmed coordination aspirations.
The analysis is diagnostic rather than prescriptive, identifying structural barriers to coordination rather than proposing solutions whose feasibility our framework questions.
1 Introduction
1.1 The Productivity Puzzle
John Maynard Keynes famously predicted in 1930 that technological progress would reduce the working week to 15 hours by century’s end (Keynes, 1930). Nearly a century later, this prediction appears dramatically wrong. Workers globally typically work 1,350-2,200+ hours annually (OECD, 2023), far exceeding Keynes’s predicted 780 hours, and in many sectors working hours have increased rather than decreased (Schor, 1992); this competitive escalation is exemplified by the “996” work culture (9 am to 9 pm, 6 days a week) in China (Wikipedia contributors, 2025), which has even been emulated by some US startups seeking to match that intensity (The New York Times, 2025).
This failure demands explanation. The technological capacity for reduced working hours clearly exists: productivity per worker-hour has increased approximately eightfold since Keynes wrote (Gordon, 2016). The material abundance Keynes envisioned has been achieved: modern economies produce far more goods and services per hour of labor than was possible in 1930. Yet this capacity for leisure remains unrealized.
Why? Conventional explanations typically invoke cultural factors or policy failures. We argue these explanations are incomplete. The persistence of long working hours despite dramatic productivity gains reflects not mere policy failure or cultural preferences, but a structural coordination problem analogous to tragedies in other domains.
This study is also motivated by a persistent empirical pattern observed by the author in the software and technology sectors over the last two decades. In these industries, leadership consistently demands output expansion following productivity gains, often reducing engineering headcount rather than stabilizing it. This ’do more with less’ dynamic persists even when cost-cutting threatens organizational sustainability.
1.2 The AI Governance Puzzle
A parallel puzzle emerges in artificial intelligence development. Leading AI researchers consistently warn of existential risks from advanced AI systems (Bostrom, 2014; Russell et al., 2015; Amodei et al., 2016). Survey evidence shows substantial minorities of AI researchers estimate non-trivial probabilities of extremely bad outcomes, including human extinction (Bengio et al., 2025; Grace et al., 2018). Major AI laboratories publicly commit to safety and responsible development (Brundage et al., 2018).
Yet development accelerates rather than slows. Investment in AI capabilities races ahead (Federal Reserve Board, 2025). Voluntary pauses fail to materialize. Safety measures, while present, appear insufficient relative to the risks researchers identify (Askell et al., 2019). International coordination remains nascent and non-binding.
Why can’t we coordinate even when the stakes appear existential? Why do warnings from the field’s leading researchers fail to slow development? Why do commitments to safety coexist with accelerating capabilities races?
1.3 The Thesis: Structural Identity
We argue these puzzles exhibit the same mathematical structure. Both represent structural tragedies, coordination failures where individually rational choices produce collectively suboptimal outcomes and where coordination faces overwhelming structural barriers.
This structural identity has important implications. First, it explains why standard solutions fail in both domains: the barriers are structural, not merely political or cultural. Second, it enables prediction: we can anticipate which coordination attempts will succeed or fail based on whether they address the underlying structure. Third, it clarifies what kinds of interventions might succeed: only those that fundamentally alter the competitive architecture itself.
1.4 Preview of Contributions
This paper makes three contributions:
First, we synthesize a unified framework from established game theory (Dawes, 1980; Jervis, 1978) identifying five necessary and sufficient conditions that create structural tragedies. These conditions provide a diagnostic tool applicable across domains. We validate this framework across canonical cases including environmental commons (Hardin, 1968), security dilemmas (Mearsheimer, 2001), bank runs (Diamond and Dybvig, 1983), and other established coordination failures.
Second, we apply this framework to productivity competition and show that it satisfies all five conditions, explaining Keynes’s failed prediction structurally rather than culturally. We formalize the coordination problem firms face when productivity increases, showing why welfare improvements remain difficult despite technological capacity. European evidence demonstrates that even under maximally favorable conditions, such as wealthy democracies, strong institutions, and supranational governance, coordination achieves only partial, costly overrides of the tragic logic that erode over time.
Third, we demonstrate that AI development exhibits the same structure but with amplified intensity. Systematic comparison with historical arms control reveals that AI faces unfavorable conditions on every dimension where nuclear, biological, chemical (NBC) weapons coordination succeeded: dual-use inseparability, verification impossibility, low entry barriers, absence of substitutes, and convergence of economic and military imperatives. The Russia-Ukraine drone war provides empirical validation, showing how rapidly competitive dynamics override governance intentions when existential stakes are involved.
1.5 Organization
The remainder of this paper is organized as follows. Section 2 reviews the related literature on social dilemmas, productivity, and AI safety. Section 3 introduces our unified framework, defining the five necessary conditions for structural tragedy and the Tragedy Index. Section 4 applies this framework to labor markets, providing a formal proof of the coordination failure in working hours. Section 5 extends the analysis to AI governance, offering a comparative analysis with historical arms control and examining the Russia-Ukraine drone conflict. Section 6 discusses testable predictions, falsifiability, and potential interventions. Section 7 concludes.
2 Related Literature
Our work synthesizes insights from multiple, often separate, literatures.
2.1 The Prisoner’s Dilemma and Social Dilemmas
The prisoner’s dilemma, introduced by Tucker and formalized by Rapoport and Chammah (1965), represents the canonical coordination failure. Two players each face a choice between cooperation and defection. Defection dominates cooperation regardless of the other’s choice, yet mutual cooperation Pareto-dominates mutual defection. The unique Nash equilibrium—mutual defection—is recognizably suboptimal.
Schelling (1960, 1973) extended this analysis to N-player settings with binary choices and externalities. When multiple actors face similar decisions where individual choices impose externalities on others, coordination failures can emerge even when all actors understand the collective problem. The key insight is that as $N$ increases, the individual incentive to deviate grows stronger—each actor’s restraint provides smaller individual benefit while the temptation to defect remains constant.
Dawes (1980) formalized the structure of “social dilemmas,” defining them as N-person games where individual rationality leads to collective irrationality. His key contribution was identifying the mathematical structure common to diverse coordination failures: defection dominates cooperation for each individual, yet universal cooperation yields higher payoffs than universal defection. Dawes emphasized that social dilemmas persist even when actors fully understand the collective harm, because rational individual choice remains incompatible with collective welfare absent enforcement mechanisms.
Jervis (1978) applied this logic to international relations, analyzing the “security dilemma” where states’ efforts to increase their own security—through arms buildups or territorial expansion—threaten other states and trigger countermeasures that leave all parties less secure. The security dilemma exhibits the core social dilemma structure: each state has incentive to arm regardless of others’ choices (dominance), yet mutual restraint would benefit all (Pareto-efficiency), but international anarchy prevents enforceable agreements. Jervis demonstrated that this tragic dynamic persists even when all states prefer peace, because the anarchic structure of international relations makes credible commitment to cooperation impossible.
The folk theorem literature (Fudenberg and Maskin, 1986) demonstrates that repeated interaction can sustain cooperation under certain conditions through trigger strategies. However, these results require strong assumptions: players must be sufficiently patient, the game must repeat infinitely, and actions must be perfectly observable. When these conditions fail—as we show they do in productivity competition and AI development—cooperation becomes unsustainable even with repetition.
This foundational literature, particularly the work of Dawes (1980) on social dilemmas and Jervis (1978) on the security dilemma, establishes the core game-theoretic structure that this paper adopts. We synthesize these insights into a diagnostic framework specifying exactly which five conditions create coordination failures that resist standard solutions including repeated interaction, reputation, and institutional design. This enables systematic comparison across domains and reveals mathematical identity where previous work identified domain-specific phenomena.
2.2 Tragedies in Specific Domains
Multiple literatures have identified coordination failures in specific domains.
Hardin (1968) demonstrated how individual rationality leads to resource depletion when property rights are absent. Ostrom (1990) showed that under certain conditions—small groups, clear boundaries, monitoring capacity, graduated sanctions—communities can overcome commons tragedies through institutional design. However, her analysis also revealed that many commons lack these favorable conditions, making coordination structurally difficult.
Jervis (1978) formalized the security dilemma where states’ defensive preparations threaten other states, generating arms races that make all parties less secure. Mearsheimer (2001) extended this into offensive realism, arguing that international anarchy creates tragic dynamics where security competition becomes structurally difficult to avoid even when all states prefer peace. The key insight is that uncertainty about intentions, combined with the anarchic structure of international relations, makes cooperation on security matters extremely difficult.
Olson (1965) analyzed collective action problems in public goods provision, showing why rational individuals undercontribute to group benefits. Diamond and Dybvig (1983) modeled bank runs as coordination failures where individual rationality leads to systemic crisis. These analyses demonstrate that coordination failures pervade economic life, not merely environmental or security domains.
Our framework makes the underlying structure explicit, treating these as variations of the same underlying social dilemma.
2.3 Productivity and Working Hours
The literature on working hours exhibits puzzlement about why productivity gains don’t translate into leisure (Keynes, 1930).
Historical research documents the trajectory of working hours. Huberman and Minns (2007) show that annual working hours in industrialized economies declined from over 3,000 hours in 1870 to approximately 1,800 hours by 2000, but this decline has essentially stalled since the 1970s despite continued productivity growth. Pencavel (2018) documents the relationship between productivity and hours, finding that after initial declines, working hours have remained stubbornly stable.
Theoretical explanations focus primarily on preferences and culture. Becker (1965) developed time allocation theory showing how rising wages could increase or decrease labor supply depending on income and substitution effects. Schnaiberg (1980) introduced the “treadmill of production” concept, arguing that competitive pressures force continuous expansion. Alesina et al. (2005) examine US-Europe differences in working hours and conclude that European labor market regulations, advocated by unions in declining industries who argued “work less, work all,” explain the bulk of the difference, rather than cultural preferences or tax structures alone.
Recent work by Acemoglu and Restrepo (2020) and Acemoglu (2025) provides important insights into technology adoption, distinguishing between automation that displaces labor (“so-so automation”) and technology that creates new tasks or augments workers (“human-augmenting”). They argue that market forces may lead to socially excessive automation due to insufficient incentives for firms to internalize benefits that accrue to workers.
While Schnaiberg (1980) intuited competitive pressure as a “treadmill,” and Acemoglu and Restrepo (2020); Acemoglu (2025) identified market failures in technology choices, no formal game-theoretic model explains the coordination problem firms face when productivity increases. Also while extensive research documents the secular decline in the labor share of income globally, (Karabarbounis and Neiman, 2014; Elsby et al., 2013), the literature lacks a unified game-theoretic explanation for why competitive pressure systematically prevents productivity gains from translating to worker welfare.
We provide the first formal game-theoretic model of productivity competition as an N-player social dilemma. Building on Schnaiberg (1980)’s treadmill intuition, we prove that firms face a structural coordination failure: each firm has a dominant strategy to expand output rather than reduce hours, yet universal restraint would Pareto-dominate universal expansion. We formalize the barriers to coordination that make this tragedy persist. This structural framework explains Schnaiberg (1980)’s empirical observation, integrates with Acemoglu and Restrepo (2020); Acemoglu (2025)’s analysis of technology choice failures, and clarifies why the regulatory interventions documented by Alesina et al. (2005) require costly, persistent institutional support to partially override market dynamics.
2.4 AI Governance and Safety
The AI safety literature has grown rapidly in recent years, identifying multiple challenges in developing safe and beneficial AI systems.
Bostrom (2014) provided a comprehensive analysis of existential risks from superintelligent AI, emphasizing the orthogonality thesis (intelligence and goals are independent) and instrumental convergence (advanced AI systems will pursue convergent subgoals regardless of final objectives). Russell (2019) framed AI safety as an inverse reinforcement learning problem and proposed provably beneficial AI through value alignment. Amodei et al. (2016) identified concrete problems in AI safety: avoiding negative side effects, reward hacking, distributional shift, safe exploration, and robustness to distributional change. Dafoe (2018) analyzed AI governance challenges, identifying the need for international coordination on development and deployment. Bengio et al. (2025) provided a comprehensive synthesis of the current evidence on the capabilities, risks, and safety of advanced AI systems. Panigrahy and Sharan (2025) proved that under reasonable mathematical definitions of safety, trust, and artificial general intelligence (AGI), a safe and trusted AI system cannot be an AGI system.
Armstrong et al. (2016) analyzed “racing to the precipice”, the concern that competitive pressure could drive unsafe AI development. Their analysis identifies the race dynamic but does not provide a formal structural model explaining why coordination fails.
While this literature thoroughly documents AI risks and governance challenges, it lacks a formal structural analysis of why coordination fails. What are the game-theoretic conditions that make AI governance difficult? Why do commitments to safety coexist with accelerating development? How does AI compare structurally to previous coordination challenges like NBC weapons?
3 A Unified Framework for Structural Tragedies
We now present a unified framework identifying necessary and sufficient conditions for structural tragedies. We formalize these conditions, prove they are necessary and sufficient to generate coordination failures, and validate the framework across canonical cases.
We formalize the payoff structure as follows. Let $s=(s_{1},s_{2},\ldots,s_{n})$ be a strategy profile, which is a vector specifying the chosen strategy $s_{i}\in\{C,D\}$ for every player $i$.
The utility (or payoff) for a specific player $i$ is given by a function $U_{i}$ that maps this complete strategy profile to a real-valued payoff:
$U_{i}(s)=U_{i}(s_{1},s_{2},\ldots,s_{n})$
For clarity, we will often separate the strategy of player $i$ from the strategies of all other players. We denote the strategy profile of all players except $i$ as $s_{-i}$. This allows us to write the utility function as $U_{i}(s_{i},s_{-i})$.
Using this formal notation, we can precisely define the five conditions for the tragedy:
3.1 The Five Conditions
A problem exhibits structural tragedy if and only if it satisfies five conditions:
Definition 1 (C1: N-Player Structure).
Multiple independent decision-makers ($n\geq 2$) each control their own choices. No single actor can solve the problem unilaterally. Choices are made in a decentralized architecture without centralized coordination.
This condition specifies the basic game structure. If a single actor controlled all decisions, coordination would be trivial. The tragedy emerges from decentralized choice architecture where multiple actors must independently decide their strategies.
Definition 2 (C2: Binary Choice with Externalities).
Each actor $i$ faces a binary choice between two strategies: Cooperate ($C$) representing restraint, or Defect ($D$) representing advantage-seeking. Individual choices create externalities affecting all actors. Specifically, one actor’s defection imposes negative externalities on others.
The binary structure simplifies the game while capturing the essential strategic tension. In practice, choices may appear continuous, but the strategic logic often reduces to a binary decision: restrain (cooperate) or expand (defect). The key is that individual choices impose externalities—one actor’s defection makes cooperation more costly for others.
Definition 3 (C3: Dominance Property).
Defection yields higher payoff than cooperation regardless of other actors’ choices. Formally,
$U_{i}(D,s_{-i})>U_{i}(C,s_{-i})\quad\text{for any possible }s_{-i}$
Dominance creates the dilemma. Each actor has strict incentive to defect regardless of what others do. This is often driven by existential stakes. In high-stakes tragedies, survival instincts override social norms; if cooperating while others defect leads to elimination (e.g., death, bankruptcy), defection becomes an absolute imperative. When survival is at stake, the incentive to defect becomes overwhelming.
Definition 4 (C4: Pareto-Inefficiency).
Universal cooperation Pareto-dominates universal defection. Formally, let $s^{C}=(C,C,\ldots,C)$ be the Universal Cooperation profile and let $s^{D}=(D,D,\ldots,D)$ be the Universal Defection profile. This condition states that every player strictly prefers the universal cooperation outcome:
$U_{i}(s^{C})>U_{i}(s^{D})\quad\text{for all players }i.$
The equilibrium outcome is recognizably suboptimal—all actors would prefer mutual cooperation to mutual defection if they could coordinate.
This condition ensures the tragedy is recognizable. Actors understand that mutual defection makes everyone worse off than mutual cooperation. The tragedy is not hidden—it is visible, predictable, and collectively unwanted. Yet the structure makes it difficult to escape.
Definition 5 (C5: Enforcement Difficulty).
Structural barriers prevent effective coordination mechanisms. These barriers may include: (i) anarchy as in no central authority with enforcement power, (ii) verification impossibility as in actions cannot be reliably monitored, (iii) monitoring infeasibility as in costs of monitoring exceed benefits of coordination, or (iv) commitment problems as in actors cannot credibly commit to future cooperation.
This condition distinguishes structural tragedies from simple coordination problems. Many Pareto-inefficient equilibria can be overcome through contracts, monitoring, or institutional design. Structural tragedies persist because enforcement mechanisms face fundamental barriers—not merely high costs but structural impossibilities.
Throughout the paper, we will refer to the conditions using the condition number and the first letter of the condition name as in “C1:N” or “C5:E”.
3.2 The Formal Result
We now prove these conditions are both necessary and sufficient to generate structural tragedy.
Theorem 6 (Necessity and Sufficiency of Tragedy Conditions).
Consider an N-player game satisfying Conditions 1-2 (structure and externalities). Then:
  1. (i)
    Universal defection $s^{D}$ is the unique Nash equilibrium.
  2. (ii)
    This equilibrium is stable against unilateral deviations.
  3. (iii)
    This equilibrium is Pareto-dominated by universal cooperation $s^{C}$.
  4. (iv)
    This equilibrium persists absent radical structural change.
if and only if Conditions 3-5 (dominance, Pareto-inefficiency, and enforcement difficulty) also hold.
Proof.
We establish sufficiency and necessity separately.
Part I: Sufficiency.
Assume Conditions 1-5 hold. We prove claims (i)-(iv).
Claim (i): Unique Nash Equilibrium
Let $s=(s_{1},\ldots,s_{n})$ be a strategy profile where $s_{i}\in\{C,D\}$ for each actor $i$.
First, we show $s^{D}=(D,\ldots,D)$ is a Nash equilibrium. For any actor $i$, their payoff from defecting is $U_{i}(D,s^{D}_{-i})$, where $s^{D}_{-i}$ is the profile where all other players also defect. If actor $i$ unilaterally deviates to cooperation, their payoff becomes $U_{i}(C,s^{D}_{-i})$. By the dominance property (Condition 3), $U_{i}(D,s^{D}_{-i})>U_{i}(C,s^{D}_{-i})$. Therefore, no actor has an incentive to deviate, and $s^{D}$ is a Nash equilibrium.
Second, we show no other profile is a Nash equilibrium. Consider any profile $s$ where at least one actor $j$ cooperates. This actor’s payoff is $U_{j}(C,s_{-j})$, where $s_{-j}$ is the profile of all other players. If actor $j$ deviates to defection, their payoff becomes $U_{j}(D,s_{-j})$. By the dominance property (Condition 3), $U_{j}(D,s_{-j})>U_{j}(C,s_{-j})$. Therefore, actor $j$ has a strict incentive to deviate. Hence, no profile with any cooperation is a Nash equilibrium.
Therefore, $s^{D}$ is the unique Nash equilibrium.
Claim (ii): Stability
Stability follows immediately from Claim (i). At the equilibrium $s^{D}$, no actor has unilateral incentive to deviate by the dominance property. The equilibrium is thus stable against unilateral deviations.
Claim (iii): Pareto-Domination
By Condition 4 (Pareto-inefficiency), $U_{i}(s^{C})>U_{i}(s^{D})$ for all actors $i$. Therefore, universal cooperation strictly Pareto-dominates the Nash equilibrium.
Claim (iv): Persistence
Even though all actors recognize that $s^{C}$ Pareto-dominates $s^{D}$, they cannot coordinate to achieve it. By Condition 5, structural barriers prevent enforcement:
  • If anarchy prevails, no authority can compel cooperation
  • If verification is impossible, actors cannot monitor compliance
  • If monitoring is infeasible, detection of defection comes too late
  • If commitment is problematic, promises to cooperate lack credibility
Without enforcement, any agreement to cooperate is unstable, as each actor has a strict incentive to defect (from Claim (i)). The folk theorem, which allows for cooperation in repeated games, does not apply because its required conditions (perfect monitoring, sufficient patience, infinite repetition) fail.
Therefore the tragedy persists absent radical structural change.
Part II: Necessity.
We prove necessity by showing the contrapositive: if any of Conditions 3-5 fails to hold, then the tragedy structure breaks (at least one of claims (i)-(iv) fails). By contrapositive logic, this establishes that each condition is necessary for tragedy.
[…]
  • A central authority exists with enforcement power, or
  • Actions are perfectly observable and verifiable, or
  • Monitoring is feasible at reasonable cost, or
  • Actors can make credible commitments to future cooperation
Under these conditions, actors can contract to implement $s^{C}$ and sustain it through enforcement mechanisms. Defection can be detected and punished. Commitments become credible. The folk theorem for repeated games may apply: if actors are sufficiently patient and monitoring is adequate, cooperation can be sustained as a subgame perfect equilibrium through trigger strategies that punish defection.
While cooperation may require institutional infrastructure, it becomes achievable through standard mechanisms—contracts, monitoring, reputation, repeated interaction. The tragedy dissolves from structurally impossible to merely institutionally challenging. Therefore C5 is necessary.
We have shown that removing any single condition eliminates the structural tragedy (contrapositive of necessity). Combined with Part I showing sufficiency, we conclude that Conditions 1-5 are necessary and sufficient for structural tragedy as characterized by Theorem 6.
This theorem establishes that Conditions 1-5 precisely characterize structural tragedies. The equilibrium is not merely Pareto-inefficient: it is recognizably suboptimal yet structurally difficult to escape. The framework provides both diagnostic power (checking conditions predicts tragedy) and intervention guidance (breaking any condition resolves tragedy).
3.3 Validation Across Canonical Cases
We now validate this framework by applying it to established coordination failures across multiple domains. Table 1 summarizes the analysis.
Table 1: Five Conditions Across Canonical Coordination Failures, including the two new ones introduced in this paper.
Domain C1:N C2:B C3:D C4:P C5:E Reference
Commons Hardin (1968)
Security Jervis (1978); Mearsheimer (2001)
Bank Runs Diamond and Dybvig (1983)
Space Debris Adilov et al. (2015)
Antibiotics Foster and Grundmann (2006)
Public Goods Olson (1965)
Hockey Helmets Schelling (1973)
Climate Change Ostrom (2010)
Productivity this paper
AI Governance this paper
We briefly discuss each case, demonstrating how it satisfies all five conditions.
Commons Tragedies (Hardin, 1968): Multiple shepherds (N-player) each decide whether to add sheep to common pasture (binary choice creating negative externality through overgrazing). Each shepherd benefits from adding sheep regardless of others’ choices (dominance driven by individual profit maximization). Yet universal restraint would preserve the commons for all (Pareto-domination). Absent property rights or enforceable agreements (enforcement difficulty), the commons is depleted.
Security Dilemmas (Jervis, 1978; Mearsheimer, 2001): Multiple states (N-player) each decide whether to arm or remain unarmed (binary choice where one state’s arms create security threats for others). Each state benefits from arming for security regardless of others’ choices (dominance driven by survival concerns in anarchy). Yet universal disarmament would provide security without arms costs (Pareto-domination). International anarchy prevents enforceable disarmament agreements (enforcement difficulty).
Bank Runs (Diamond and Dybvig, 1983): Multiple depositors (N-player) each decide whether to withdraw or keep deposits in a bank (binary choice where withdrawals create panic). Each depositor benefits from withdrawing if others might withdraw (dominance driven by fear of losing deposits). Yet if no one panics, all maintain deposits safely (Pareto-domination). Absence of instant coordination during panics (enforcement difficulty) allows runs to occur. Deposit insurance (FDIC) broke this tragedy (for deposits up to the insurance limit) by eliminating existential stakes, removing Condition 3. Note that the wholesale funding and equities markets are not covered under the FDIC insurance so they may still be vulnerable to bank runs.
Space Debris (Adilov et al., 2015): Multiple states and firms (N-player) each decide whether to limit launches or expand operations (binary choice where debris from launches threatens all satellites). Each actor benefits from launches regardless of others’ choices (dominance driven by commercial and security advantages). Yet universal restraint would preserve orbital space (Pareto-domination). International space law lacks enforcement mechanisms (enforcement difficulty).
Antibiotic Resistance (Foster and Grundmann, 2006): Multiple actors including doctors, patients, farmers (N-player) each decide whether to limit or freely use antibiotics (binary choice where overuse breeds resistance affecting all). Each actor benefits from use regardless of others’ choices (dominance driven by immediate health or productivity gains). Yet universal restraint would preserve antibiotic effectiveness (Pareto-domination). Distributed usage across multiple sectors resists coordination (enforcement difficulty).
Public Goods (Olson, 1965): Multiple individuals (N-player) each decide whether to contribute to a public good (Cooperate) or withhold contribution (Defect). Withholding allows an individual to save resources (free-ride) while still enjoying any good provided by others (externality). Each individual has a dominant incentive to free-ride, as their personal contribution is costly and has minimal impact on the total outcome (dominance). Yet, universal contribution would fund the public good, making everyone better off than if the good is not provided (Pareto-domination). Absent a coercive mechanism like taxation (enforcement difficulty), the public good is under-provided or not provided at all.
Hockey Helmets (Schelling, 1973): Multiple players (N-player) each decide whether to wear a helmet (Cooperate/Restrain) or play without one (Defect/Advantage-seeking). Playing without a helmet provides a slight competitive advantage (e.g., better vision, “toughness”) but increases the overall risk of injury for all (negative externality). Each player benefits from the competitive edge of not wearing one, regardless of others’ choices (dominance). Yet, universal helmet use would neutralize the competitive effects and make everyone safer (Pareto-domination). Absent a league-wide, binding mandate (enforcement difficulty), players are incentivized to play without helmets despite the collective danger.
Climate Change (Ostrom, 2010): Multiple nations and firms (N-player) each decide whether to reduce emissions (Cooperate) or pollute freely (Defect). Polluting provides immediate economic advantages while imposing global climate damage (negative externality). Each actor benefits from polluting to maintain economic competitiveness, regardless of others’ choices (dominance). Yet, universal restraint would prevent catastrophic climate costs, making all better off (Pareto-domination). International anarchy and sovereignty concerns prevent binding, enforceable agreements (enforcement difficulty).
Result: Perfect verification across all canonical cases. Each tragedy satisfies all five conditions. This suggests the framework captures the essential structure that creates coordination failures.
Notably, cases that do not satisfy all conditions either are not tragedies or have been successfully resolved. For example, bank runs violated Condition 3 (dominance) when deposit insurance eliminated existential stakes (up to the insurance limit) since depositors no longer benefit from withdrawing if others withdraw. This structural change broke the tragedy, validating our framework’s diagnostic power.
3.4 Implications of the Framework
This unified framework provides several analytical advantages:
Identification: We can identify new tragedies by checking whether a situation satisfies all five conditions. If it does, coordination will be structurally difficult. If any condition fails, standard solutions may suffice.
Prediction: We can predict difficulty of coordination by examining the intensity with which each condition is satisfied. Situations with extreme dominance (high existential stakes) or severe enforcement problems will resist coordination more strongly than situations where conditions are weakly satisfied.
Intervention Design: We can design interventions by targeting specific conditions. Breaking any single condition may resolve the tragedy.
3.5 Extension: Condition Intensity and Tragedy Severity
While the framework treats conditions as binary (present or absent) for analytical clarity, each condition varies in intensity in real-world applications. The severity of a structural tragedy increases with the intensity of each condition. This intensity dimension provides additional predictive and diagnostic power.
C1: N-Player Structure - Number and Heterogeneity of Actors.
Intensity spectrum:
  • Low intensity: Few actors (e.g., n=3n=3-5 firms in an oligopoly)
  • Medium intensity: Moderate number of actors (e.g., n=20n=20-50)
  • High intensity: Large number of actors (e.g., thousands of firms, hundreds of nations)
Impact on tragedy severity: More actors make coordination exponentially more difficult. Each actor’s individual contribution to the collective good diminishes as $n$ increases, while the temptation to defect remains constant. Additionally, heterogeneity among actors (different sizes, capabilities, preferences) further intensifies coordination difficulty.
Examples:
  • Low: OPEC coordination among  13 major oil producers—difficult but achievable
  • High: Climate change with 195+ nations plus countless firms—extremely difficult
C2: Externalities - Magnitude and Scope.
Intensity spectrum:
  • Low intensity: Small negative externalities affecting few others
  • Medium intensity: Moderate externalities with delayed effects
  • High intensity: Large, immediate externalities affecting all actors severely
Impact on tragedy severity: Larger externalities create stronger feedback loops. When one actor’s defection severely harms others, it triggers more aggressive competitive responses, accelerating the race to the bottom. The speed and magnitude of externality transmission determines how rapidly the tragedy unfolds.
Examples:
  • Low: Overfishing in a large ocean—gradual depletion over decades
  • High: Bank runs—panic spreads within hours, complete collapse within days
  • High: AI capabilities races—competitor’s breakthrough immediately threatens all others
C3: Dominance - Strength of Existential Stakes.
Intensity spectrum:
  • Low intensity: Defection provides modest advantage; cooperation causes minor disadvantage
  • Medium intensity: Defection provides significant advantage; cooperation causes substantial losses
  • High intensity: Defection is existentially necessary; cooperation risks elimination (bankruptcy, conquest, extinction)
Impact on tragedy severity: The intensity of existential stakes determines how strongly dominance holds. When cooperation means death or bankruptcy, actors will defect regardless of collective costs. Lower existential stakes create space for partial coordination or voluntary restraint.
Examples:
  • Low: Hockey helmets—competitive disadvantage is marginal visibility/comfort
  • Medium: Productivity competition—firms face market share loss, not immediate bankruptcy
  • High: Security dilemmas in anarchy—states face conquest and political death
  • Extreme: AI development with AGI—actors face permanent strategic obsolescence or existential risk
C4: Pareto-Inefficiency - Magnitude of Collective Loss.
Intensity spectrum:
  • Low intensity: Universal cooperation slightly better than universal defection
  • Medium intensity: Universal cooperation moderately better—significant welfare gains available
  • High intensity: Universal cooperation vastly superior—catastrophic outcomes from mutual defection
Impact on tragedy severity: Greater Pareto-inefficiency means larger collective losses from coordination failure. This affects both the normative urgency (how important it is to solve) and practical difficulty (actors may invest more in solving severe tragedies). However, high inefficiency alone doesn’t make coordination easier—it only makes failure more costly.
Examples:
  • Low: Minor commons problems—modest resource depletion
  • Medium: Productivity competition—substantial leisure forgone, wage stagnation
  • High: Climate change—potential civilizational collapse
  • Extreme: AI misalignment—potential human extinction
C5: Enforcement Difficulty - Severity of Structural Barriers.
Intensity spectrum:
  • Low intensity: Enforcement difficult but feasible; monitoring possible with effort and cost
  • Medium intensity: Enforcement very difficult; verification problematic; commitment fragile
  • High intensity: Enforcement structurally impossible; verification impossible; no credible commitment mechanism exists
Impact on tragedy severity: The impossibility of enforcement determines whether the tragedy is merely difficult or truly structural. When enforcement is difficult but possible (low intensity), sufficient institutional investment can achieve coordination. When enforcement is structurally impossible (high intensity), no amount of effort suffices without changing the underlying structure.
Subcategories of enforcement difficulty:
  • Anarchy intensity: No authority $\rightarrow$ weak authority $\rightarrow$ strong authority
  • Verification impossibility: Perfect hiding $\rightarrow$ difficult detection $\rightarrow$ easy monitoring
  • Commitment problems: No credible commitment $\rightarrow$ costly commitment $\rightarrow$ binding commitment
Examples:
  • Low: National labor regulations—government can monitor and enforce within borders
  • Medium: Nuclear non-proliferation—verification possible but imperfect through inspections
  • High: International labor standards—no global authority, capital mobility enables evasion
  • Extreme: AI development—software is copyable, training is concealable, dual-use is inseparable
3.6 Implications of Intensity Analysis
This intensity framework provides several analytical advantages:
Comparative prediction: We can predict relative difficulty of coordination across domains by comparing condition intensities. AI governance faces high intensity on all five conditions simultaneously, while nuclear non-proliferation faced high intensity on only some conditions (dominance, Pareto-inefficiency) but lower intensity on others (entry barriers created enforcement possibility).
Intervention prioritization: Interventions should target the highest-intensity conditions. For AI, the extreme intensity of C5 (enforcement impossibility due to verification) suggests compute governance (making verification partially possible) may be most promising despite remaining severe challenges.
Partial coordination explanation: The European case shows lower intensity on some conditions (reduced existential stakes through wealthy welfare states, supranational governance reducing anarchy) enables partial coordination despite high intensity on others (global competition, verification difficulty).
Temporal variation explanation: Condition intensities change over time. Productivity competition intensified after 1980 as globalization increased N (more actors), strengthened externalities (faster competitive transmission), and weakened enforcement (capital mobility). This explains why coordination that worked 1945-1980 eroded subsequently.
Falsification refinement: The framework predicts that reducing condition intensity should reduce tragedy severity proportionally. Evidence of coordination success without reducing any condition’s intensity would falsify the framework. The European case confirms: partial reduction in intensity (some conditions) produces partial coordination (some dimensions).
3.7 Tragedy Index: Intensity Quantification
The intensity framework adds a continuous dimension to our binary framework, enabling more nuanced predictions while maintaining the core insight that all five conditions must be satisfied (at non-zero intensity) for structural tragedy to occur.
The intensity framework also allows for two analytical advances:
  • Structural Fingerprinting: By plotting intensities on a radar chart, we can visualize the ’shape’ of the coordination failure, revealing which specific conditions are the hardest to break for a given domain.
  • The Tragedy Index: We propose a quantitative metric to compare severity across domains, acknowledging that coordination difficulty plausibly scales non-linearly with condition intensity.
We propose the Tragedy Index ($T$) as a heuristic quantification of coordination difficulty. Two methodological choices deserve acknowledgment:
Exponential scaling: We use base-10 exponential intensities ($I_{c}\in\{0,10,10^{2},10^{3},10^{4}\}$) because coordination difficulty plausibly scales non-linearly: a tragedy with “extreme” conditions on multiple dimensions may be orders of magnitude harder than one with “low” conditions, not merely several times harder. The base-10 choice is somewhat arbitrary but enables intuitive order-of-magnitude interpretation.
Geometric aggregation: We combine intensities via geometric mean (the 5th root of the product):
$T=\left(\prod_{c=1}^{5}I_{c}\right)^{\frac{1}{5}}$
where $I_{c}$ corresponds to Solved/Absent (0), Low (10), Medium (100), High (1,000), or Extreme (10,000). This multiplicative formulation handles the zero case naturally (solved tragedies yield $T=0$) and penalizes extreme conditions heavily, reflecting their disproportionate contribution to coordination difficulty. Alternative aggregations such as logarithmic sum would yield different absolute values but similar ordinal rankings.
The value of this index lies not in precise quantification but in systematic comparison. We interpret the resulting values as “Tragedy Magnitude”:
  • Magnitude 0: Solved (Stable)
  • Magnitude 10–100: Manageable Nuisance (Chronic)
  • Magnitude 100–1,000: Systemic Crisis (Acute)
  • Magnitude 1,000+: Existential/Terminal Threat
Under this heuristic measure, AI Governance (T=3,981T=3{,}981) faces orders-of-magnitude greater coordination difficulty than climate change ($T=631$) or productivity competition ($T=631$), which themselves substantially exceed manageable cases like public goods ($T=40$). This quantification reinforces the qualitative analysis: AI governance presents an unprecedented coordination challenge.
Table 2 compares condition intensities across our main cases.
Table 2: Condition Intensity Comparison Across Domains, including the two new ones introduced in this paper. Intensities: L=Low (10), M=Medium (100), H=High (1,000), E=Extreme (10,000).
Domain C1:N C2:B C3:D C4:P C5:E Index ($T$)
Commons M M M M M 100
Security L H H H H 398
Bank Runs H H H M H 631
Space Debris M M M M H 158
Antibiotics H M M M H 251
Public Goods M L L M M 40
Hockey Helmets L L L L M 16
Climate Change H H M H H 631
Productivity H H H M H 631
AI Governance H H E E E 3,981
We now apply the basic framework (without the intensity extension) to productivity competition and AI development, demonstrating that both satisfy all five conditions, and that both are structural tragedies.
4 The Tragedy of Productivity
We now apply the unified framework to productivity competition, demonstrating that all five conditions are satisfied. This explains structurally why Keynes’s prediction failed: not from cultural preferences or policy failures, but from coordination barriers inherent in competitive markets.
Although worker welfare includes wages and benefits in addition to working hours, we will use working hours as a proxy for worker welfare as well as to simplify the formal analysis.
4.1 Mapping the Five Conditions to Firm Competition
C1: N-Player Structure.
Productivity competition involves multiple independent decision-makers:
  • Multiple firms compete in domestic and global markets
  • National governments face political pressure from firms and workers
  • No single actor can coordinate welfare improvements unilaterally
Even large firms cannot unilaterally reduce hours without risking competitive disadvantage. Even powerful governments face capital flight and reduced tax revenue if they impose work reductions that competitors don’t match. The decision architecture is fundamentally decentralized.
C2: Binary Choice with Externalities.
When productivity increases, each firm faces essentially a binary choice:
  • Strategy $C$ (Cooperate/Restrain): Reduce working hours proportionally to productivity gains while maintaining output
  • Strategy $D$ (Defect/Expand): Maintain or increase working hours to expand output and market share
Individual choices create externalities. When one firm expands output (defects), it increases competitive pressure on all rivals:
  • Expanded output captures market share from competitors
  • Increased competition reduces rivals’ profit margins
  • Rivals face pressure to expand output to avoid displacement
The externality is negative: one firm’s defection harms others by intensifying competition.
C3: Dominance Property.
Defection (output expansion) dominates cooperation (welfare improvements) regardless of competitors’ choices:
Case 1—If competitors reduce hours: Firm gains decisive market share advantage by maintaining output. Competitors who reduce hours produce less, creating market opportunity. The expanding firm captures this market share, increasing profits and market position.
Case 2—If competitors expand output: Firm must expand to avoid displacement. Competitors’ expansion increases supply and reduces prices. Reducing hours while competitors expand means producing less in a more competitive market—a recipe for bankruptcy.
This is often driven by existential stakes. In high-stakes tragedies, survival instincts override social norms; if cooperating while others defect leads to elimination, defection becomes an absolute imperative. When survival is at stake, the incentive to defect becomes overwhelming. As Alchian (1950) famously argued, market competition functions as an evolutionary selection mechanism; firms that fail to maximize profit risk elimination.
Historical evidence confirms firms consistently choose expansion over work reduction when facing competitive pressure. Manufacturing plants consistently respond to productivity improvements by expanding output rather than reducing labor input (Baily et al., 1992), and even during periods of declining work hours (1890s-1940s), reductions came primarily through labor movement pressure and legislation rather than voluntary firm decisions (Costa, 2000). The overwork pattern Schor (1992) documented reflects this systematic preference for expansion over welfare improvements.
C4: Pareto-Inefficiency.
Universal welfare improvements Pareto-dominate universal output expansion:
If all firms reduce hours proportionally to productivity gains:
  • Workers gain leisure time without income loss
  • Firms maintain relative market positions (no competitive disadvantage)
  • Total output remains sufficient for material abundance
  • Environmental externalities may decrease with reduced production
Compare this to mutual expansion (the equilibrium):
  • Workers work longer hours despite productivity gains
  • Firms engage in costly competition, reducing profit margins
  • Excess output may exceed demand, requiring artificial stimulus
  • Negative externalities from production increase
The equilibrium is recognizably suboptimal. This is precisely what Keynes observed, i.e., the capacity for leisure exists but remains unrealized. Workers could work less, firms could coordinate restraint, yet competition drives continued expansion.
C5: Enforcement Difficulty.
Multiple structural barriers prevent coordination:
International Anarchy: No global authority can enforce coordinated work reduction. National regulations face race-to-bottom dynamics as firms relocate to jurisdictions with longer hours or fewer restrictions. Capital mobility undermines any single nation’s ability to impose work reductions that competitors don’t match.
Verification Challenges: Even with agreements, compliance is difficult to verify. Firms can disguise overtime, create parallel structures, or use contractors. Workers facing job insecurity may voluntarily work longer hours to demonstrate commitment. Actual working hours may diverge substantially from official regulations.
Monitoring Infeasibility: The distributed nature of modern work, such as remote work, knowledge work, and, always-on digital connectivity, makes monitoring difficult. When work becomes cognitive rather than physical, and occurs outside traditional workplace boundaries, enforcement costs escalate.
Commitment Problems: Even if firms agree to coordinate, commitments lack credibility. Each firm has overwhelming incentive to defect once others reduce hours. Without binding enforcement, any agreement is unstable. First-mover disadvantage discourages unilateral action—the firm that reduces hours first suffers competitive disadvantage if others don’t follow.
Result:
Productivity competition satisfies all five conditions. Therefore, by Theorem 6, universal output expansion (long working hours) is the unique Nash equilibrium despite being Pareto-dominated by universal work reduction. This explains Keynes’s failed prediction structurally.
4.2 Formalization: The N-Firm Game
We now formalize the heuristic arguments from the previous section to demonstrate that the tragedy structure can be derived from first principles under standard assumptions.
Consider $n$ firms indexed by $i\in\{1,\ldots,n\}$ competing in a market. Firms are heterogeneous: firm $i$ has initial output $Q_{i}$ and working hours $H_{i}$ such that productivity is output per hour. We assume that working hours do not exceed the physiological threshold where marginal productivity turns negative.
We define two separate utility functions: one for the firm ($U_{F}$), driven by profit, and one for the workers ($U_{W}$), driven by welfare.
  • Firm Utility ($U_{F}$): This is pure profit: $\text{Revenue}-\text{Labor Cost}$, after ignoring other typical costs such as taxes and costs that may arise in mutual defection state such as additional advertising costs and price wars.
    • $\pi$ is the total market-wide profit.
    • $c$ is the cost per hour of labor.
  • Worker Utility ($U_{W}$): This is welfare: $\text{Wage}+\text{Leisure Value}$.
    • We assume wages are a component of $c$ and are maintained in both scenarios.
    • $\beta$ is the monetized utility value of one hour of leisure. (This is based on Akerlof’s partial gift exchange model (Akerlof, 1982)).
Both utility equations above ignore other components that stay constant between strategy changes.
When productivity increases by factor $\alpha>1$, each firm chooses between two strategies, $s_{i}\in\{C,D\}$:
  • $C$ (Cooperate): Reduce hours proportionally to $H_{i}/\alpha$, maintaining output $Q_{i}$
  • $D$ (Defect): Maintain hours $H_{i}$, expanding output to $\alpha Q_{i}$
Payoff Structure
The utility for firm $i$ depends on its own strategy $s_{i}\in\{C,D\}$ and the strategy profile of its competitors, $s_{-i}$. In this specific market game, the impact of $s_{-i}$ is fully captured by the total output of all other firms, which we denote as $Q_{-i}(s_{-i})$.
The utility functions for firm $i$ and a worker in the firm are therefore:
  • If Cooperating ($s_{i}=C$): Output is $Q_{i}$, Hours are $H_{i}/\alpha$.
    • $U_{F,i}(C,\mathbf{s_{-i}})=\pi\cdot\frac{Q_{i}}{Q_{i}+Q_{-i}(\mathbf{s_{-i}})}-c\cdot\frac{H_{i}}{\alpha}$
    • $U_{W,i}(C)=\text{Wage}+\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha}$
  • If Defecting ($s_{i}=D$): Output is $\alpha Q_{i}$, Hours are $H_{i}$.
    • $U_{F,i}(D,\mathbf{s_{-i}})=\pi\cdot\frac{\alpha Q_{i}}{\alpha Q_{i}+Q_{-i}(\mathbf{s_{-i}})}-c\cdot H_{i}$
    • $U_{W,i}(D)=\text{Wage}$
Proposition 7 (Dominance in Productivity Competition).
For any firm $i$, any competitor strategy profile $s_{-i}$ resulting in total competitor output $Q_{-i}(s_{-i})>0$, and any productivity factor $\alpha>1$:
  1. (i)
    Defection yields strictly higher market share than cooperation.
  2. (ii)
    Defection is the dominant strategy, i.e., $U_{i}(D,s_{-i})>U_{i}(C,s_{-i})$ for all $s_{-i}$, if the profit gain from this increased market share always exceeds the (constant) net cost of forgoing leisure.
Proof.
(i) The proof that market share for Defecting is strictly greater is unchanged and holds by $\alpha>1$. (ii) We find the net profit gain for the firm from defecting:
$U_{F,i}(D,s_{-i})-U_{F,i}(C,s_{-i})$ $=\left(\pi\cdot\frac{\alpha Q_{i}}{\alpha Q_{i}+Q_{-i}(s_{-i})}-c\cdot H_{i}\right)$
$\quad-\left(\pi\cdot\frac{Q_{i}}{Q_{i}+Q_{-i}(s_{-i})}-c\cdot\frac{H_{i}}{\alpha}\right)$
Grouping terms:
$U_{F,i}(D)-U_{F,i}(C)$ $=\underbrace{\pi\left(\frac{\alpha Q_{i}}{\alpha Q_{i}+Q_{-i}(s_{-i})}-\frac{Q_{i}}{Q_{i}+Q_{-i}(s_{-i})}\right)}_{\text{Market Share Gain (Always Positive)}}$
$\quad-\underbrace{\left(c\cdot H_{i}-c\cdot\frac{H_{i}}{\alpha}\right)}_{\text{Increased Labor Cost (Always Positive)}}$
Simplifying the cost term:
$U_{F,i}(D)-U_{F,i}(C)=\text{Market Share Gain}-\text{Increase Labor Cost}$
Defection is the dominant strategy (C3 is met) with the assumption that the profit from the Market Share Gain is greater than the Increased Labor Cost from not reducing hours.
Historical evidence confirms this assumption holds in practice: firms consistently choose output expansion when productivity increases (Baily et al., 1992; Costa, 2000; Schor, 1992), demonstrating that competitive pressure overwhelms leisure preferences across diverse market structures and firm sizes.
Note that for workers, defection leads to a decrease in utility since
$U_{W,i}(D)-U_{W,i}(C)=(\text{Wage})-(\text{Wage}+\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha})=-\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha}$
This indicates a welfare opportunity loss for workers.
Proposition 8 (Pareto-Inefficiency in Productivity Competition).
Universal Cooperation ($s^{C}$) strictly Pareto-dominates Universal Defection ($s^{D}$).
Proof.
We must show that $U_{i}(s^{C})>U_{i}(s^{D})$ for all actors (both firms and workers). The key insight is that when all firms scale output by the same factor, relative market shares remain unchanged.
1. For Firms ($U_{F}$):
  • $U_{F,i}(s^{C})=\pi\cdot\frac{Q_{i}}{\sum_{j}Q_{j}}-c\cdot\frac{H_{i}}{\alpha}$
  • $U_{F,i}(s^{D})=\pi\cdot\frac{\alpha Q_{i}}{\sum_{j}\alpha Q_{j}}-c\cdot H_{i}$
The difference is:
$U_{F,i}(s^{C})-U_{F,i}(s^{D})$ $=\left(-c\cdot\frac{H_{i}}{\alpha}\right)-\left(-c\cdot H_{i}\right)$
$=\underbrace{c\cdot\frac{(\alpha-1)H_{i}}{\alpha}}_{\text{Labor Cost Savings}}$
This value is strictly positive (assuming $c>0$). Firms are better off under $s^{C}$.
2. For Workers ($U_{W}$):
  • $U_{W,i}(s^{C})=\text{Wage}+\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha}$
  • $U_{W,i}(s^{D})=\text{Wage}$
The difference is:
$U_{W,i}(s^{C})-U_{W,i}(s^{D})=\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha}$
This value is strictly positive (assuming $\beta>0$). Workers are better off under $s^{C}$.
Since both firms and workers are strictly better off, $s^{C}$ Pareto-dominates $s^{D}$. The tragedy holds: firms’ individually rational dominant strategy (Proposition 7) leads to a unique Nash equilibrium ($s^{D}$) that is worse for everyone.
Crucially, this result holds for all firms simultaneously regardless of their sizes, initial hours, or market positions. Large and small firms, those with long and short hours, all benefit from coordinated restraint. The Pareto-domination is genuine as there exists no firm that prefers mutual defection to mutual cooperation. ∎
Result: Combined with the N-player structure, externalities, and enforcement difficulty established in Section 4, productivity competition satisfies all five conditions for structural tragedy. The coordination failure emerges from the fundamental competitive structure of markets facing productivity shocks, not from special assumptions about firm homogeneity or market structure.
4.3 Formalization: Extensions, Robustness, and Limitations
Having established the core coordination failure for working hours under uniform productivity gains, we now examine extensions to other welfare dimensions, robustness to relaxed assumptions, and limitations of the analysis.
4.3.1 Extension to Multiple Welfare Dimensions
While we formalize the coordination failure for working hours, the same structural logic applies to other dimensions of worker welfare. When productivity increases, firms face analogous binary choices:
  • Wages: Raise proportionally (C) vs. maintain while expanding output (D)
  • Benefits: Enhance (C) vs. maintain while expanding output (D)
  • Job security: Stabilize employment (C) vs. maintain flexibility while expanding (D)
Each dimension exhibits the same five conditions for structural tragedy. Which dimension manifests the coordination failure depends on institutional constraints. Hour regulations shift pressure to wages; wage floors shift pressure to employment; strong employment protections shift pressure to benefits. The declining labor share of income globally (Karabarbounis and Neiman, 2014; Elsby et al., 2013) provides evidence that productivity gains systematically flow away from worker welfare across all dimensions, though the specific manifestation varies by institutional context.
4.3.2 Relaxing the Uniform Productivity Assumption
The baseline analysis assumes all firms experience the same productivity gain $\alpha$. This simplification is most realistic for economy-wide technological shifts (e.g., electrification, computerization, internet) that eventually diffuse broadly, or for long-run analysis after differential adoption rates have equilibrated through competitive selection.
Dominance under heterogeneity. With heterogeneous productivity gains $\alpha_{i}$ varying across firms, the dominance result (Proposition 7) remains fully valid: each firm $i$ has incentive to expand given its own $\alpha_{i}>1$, regardless of others’ productivity gains. The proof is identical, replacing $\alpha$ with $\alpha_{i}$ throughout. Therefore the core mechanism driving the tragedy—individual incentive to defect regardless of what others do—persists under heterogeneous productivity shocks.
Pareto-inefficiency under heterogeneity. The Pareto-inefficiency result (Proposition 8) requires modification. With heterogeneous $\alpha_{i}$, workers remain strictly better off under $s^{C}$, but the outcome for firms depends on the distribution of productivity gains:
$U_{F,i}(s^{C})-U_{F,i}(s^{D})$ $=\left(\frac{Q_{i}}{\sum_{j}Q_{j}}-\frac{\alpha_{i}Q_{i}}{\sum_{j}\alpha_{j}Q_{j}}\right)+\left(-c\cdot\frac{H_{i}}{\alpha_{i}}\right)-\left(-c\cdot H_{i}\right)$
$=\underbrace{\pi\cdot Q_{i}\left(\frac{\sum_{j}(\alpha_{j}-\alpha_{i})Q_{j}}{(\sum_{j}Q_{j})\cdot(\sum_{j}\alpha_{j}Q_{j})}\right)}_{\text{Market Share Change}}+\underbrace{c\cdot\frac{(\alpha_{i}-1)H_{i}}{\alpha_{i}}}_{\text{Labor Cost Savings}}$
Firms with higher productivity gains capture larger market shares under universal defection. Whether universal cooperation Pareto-dominates then depends on the distribution of productivity gains. If productivity differences are modest, cooperation can still dominate. If productivity differences are extreme, some high-productivity firms might prefer universal defection as it allows market share capture from lower-productivity rivals.
Critically, heterogeneity makes coordination more difficult, not easier. Firms with productivity advantages face even stronger temptation to defect and capture market share from lower-productivity rivals. This reinforces rather than undermines our structural analysis, as the tragedy becomes more severe when productivity gains are unequally distributed.
4.3.3 Generality and Robustness
The formalization demonstrates that productivity competition exhibits structural tragedy under general conditions:
Beyond binary choices. The dominance property is not an artifact of the binary choice model. It extends to continuous strategy spaces where firms choose specific hours $h_{i}\in[0,H_{max}]$. In such settings, the marginal incentive to increase output (to capture market share) persists at every level of $h_{i}$ provided the competitive elasticity is sufficiently high.
Risk aversion strengthens the tragedy. Since the utility loss from losing market share (existential risk) outweighs the utility gain from leisure, risk-averse firms face even stronger pressures to defect than risk-neutral ones.
Generality across market structures. The tragedy holds for:
  • Firms of any size distribution (concentrated or competitive markets)
  • Any initial working hours across firms
  • Any productivity gain $\alpha>1$
  • Any configuration of competitor choices
  • Heterogeneous productivity gains $\alpha_{i}$ across firms
4.3.4 Welfare Cost Increases with Productivity
Sensitivity analysis of the utility functions reveals a counter-intuitive dynamic: the welfare cost of the coordination failure increases with technological progress. We define the “cooperative surplus” as the difference between welfare in the cooperative outcome ($s^{C}$) and the Nash equilibrium ($s^{D}$). As the productivity factor $\alpha$ rises, this surplus grows asymptotically toward the total labor cost, meaning the gap between potential leisure and actual leisure widens.
This implies that highly productive economies sacrifice significantly more potential welfare than low-productivity ones. Technological abundance amplifies rather than resolves the structural waste. Modern economies with eightfold productivity gains since Keynes could provide far more leisure than 1930s economies, yet the coordination failure prevents realizing this potential.
4.4 Empirical Validation: The European Case
The European experience provides crucial empirical validation as a best-case scenario. If our structural analysis is correct, even under maximally favorable conditions for coordination—wealthy democracies with strong institutions—we should observe:
  1. 1.
    Partial override: coordination achieves some welfare improvements but not full Keynesian leisure
  2. 2.
    High costs: coordination requires substantial institutional apparatus
  3. 3.
    Erosion over time: competitive pressure gradually undermines coordination
The evidence strongly supports these predictions.
Favorable Conditions.
Europe represents the most favorable case for coordination:
  • Wealthy democracies with strong institutions
  • Supranational governance (European Union) providing coordination capacity
  • Cultural consensus supporting work-life balance
  • Strong labor unions with political influence
  • Legal frameworks mandating maximum hours and minimum vacation
If coordination is possible anywhere, it should succeed in Europe.
Partial Override.
European workers have achieved shorter hours than American workers, e.g., about 1,350 hours in Germany vs. about 1,800 hours in the US (OECD, 2025). Workers in many developing economies face even longer hours with weaker protections, demonstrating that the coordination failure intensifies under less favorable institutional conditions.
This represents real achievement: European coordination reduced hours by approximately 450 compared to the United States. However, even German workers work far more than Keynes’s predicted 15 hours weekly (780 hours annually). The coordination is partial, not complete.
Moreover, productivity-welfare decoupling manifests differently across institutional contexts. While European workers gained some leisure relative to their US counterparts, Europe also exhibits:
  • Declining labor share of income since 1980, similar to the US trend
  • Slower wage growth relative to productivity in many sectors
  • Persistent pressure on social benefits and employment protections
  • Higher unemployment rates in several major economies
The tragedy is not escaped but redistributed: European workers gained some leisure at the cost of wage growth and employment, while US workers experienced stagnant wages and benefits despite longer hours. Both regions show the fundamental pattern our framework predicts: productivity-welfare decoupling driven by competitive pressure, manifesting through different dimensions (hours vs. wages vs. benefits vs. job security) depending on institutional context.
High Costs.
This coordination came with substantial costs:
Economic costs: European GDP growth has lagged the United States annually over recent decades; for example, data indicates a significant divergence in aggregate growth trajectories between EU and the USA since 2008 (World Bank, 2024). While this divergence is driven by multiple factors, shorter working hours impose a mechanical ceiling on total output.
Regulatory costs: Maintaining shorter hours requires extensive regulation. France’s 35-hour workweek involves complex rules about overtime, exceptions, and compliance monitoring (L&E Global, 2024). Germany’s system involves workplace councils, sector-level bargaining, and intricate labor law.
Competitive pressure: European firms face disadvantages competing against firms from regions with longer working hours. This shows up in market share, innovation metrics, and firm valuations. The competitive pressure remains even under coordination.
Erosion Over Time.
Most tellingly, the coordination shows signs of erosion:
Enforcement weakening: France’s 35-hour law contains numerous exceptions and workarounds. Actual hours often exceed nominal limits through overtime, managerial exemptions, and definitional ambiguities (L&E Global, 2024).
Pressure for reform: Germany’s Hartz reforms increased flexibility and effectively lengthened working hours for some sectors (IAB-Forum, 2023).
Competitive concerns: European policy discussions increasingly focus on competitiveness relative to the United States and China, with pressure to loosen labor regulations to boost growth.
Hours stabilization, not decline: European working hours have largely stabilized since the 1980s despite continued productivity growth (Huberman and Minns, 2007). The decline in hours observed from 1870-1980 has stalled, suggesting coordination has reached its limits.
Interpretation.
The European case validates our framework in three ways:
First, even under maximally favorable conditions, coordination is partial. European workers still work far more than technological capacity allows. The tragedy is attenuated but not resolved.
Second, coordination is costly. Substantial institutional apparatus is required, and even this apparatus faces continuous pressure and erosion. The coordination is fragile, not robust.
Third, competitive pressure persists and manifests across multiple dimensions. European firms face disadvantages that generate political pressure to loosen regulations. Even where hours were reduced, other dimensions of worker welfare (wages, employment, benefits) came under pressure. The underlying tragic structure remains even when temporarily overridden by institutions.
This demonstrates that the productivity tragedy is structural, not merely cultural or policy-related. Even when culture supports leisure, institutions enable coordination, and wealth permits reduced hours, the competitive structure creates barriers that limit welfare improvements and erode over time.
4.5 Why Standard Solutions Face Structural Barriers
Our structural analysis explains why proposed solutions face barriers:
Solution 1: Individual Firm Action
Problem: Violates dominance condition. Any firm reducing hours unilaterally suffers competitive disadvantage. Competitors capture market share, potentially forcing the cooperative firm out of business. First-mover disadvantage is severe.
Solution 2: Industry-Level Agreements
Problem: Violates enforcement condition. Agreements lack binding force absent government enforcement. Each firm has incentive to defect secretly. Even with monitoring, defection detection may come too late to prevent competitive harm.
Solution 3: National Regulation
Problem: Faces international competition. Capital mobility enables firms to relocate to jurisdictions with longer hours or lower labor costs. This race-to-the-bottom dynamic affects both wealthy and developing nations, with the latter often experiencing more intense pressure to attract investment through weaker labor protections.
Solution 4: International Agreements
Problem: Violates enforcement condition. International labor standards (ILO conventions) lack enforcement mechanisms. Compliance is voluntary. Countries competing for investment have incentives to defect through weak enforcement or exceptions.
Solution 5: Cultural Change
Problem: Insufficient against structural pressure. Even if workers prefer leisure, competitive pressure compels long hours. Firms facing bankruptcy cannot afford to honor worker preferences. Cultural change may be necessary but is insufficient absent structural change.
Solution 6: Technological Unemployment
Problem: Historical evidence suggests employment adjusts rather than hours reducing. When technology displaces workers from one sector, they typically move to other sectors rather than working less (Autor, 2015). Unemployment may increase but this differs from coordinated work reduction.
These solutions all fail because they don’t address the underlying structure. The tragedy persists because the five conditions remain satisfied. Only interventions that break at least one condition could resolve the tragedy.
4.6 What Would Work
Based on our framework, effective interventions must break at least one of the five conditions:
Break Condition 1 (N-Player): Reduce the number of independent actors. This is difficult—markets inherently involve multiple firms, and consolidation faces antitrust constraints. Complete monopoly might enable work reduction but at costs to competition and innovation.
Break Condition 2 (Externalities): Eliminate the externality where one firm’s expansion harms others. This seems impossible since competitive pressure is inherent to markets. Work reduction by one firm creates market share opportunities for others by definition.
Break Condition 3 (Dominance): Reduce existential stakes making defection dominant. This is most promising. If firms didn’t face bankruptcy risk from work reduction, coordination becomes feasible. Mechanisms:
  • Universal Basic Income reduces worker pressure for long hours
  • Stronger social safety nets reduce survival stakes
  • Reduced international competition lessens competitive pressure
  • Market structures that reward restraint (currently absent)
Break Condition 4 (Pareto-Inefficiency): Accept the equilibrium rather than seeking coordination. Some economists argue long hours reflect preferences rather than coordination failure (Pencavel, 2018). Our framework suggests this is incorrect as the Pareto-domination is real, and workers would benefit from coordination.
Break Condition 5 (Enforcement): Solve the enforcement problem through effective global labor standards with monitoring and sanctions. This requires:
  • International agreement on maximum hours
  • Verification mechanisms for compliance
  • Sanctions for defection sufficient to deter
  • Sufficient participation to prevent race-to-bottom
This is theoretically possible but practically difficult given international anarchy and sovereignty concerns.
The productivity tragedy can be solved, but only through interventions that address its structural foundations. Incremental reforms that leave the five conditions intact will achieve at most partial, costly, eroding coordination as observed in Europe.
5 The Tragedy of AI Governance
We now apply the unified framework to AI development, demonstrating that AI governance faces the same five conditions but with amplified intensity. This makes coordination structurally more difficult than for historical arms control cases including nuclear, biological, and chemical (NBC) weapons.
5.1 Mapping the Five Conditions to AI Development
Condition 1: N-Player Structure
AI development involves numerous independent actors: AI labs, tech companies, universities and research institutions, nations, and open-source communities.
The number of actors is large and growing. No single actor can coordinate AI development unilaterally. The architecture is fundamentally decentralized across private sector, public sector, and international actors.
Condition 2: Binary Choice with Externalities
When AI capabilities advance, each actor faces essentially a binary choice:
  • Strategy $C$ (Cooperate/Restrain): Slow development to ensure safety, conduct thorough testing, wait for governance mechanisms
  • Strategy $D$ (Defect/Advance): Accelerate capabilities development to maintain competitive position
Individual choices create negative externalities. When one actor accelerates development:
  • Competitors face pressure to accelerate or lose market/strategic position
  • Safety margins narrow as capabilities race ahead of governance
  • Risks from powerful AI systems affect everyone regardless of who develops them
  • First-mover advantages create winner-take-all dynamics
The externality is severe since one actor’s defection substantially increases risks for all.
Condition 3: Dominance Property
Acceleration dominates restraint regardless of competitors’ choices:
Case 1—If competitors restrain: Actor gains decisive first-mover advantage by accelerating. In AI markets characterized by network effects, data advantages, and talent concentration, first-mover advantages are enormous. The actor that reaches advanced capabilities first may capture the entire market.
Case 2—If competitors accelerate: Actor must accelerate to avoid strategic elimination. Falling behind in AI capabilities may mean permanent market displacement for firms or strategic vulnerability for nations. When competitors accelerate, restraint means obsolescence.
The dominance is driven by existential stakes for firms, nations, and researchers. For example, a firm who cannot raise capital for massive compute (or acceleration) risk obsolescence. This structural pressure forced even the leading non-profit actor to restructure first into a “capped-profit” entity (OpenAI, 2019) and eventually into a Public Benefit Corporation to remove fundraising limits (OpenAI, 2025).
Condition 4: Pareto-Inefficiency
Universal restraint Pareto-dominates universal acceleration:
If all actors coordinate to slow development and prioritize safety:
  • Risks from advanced AI are reduced through careful development
  • Relative competitive positions are preserved (no disadvantage from restraint)
  • Time is available for governance mechanisms to develop
  • Benefits from AI still accrue but with better risk management
Compare this to mutual acceleration (the equilibrium):
  • Risks accumulate faster than governance mechanisms develop
  • Safety testing is compressed or skipped under time pressure
  • First-mover races waste resources on competitive positioning
  • Potential for catastrophic outcomes that benefit no one
Leading AI researchers explicitly recognize this Pareto-domination. Surveys show substantial minorities estimating non-trivial probabilities of catastrophic outcomes (Bengio et al., 2025; Grace et al., 2018). The AI safety community argues current trajectories are collectively harmful (Askell et al., 2019). Yet development accelerates.
Asymmetric Gains and the Winner-Take-All Problem AI development creates potential for extremely asymmetric outcomes. Unlike productivity competition where firms maintain roughly proportional positions, AI capabilities races may produce winner-take-all dynamics where the first to achieve advanced AI captures decisive advantages, potentially controlling entire markets, achieving strategic dominance, or even technological singularity.
This asymmetry threatens the Pareto-inefficiency condition. If an actor believes they will be the winner, they might rationally prefer the racing outcome (where they capture enormous gains) over a cooperative outcome (where advantages are constrained or shared). For prospective winners, the situation appears not as a tragedy but as a rational pursuit of dominance.
However, three factors restore the tragedy structure even for prospective winners:
First, existential risk is non-excludable. The catastrophic risks from misaligned or insufficiently robust advanced AI systems affect all actors regardless of who develops them first (Bostrom, 2014). A winner who achieves advanced AI through hasty development faces the same risks of loss of control, unintended consequences, or catastrophic failure as any other actor. The instrumental convergence thesis (Bostrom, 2014) suggests that advanced AI systems will pursue subgoals (self-preservation, resource acquisition, goal preservation) that may conflict with human interests regardless of which actor deploys them. Winner and losers alike face potential extinction or permanent disempowerment.
Crucially, the tragedy persists because the rate of capability gains in a race condition naturally outpaces the rate of safety verification. Consequently, the winner is statistically likely to deploy an unsafe system before establishing control, rendering the expected utility of “winning” negative even for the victor. This conclusion is reinforced by empirical data: Bengio et al. (2025); Grace et al. (2018) found that substantial minorities of AI researchers estimate non-negligible probabilities of extremely bad outcomes including human extinction. If these probability estimates are even roughly accurate, the astronomical loss associated with an unsafe deployment outweighs even enormous competitive gains, making the race irrational for all participants.
Second, winner identity is uncertain ex ante. While racing, no actor knows with certainty they will be the winner. Multiple actors have credible capabilities. Geopolitical and corporate competition involves numerous sophisticated actors, any of whom might achieve breakthroughs. The “DeepSeek Shock” of early 2025 serves as a potent empirical validator of this uncertainty. By achieving frontier-level reasoning capabilities with orders-of-magnitude greater training efficiency, a resource-constrained actor demonstrated that massive capital and hardware “moats” are not impregnable (DeepSeek-AI, 2025). This reintroduced volatility into the race, proving that “winner” status is fragile and susceptible to algorithmic disruption. From behind the veil of this uncertainty, all actors face expected losses from racing even if the ex post winner (whoever that turns out to be) might retrospectively have benefited.
This parallels arms races where no state knows ex ante whether it will achieve dominance. Even if one state would retrospectively prefer the outcome where it achieved hegemony, all states prefer coordinated restraint to a race with uncertain winners and certain risks.
Third, even decisive technological leads may be temporary. Unlike traditional military conquest where territory captured remains under control, AI capabilities can potentially be copied, stolen, or independently rediscovered. Winner-take-all assumptions may prove mistaken if multiple actors achieve advanced AI within relatively short timeframes, or if AI capabilities prove impossible to monopolize due to espionage, independent discovery, or proliferation.
Therefore, even with extremely asymmetric potential gains, the Pareto-inefficiency condition holds when existential risks are incorporated into rational calculation. The racing equilibrium threatens all actors including prospective winners, making coordinated restraint collectively preferable despite competitive temptations.
Critical caveat: This defense of Pareto-inefficiency relies on two empirical claims: (1) catastrophic risks from advanced AI are sufficiently probable and severe to dominate even winner-take-all competitive gains, and (2) actors incorporate these risks into their decision-making rather than pure output or power maximization. The first claim is defended by substantial technical literature on AI safety (Bostrom, 2014; Russell, 2019; Amodei et al., 2016). The second claim is more tenuous—if actors systematically discount existential risks or believe they can control advanced AI despite expert warnings, the tragedy structure may indeed break for prospective winners, transforming the coordination problem into a more traditional conquest dynamic.
Condition 5: Enforcement Difficulty
Multiple structural barriers prevent coordination:
International Anarchy: No global authority can enforce coordinated AI governance. Nations compete for AI advantages. International agreements lack binding force. AI is dual-use, making arms control agreements difficult to negotiate.
Verification Impossibility: AI development occurs in software, making verification extremely difficult. Model weights can be copied and modified. Training runs can occur covertly.
Monitoring Infeasibility: Even with agreements, monitoring compliance faces severe challenges. What constitutes “dangerous” AI capabilities may be ambiguous until after development. The distributed nature of AI research across many institutions resists centralized monitoring.
Commitment Problems: Even if actors agree to coordinate, commitments lack credibility. Each actor has overwhelming incentive to defect and gain first-mover advantages once others restrain. Secret defection is technologically feasible. First-mover disadvantage from restraint is severe.
Result: AI development satisfies all five conditions with high intensity when catastrophic risks are incorporated into rational assessment. The Pareto-inefficiency condition is satisfied for all actors including prospective winners because existential risks are non-excludable, winner identity is uncertain ex ante, and even decisive leads may prove temporary. Therefore, by Theorem 6, universal acceleration is the unique Nash equilibrium despite being Pareto-dominated by coordinated restraint.
This analysis assumes actors incorporate existential risks into rational calculations, consistent with survey evidence showing substantial minorities of AI researchers estimate non-trivial extinction probabilities (Bengio et al., 2025; Grace et al., 2018). Actors who systematically discount such risks would not face the structural tragedy we identify, transforming the problem from coordination failure to conquest dynamics. However, the empirical evidence suggests leading actors do recognize these risks even while competitive pressure compels continued development—precisely the tragedy structure our framework predicts.
Robustness of the Racing Dynamic.
Just as with productivity, this tragedy structure is robust to relaxed assumptions. The dominance of acceleration holds in continuous strategy spaces: if actors choose a level of investment or compute $k_{i}\in[0,K_{max}]$, the marginal strategic advantage of increasing $k_{i}$ remains positive as long as competitors are advancing. Consequently, “slowing down” is not a stable equilibrium; the Nash equilibrium collapses to maximum feasible acceleration ($K_{max}$). Furthermore, strategic risk aversion (fear of falling behind) amplifies the race. Because the loss of sovereignty or market relevance is viewed as a certain, immediate catastrophe, while safety failure is viewed as a probabilistic, future catastrophe, risk-averse actors rationally prioritize avoiding the immediate deterministic loss, thereby intensifying the acceleration.
5.2 AI Governance vs. Historical Arms Control: An Eight-Dimension Comparison
To assess whether AI governance is more or less difficult than previous coordination challenges, we systematically compare AI to successful arms control cases. Nuclear, biological, and chemical (NBC) weapons, all achieved substantial (though imperfect) international coordination. How does AI compare?
Table 3 summarizes the comparison across eight crucial dimensions.
Table 3: AI Governance vs. Historical Arms Control Across Eight Dimensions
Dimension Chemical / Biological Nuclear AI
Strategic Value Military only, limited Military, high Military + Economic, extreme
Dual-Use Separable (some) Separable (enrichment) Inseparable
Verification Observable (facilities) Observable (materials) Unobservable (software)
Reputational Costs High (taboo) High (stigma) Low (prestige)
Entry Barriers Moderate Very High Low
Substitutes Yes (conventional) Yes (conventional) None
Time Scale Decades warning Decades warning Years warning
Convergence Military only Military only Military + Economic
We now discuss each dimension in detail.
Dimension 1: Strategic Value
Chemical and biological weapons have limited strategic value. They are unreliable, difficult to control, and often militarily ineffective compared to conventional weapons. This limited value reduced incentives for development and made restraint easier.
Nuclear weapons have high strategic value for deterrence, making complete elimination difficult. However, their value is purely military. Economic incentives for nuclear development are weak.
AI has extreme strategic value in both military and economic domains. Military applications include autonomous weapons, intelligence analysis, cyber operations, and logistics optimization. Economic applications include productivity enhancement, automation, market advantage, and platform dominance. This dual-domain value creates reinforcing incentives far stronger than for any previous arms control case.
Dimension 2: Dual-Use Inseparability
Chemical and biological agents have some separability between civilian and military applications. Industrial chemicals differ from weaponized agents. Civilian biology can be distinguished from weaponization programs with moderate effort.
Nuclear programs have some separability. Civilian nuclear power uses low-enriched uranium; weapons require highly enriched uranium or plutonium. Monitoring enrichment levels provides some verification capability.
AI research and development is completely inseparable between civilian and military applications. The same models, training techniques, and hardware serve both purposes. Foundational research benefits both domains simultaneously. Attempting to restrict military AI while permitting civilian AI is structurally impossible—they are identical technologies with different deployment contexts.
This inseparability creates severe enforcement problems. Any agreement to restrict military AI would face immediate problems: civilian AI companies would argue their research is purely civilian, verification would be impossible, and the distinction itself would be meaningless.
Dimension 3: Verification
Chemical and biological weapons require observable facilities leaving physical traces. Nuclear weapons require large enrichment facilities and detectable materials, enabling moderate verification success. AI development occurs in software and is fundamentally unverifiable—model weights can be copied instantly, training runs can occur covertly, and testing for dangerous capabilities requires developing the very systems one seeks to restrict.
Dimension 4: Reputational Costs
Chemical and biological weapons face powerful taboos with strong international norms (Price, 1995; Organisation for the Prohibition of Chemical Weapons (2025), OPCW; United Nations Office for Disarmament Affairs, 2025). Nuclear weapons face mixed stigma and prestige. AI development faces inverse reputational dynamics, leading development confers prestige rather than stigma, making restraint appear as weakness rather than responsibility.
Dimension 5: Entry Barriers
Nuclear weapons face very high barriers (billions of dollars, advanced technology, years of development) that concentrated capabilities in few hands, enabling coordination. AI faces low and declining entry barriers: frontier development costs hundreds of millions but is accessible to many actors and declining with algorithmic progress, multiplying the coordination problem.
Dimension 6: Substitutes
Chemical, biological, and nuclear weapons have conventional substitutes for most military objectives. AI has no substitutes for either economic or military applications—actors who restrain fall irreversibly behind, amplifying the dominance condition.
Dimension 7: Time Scale
Chemical and nuclear weapons had decades between emergence and arms control (1925-1993 for chemical, 1942-1968 for nuclear). AI governance faces compressed timelines: modern deep learning emerged around 2012, with expert surveys estimating a 50% chance of human-level AI within roughly two decades (Grace et al., 2018), leaving far less time for institutional development.
Dimension 8: Economic-Military Convergence
Chemical and biological weapons converged only military considerations. Nations developed and deployed them (when they did) for military advantage, making the coordination problem purely security-focused.
Nuclear weapons similarly converged only military considerations, with the exception of some civilian nuclear power development which proved separable.
AI represents complete convergence of economic and military imperatives. Every advance in civilian AI capabilities immediately enhances military capabilities. Every AI laboratory contributes simultaneously to economic competitiveness and military potential. Nations cannot choose to prioritize economic development over military security or vice versa—AI advancement serves both goals identically.
This convergence creates reinforcing pressures absent in previous coordination challenges. Even if nations could overcome military competition (as partial nuclear arms control achieved), economic competition would independently drive AI development. Even if firms could coordinate to slow capabilities races for safety, military imperatives would independently accelerate development. The dual reinforcement makes coordination structurally more difficult than previous single-domain challenges.
Summary Across Dimensions
AI faces unfavorable conditions on all eight dimensions compared to successful historical arms control:
  • Higher strategic value (dual-domain)
  • Complete dual-use inseparability
  • Fundamental verification impossibility
  • Reversed reputational incentives
  • Lower entry barriers multiplying actors
  • No substitutes increasing opportunity costs
  • Compressed timelines preventing institutional development
  • Economic-military convergence creating reinforcing pressures
This systematic comparison demonstrates that AI governance faces coordination challenges structurally worse than any previous arms control case. AI faces unfavorable conditions on every dimension simultaneously. The inseparability of economic and military AI (the dual-use dilemma) is the decisive structural barrier: stopping military AI requires stopping economic AI, which is suicidal for a nation’s productivity.
5.3 Empirical Validation: The Russia-Ukraine Drone War
The Russia-Ukraine conflict validates the framework’s structural predictions, not for autonomous AI specifically, but for the competitive dynamics that will govern its development. When Russia invaded Ukraine in February 2022, both sides had minimal drone capabilities. Within two years, deployment escalated from dozens to thousands of drones monthly (Center for Strategic and International Studies, 2024; Watling and Reynolds, 2023), with Ukraine expected to produce 4.5 million drones in 2025 (Nieczypor and Matuszak, 2025). This exponential growth occurred despite years of prior international dialogue on weapons governance. While current drones remain predominantly remotely piloted rather than AI-controlled, the success of drone warfare in this and other recent conflicts is expected to increase drone use in future warfare (Fukuyama, 2024), with autonomous AI-controlled systems a likely next stage as capabilities mature.
The escalation pattern demonstrates our framework’s core dynamic: once a technology confers decisive military advantage under existential stakes, competitive pressure compels rapid adoption regardless of prior governance intentions. Each side faced overwhelming incentive to match the other’s deployment (C3: dominance driven by existential stakes). Drone deployment by one side created immediate pressure for counter-deployment (C2: negative externalities). No international authority could enforce restraint during active conflict (C5: enforcement impossibility). Pre-war governance discussions proved irrelevant once competitive pressure intensified—precisely the trajectory our framework predicts for autonomous AI systems as capabilities mature.
This case validates key structural predictions: when existential stakes are present, competitive dynamics overwhelm governance aspirations; escalation occurs rapidly when technology enables it; dual-use technology resists governance attempts; and verification and enforcement fail during intense competition. The drone war thus serves as both confirmation of the framework and a leading indicator for AI governance challenges ahead.
5.4 Why Standard Solutions Face Structural Barriers
Our framework explains why proposed AI governance solutions face fundamental challenges:
Solution 1: Voluntary Pauses
Problem: Violates dominance condition. As noted in the context of military AI arms control, strategic stability concerns make voluntary restraint inherently unstable (Maas, 2019). Consequently, the 2023 open letter calling for an AI pause produced no actual pause despite thousands of signatures (Future of Life Institute, 2023).
Solution 2: International Treaties
Problem: Violates enforcement condition. Treaties face international anarchy plus verification impossibility given software’s intangible nature. Unlike nuclear non-proliferation (which had high entry barriers and decades for institution-building), AI lacks favorable conditions.
Solution 3: Laboratory Self-Governance
Problem: Violates dominance condition. Safety-conscious labs face first-mover disadvantage as they may be eliminated by competitive pressure, creating tragic selection for less cautious actors.
Solution 4: National Regulation
Problem: Violates enforcement condition. AI development can relocate to permissive jurisdictions. No single nation can regulate global development. Race-to-bottom dynamics undermine unilateral regulation.
Solution 5: Technical Safety Research
Problem: While crucial, technical safety research cannot solve the coordination problem. Even with technically safe systems, competitive pressure drives deployment before adequate verification. Safety becomes a constraint that competitive actors work around rather than a prerequisite they satisfy.
Solution 6: Compute Governance
Problem: Restricting access to specialized AI hardware represents the most physically verifiable intervention point, as advanced chips provide a tangible chokepoint that software development lacks. However, this advantage erodes under competitive pressure: nations race to develop domestic semiconductor capacity, algorithmic efficiency reduces hardware requirements over time, and the dual-use nature of computing infrastructure prevents meaningful restrictions without crippling broader economic development. Compute governance may offer a temporary window but faces the same structural dynamics—competitive pressure overwhelming coordination—that the framework identifies across all five conditions.
Solution 6: Energy Monitoring
Problem: Energy monitoring faces similar dynamics. Data center energy consumption is observable and growing rapidly, suggesting energy infrastructure as a potential verification mechanism. However, nations must invest in energy infrastructure for reasons entirely independent of AI such as grid modernization, electrification, industrial policy. Restricting energy investment to limit AI development is economically and politically infeasible. Any agreement to monitor energy for AI governance purposes would be trivially circumvented by dual-use justifications.
These solutions fail because they leave underlying conditions intact. Only interventions that break at least one of the five conditions have prospects for success.
5.5 What Would Work
Based on our framework, effective interventions must break at least one of the five conditions. For AI governance, this is structurally more difficult than for productivity due to amplified intensity across all conditions.
Break Condition 1 (N-Player): Consolidate AI development in a single international consortium or heavily regulated monopoly. This faces severe feasibility constraints: national security concerns prevent international consolidation, while domestic monopoly risks concentrating existential-risk technologies in single entities. The structural barriers are formidable.
Break Condition 2 (Externalities): Eliminate the externality where one actor’s AI advancement threatens others. This appears structurally impossible since AI capabilities inherently create strategic advantages that disadvantage competitors, whether in economic or military domains.
Break Condition 3 (Dominance): Reduce existential stakes making defection dominant. Mechanisms might include: (1) credible international agreements reducing security competition, (2) economic restructuring reducing winner-take-all dynamics, or (3) demonstrated proof that advanced AI poses catastrophic risks even to developers. This is theoretically possible but requires fundamental shifts in international relations and economic structure.
Break Condition 4 (Pareto-Inefficiency): Accept the racing equilibrium rather than seeking coordination. Some argue racing toward advanced AI is inevitable and should be embraced. However, our framework suggests this accepts catastrophic risks that all actors would prefer to avoid if coordination were possible.
Break Condition 5 (Enforcement): Solve the enforcement problem through effective monitoring and verification mechanisms. Unlike software, which is inherently unverifiable, physical infrastructure (compute, energy, facilities) offers potential verification points. However, as discussed in the previous section, these advantages face structural erosion. Effective enforcement would require not merely technical verification capabilities but sustained international cooperation to prevent the competitive dynamics that undermine physical chokepoints, a coordination challenge as severe as the underlying AI governance problem itself.
The analysis suggests AI governance requires interventions of unprecedented scope and coordination. Breaking any single condition might suffice in principle, but the amplified intensity across all five conditions makes each intervention structurally more difficult than for historical arms control cases. This is not pessimism but structural realism: understanding the barriers is prerequisite to overcoming them.
5.6 Reconciliation with Conservative Productivity Estimates
Acemoglu (2025) provides conservative estimates of AI’s total productivity impact (approximately 0.55% annual growth over ten years). However, our framework operates independently of magnitude: even modest advantages create competitive pressure under competitive conditions because, mathematically speaking, the gap between two exponential growths widens with time, meaning economic and military advantages due to AI compound. Even conservative estimates show AI confers sufficient advantages that actors cannot afford restraint while competitors advance.
Uncertainty about AI’s potential also amplifies this: actors race based on potential upside or fear of loss rather than strict expected value. Consistent with Prospect Theory (Kahneman and Tversky, 1979), actors facing the ’sure loss’ of strategic obsolescence will gamble on risky acceleration rather than accept a certain disadvantage.
6 Implications and Falsifiability
Having established that both productivity competition and AI development satisfy the five conditions for structural tragedy, we now discuss implications, falsifiability, and the crucial distinction between structural pressure and metaphysical necessity.
6.1 What the Framework Predicts
Our framework generates strong predictions testable against empirical evidence:
For Productivity:
  1. 1.
    Working hours will not decline substantially in competitive market economies absent structural intervention
  2. 2.
    Productivity gains will primarily flow to output expansion rather than welfare improvements
  3. 3.
    Attempts to coordinate work reduction through voluntary agreements will fail
  4. 4.
    National regulations reducing hours will face competitive pressure and erosion
  5. 5.
    Cultural shifts favoring leisure will be insufficient absent structural change
For AI:
  1. 1.
    Voluntary pauses on capabilities development will fail to materialize or will collapse quickly
  2. 2.
    Safety commitments will erode under competitive pressure as capabilities approach advanced levels
  3. 3.
    Racing dynamics will intensify as capabilities approach human-level performance
  4. 4.
    National regulations alone will be insufficient and will face circumvention
  5. 5.
    International coordination attempts will face structural barriers absent enforcement mechanisms
Falsifiability:
The framework is falsified if we observe:
  • Sustained and verifiable voluntary pause on AI capabilities development lasting 12+ months with major actors participating
  • Verifiable global cap on compute or capabilities enforced without defection
  • Coordination achieving substantial progress on any of the five conditions without new enforcement mechanisms
  • Removal of any one of the five conditions through policy or institutional change
  • Working hours declining substantially despite continued competition
These predictions are specific enough to be empirically testable. The framework does not claim coordination is metaphysically impossible but only structurally difficult. Evidence of successful coordination that overcomes structural barriers would falsify our analysis.
6.2 Structural Pressure vs. Metaphysical Necessity
This is not a deterministic claim about inevitable catastrophe. Rather, it identifies structural pressures creating strong tendencies toward particular outcomes. A ball will roll downhill unless an opposing force is applied.
What’s Required for Escape:
Successfully escaping the tragedy requires eliminating at least one of the five conditions:
Option 1—Reduce N: Consolidate actors to enable coordination. For productivity: reduce number of competing firms (faces antitrust concerns). For AI: consolidate development in single regulated entity (faces feasibility and risk concerns).
Option 2—Eliminate Externalities: Make individual choices not impose negative externalities on others. For both domains, this appears structurally impossible—competition inherently involves externalities.
Option 3—Break Dominance: Reduce existential stakes making defection dominant. For productivity: social safety nets, basic income, reduced competition, limit on capital mobility. For AI: reduce economic and military stakes (unclear how). This is perhaps most feasible.
Option 4—Accept Equilibrium: Give up on coordination, accepting mutual defection as outcome. For productivity: accept that leisure won’t increase. For AI: accept development risks as unavoidable.
Option 5—Solve Enforcement: Create effective monitoring and enforcement mechanisms. For productivity: global labor standards with verification and sanctions. For AI: international monitoring of compute and capabilities with enforcement. This is theoretically possible but practically difficult given international anarchy.
This is not pessimism but structural analysis. Understanding the structure is prerequisite for changing it. Interventions that ignore structural conditions will fail; interventions that address structural conditions have prospects for success.
6.3 Scope and Limitations
This framework identifies structural conditions that create coordination difficulties, but we do not claim to provide comprehensive causal explanations for all historical variation in productivity-welfare relationships or coordination outcomes across domains.
Temporal variation: The intensification of the tragedy after 1980 coincides with globalization, union decline, and increased capital mobility strengthening the conditions our framework identifies. However, detailed historical analysis is needed to establish definitive causal mechanisms for these temporal shifts.
Cross-national variation: Different institutional arrangements produce different manifestations of the productivity tragedy. Some countries exhibit the tragedy through stagnant wages despite productivity growth, others through stable hours but higher unemployment, and still others through pressure on benefits and job security. Our framework predicts that productivity gains will not flow fully to workers when competitive conditions are strong, but does not predict which specific dimension will bear the adjustment cost in each institutional context. Analysis of these institutional details and their interaction with competitive pressure is beyond our scope.
Condition intensity: As we have shown with the tragedy intensity extension, real-world conditions vary in intensity. When conditions hold weakly rather than strongly, partial coordination may be more feasible and sustainable. For instance, European labor market coordination succeeds in maintaining lower working hours than the United States, suggesting that weaker global competitive pressure (in non-tradable sectors) or stronger institutions can attenuate (though not eliminate) the tragedy. Future research should investigate systematically how condition intensity relates to the degree of coordination difficulty and the sustainability of partial coordination arrangements.
Empirical validation: We provide initial empirical evidence through declining labor share of income, productivity-wage decoupling, the European case for productivity, and the Russia-Ukraine drone war for AI development. However, comprehensive testing across historical periods, countries, sectors, and coordination attempts is necessary to fully validate our framework and identify boundary conditions. We view this paper as establishing a diagnostic framework and providing suggestive evidence, with systematic empirical validation remaining an important direction for future research.
7 Conclusion
7.1 Summary of Contributions
We have established three contributions:
First, we synthesized a unified diagnostic framework from established game theory (the social dilemma) by identifying five necessary and sufficient conditions. Second, we applied this framework to productivity, providing a formal game-theoretic model (building on (Schnaiberg, 1980)) that proves the coordination failure compelling firms to expand output is a structural tragedy, explaining Keynes’s failed prediction. The European case shows the tragedy is not escaped but merely redistributed at high cost. Third, we applied this framework to AI, demonstrating it faces the same structure but with amplified intensity. Our systematic comparison showed AI governance is structurally more difficult than historical arms control. The drone war (on the way towards full autonomy) validates this.
7.2 Theoretical and Practical Significance
This structural framework enables several advances:
Prediction and Comparison: The framework enables systematic prediction of coordination difficulty and comparison across domains.
Intervention Design: Understanding which conditions create tragedy clarifies intervention targets. Only interventions targeting the five structural conditions have prospects for success.
Falsifiability: The framework generates testable predictions about which coordination attempts will succeed or fail based on structural conditions.
Clear Diagnosis: The framework explains why standard solutions fail and provides realistic assessment of coordination difficulty.
7.3 Future Research
This framework enables several research directions:
Identification of new tragedies: By checking the five conditions, can new coordination failures be identified in new domains?
Intensity measurement: What is the best way to quantify intensity? Is our extension a good starting point? What is the best formula to compute it and what does the best mean? How does intensity help in predictions of coordination difficulty?
Dynamic Modeling: How do these systems evolve? Does a tragedy in one domain (Productivity) accelerate the tragedy in another (AI) by providing the resources for the arms race?
Comparative governance analysis: Would a systematic comparison of governance attempts across domains reveal which structural conditions most resist coordination and which interventions most effectively target them?
Historical validation: Could an analysis of past coordination successes and failures through the framework reveal whether structural changes preceded successful coordination?