### v3

# Ali Dasdan KD Consulting Saratoga, CA, USA

Abstract
In 1930, John Maynard Keynes predicted that technological progress would reduce the working week to 15 hours by century’s end. Despite productivity increasing eightfold since then, workers globally still work roughly double these hours. Separately, leading AI researchers warn of existential risks from artificial intelligence, yet development accelerates rather than slows. We demonstrate these failures share identical game-theoretic structure: coordination failures arising when individual rationality produces collective harm.
We synthesize from established game theory five necessary and sufficient conditions that characterize such coordination failures as structural tragedies: N-player structure with multiple independent actors, binary choices generating negative externalities, dominance property where defection yields higher payoffs, Pareto-inefficiency where cooperation dominates mutual defection, and enforcement difficulty from structural barriers. We extend this framework to quantify tragedy severity through condition intensities, introducing a Tragedy Index that reveals governance of transformative AI breakthroughs faces orders-of-magnitude greater coordination difficulty than historical cases including climate change, nuclear weapons, or bank runs. We validate this framework across canonical cases including commons tragedies, security dilemmas, bank runs, and space debris.
Applied to productivity competition, we prove firms face a coordination failure preventing productivity gains from translating fully to worker welfare. We show this formally for working hours; the same structural logic extends to wages, benefits, and job security. European evidence confirms that, even under maximally favorable institutional conditions, productivity welfare decoupling persists, manifesting differently across institutional contexts.
Applied to AI governance, we show development faces the same structure but with amplified intensity across eight dimensions compared to successful arms control cases, making coordination structurally more difficult than for nuclear weapons. The Russia-Ukraine drone war provides empirical validation: despite international dialogue on weapons governance, both sides escalated from dozens to thousands of drones monthly within two years, driven by competitive necessity that overwhelmed coordination aspirations.
The analysis is diagnostic rather than prescriptive: we synthesize from the literature structural barriers both to standard solutions and to the fundamental changes that would be required to resolve these tragedies, rather than proposing implementable solutions whose feasibility our framework questions.
1 Introduction
1.1 The Productivity Puzzle
John Maynard Keynes famously predicted in 1930 that technological progress would reduce the working week to 15 hours by century’s end [^1] While Keynes’s phrasing was speculative (“Three-hour shifts or a fifteen-hour week may put off the problem”), the literature has standardly interpreted this as a prediction of future working hours (e.g., (Crafts, 2022)). (Keynes, 1930). Nearly a century later, this prediction appears dramatically wrong. Workers globally typically work 1,350-2,200+ hours annually (OECD, 2025), far exceeding Keynes’s predicted 780 hours, and in many sectors working hours have increased rather than decreased (Schor, 1992); this competitive escalation is exemplified by the ‘996’ work culture (9 am to 9 pm, 6 days a week) in China (Wikipedia contributors, 2025), which has even been emulated by some US startups seeking to match that intensity (The New York Times, 2025).
This failure demands explanation. The technological capacity for reduced working hours clearly exists: productivity per worker-hour has increased approximately eightfold since Keynes wrote (Gordon, 2016). The material abundance Keynes envisioned has been achieved: modern economies produce far more goods and services per hour of labor than was possible in 1930. Yet this capacity for leisure remains unrealized.
Why? Conventional explanations typically invoke cultural factors or policy failures. We argue these explanations are incomplete. The persistence of long working hours despite dramatic productivity gains reflects not mere policy failure or cultural preferences, but a structural coordination problem analogous to tragedies in other domains.
This study is also motivated by a persistent empirical pattern observed by the author in the software and technology sectors over the last two decades. In these industries, leadership consistently demands output expansion following productivity gains, often reducing engineering headcount rather than stabilizing it. This ‘do more with less’ dynamic persists even when cost-cutting threatens organizational sustainability.
1.2 The AI Governance Puzzle
A parallel puzzle emerges in artificial intelligence development. Leading AI researchers consistently warn of existential risks from advanced AI systems (Bostrom, 2014; Russell et al., 2015; Amodei et al., 2016). Survey evidence shows substantial minorities of AI researchers estimate non-trivial probabilities of extremely bad outcomes, including human extinction (Bengio et al., 2025; Grace et al., 2018). Major AI laboratories publicly commit to safety and responsible development (Brundage et al., 2018).
Yet development accelerates rather than slows. Investment in AI capabilities races ahead (Federal Reserve Board, 2025). Voluntary pauses fail to materialize. Safety measures, while present, appear insufficient relative to the risks researchers identify (Askell et al., 2019). International coordination remains nascent and non-binding.
Why can’t we coordinate even when the stakes appear existential? Why do warnings from the field’s leading researchers fail to slow development? Why do commitments to safety coexist with accelerating capabilities races?
1.3 The Thesis: Structural Identity
We argue these puzzles exhibit the same mathematical structure. Both represent structural tragedies, coordination failures where individually rational choices produce collectively suboptimal outcomes and where coordination faces overwhelming structural barriers.
This structural identity has important implications. First, it explains why standard solutions fail in both domains: the barriers are structural, not merely political or cultural. Second, it enables prediction: we can anticipate which coordination attempts will succeed or fail based on whether they address the underlying structure. Third, it clarifies what kinds of interventions might succeed: only those that fundamentally alter the competitive architecture itself.
1.4 Preview of Contributions
This paper makes four contributions:
First, we synthesize a unified framework from established game theory (Dawes, 1980; Jervis, 1978) identifying five necessary and sufficient conditions that create structural tragedies. These conditions provide a diagnostic tool applicable across domains. We validate this framework across canonical cases including environmental commons (Hardin, 1968), security dilemmas (Mearsheimer, 2001), bank runs (Diamond and Dybvig, 1983), and other established coordination failures.
Second, we extend this framework to quantify tragedy severity through condition intensities. While the framework treats conditions as binary (present or absent) for analytical clarity, each condition varies in intensity in real-world applications. We introduce a Tragedy Index that systematically compares coordination difficulty across domains, revealing that governance of transformative AI breakthroughs faces orders-of-magnitude greater challenges than historical cases including climate change, nuclear weapons, or bank runs.
Third, we apply this framework to productivity competition and show that it satisfies all five conditions, explaining Keynes’s failed prediction structurally rather than culturally. We formalize the coordination problem firms face when productivity increases, showing why welfare improvements remain difficult despite technological capacity. European evidence shows that, even under maximally favorable institutional conditions such as wealthy democracies, strong institutions, and supranational governance, coordination achieves only partial, costly overrides of the tragic logic that erode over time.
Fourth, we demonstrate that AI development exhibits the same structure but with amplified intensity. Systematic comparison with historical arms control reveals that AI faces unfavorable conditions on every dimension where nuclear, biological, chemical (NBC) weapons coordination succeeded: dual-use inseparability, verification impossibility, low entry barriers, absence of substitutes, and convergence of economic and military imperatives. The Russia-Ukraine drone war provides empirical validation, showing how rapidly competitive dynamics override governance intentions when existential stakes are involved.
1.5 Organization
The remainder of this paper is organized as follows. Section 2 reviews the related literature on social dilemmas, productivity, and AI safety. Section 3 introduces our unified framework, defining the five necessary conditions for structural tragedy, proving these conditions are necessary and sufficient, and introducing the Tragedy Index. Section 4 applies this framework to labor markets, providing a formal proof of the coordination failure in working hours and examining the European case as empirical validation. Section 5 extends the analysis to AI governance, offering a comparative analysis with historical arms control and examining the Russia-Ukraine drone conflict. Section 6 discusses testable predictions, falsifiability, and potential interventions. Section 7 concludes.
2 Related Literature
Our work synthesizes insights from multiple, often separate, literatures.
2.1 The Prisoner’s Dilemma and Social Dilemmas
The prisoner’s dilemma, introduced by Tucker and formalized by Rapoport and Chammah (1965), represents the canonical coordination failure. Two players each face a choice between cooperation and defection. Defection dominates cooperation regardless of the other’s choice, yet mutual cooperation Pareto-dominates mutual defection. The unique Nash equilibrium—mutual defection—is recognizably suboptimal.
Schelling (1960, 1973) extended this analysis to N-player settings with binary choices and externalities. When multiple actors [^2] We use ‘player’ and ‘actor’ interchangeably for game participants. face similar decisions where individual choices impose externalities on others, coordination failures can emerge even when all actors understand the collective problem. The key insight is that as N increases, the individual incentive to deviate grows stronger—each actor’s restraint provides smaller individual benefit while the temptation to defect remains constant.
Dawes (1980) formalized the structure of “social dilemmas,” defining them as N-person games where individual rationality leads to collective irrationality. His key contribution was identifying the mathematical structure common to diverse coordination failures: defection dominates cooperation for each individual, yet universal cooperation yields higher payoffs than universal defection. Dawes emphasized that social dilemmas persist even when actors fully understand the collective harm, because rational individual choice remains incompatible with collective welfare absent enforcement mechanisms.
Jervis (1978) applied this logic to international relations, analyzing the “security dilemma” (originally coined by Herz (1950)) where states’ efforts to increase their own security—through arms buildups or territorial expansion—threaten other states and trigger countermeasures that leave all parties less secure. The security dilemma exhibits the core social dilemma structure: each state has incentive to arm regardless of others’ choices (dominance), yet mutual restraint would benefit all (Pareto-efficiency), but international anarchy prevents enforceable agreements. Jervis demonstrated that this tragic dynamic persists even when all states prefer peace, because the anarchic structure of international relations makes credible commitment to cooperation impossible.
The folk theorem literature (Fudenberg and Maskin, 1986) proves that repeated interaction can sustain cooperation under certain conditions through trigger strategies. However, these results require strong assumptions: players must be sufficiently patient, the game must repeat infinitely, and actions must be perfectly observable. When these conditions fail—as we show they do in productivity competition and AI development—cooperation becomes unsustainable even with repetition.
This foundational literature, particularly the work of Dawes (1980) on social dilemmas and Jervis (1978) on the security dilemma, establishes the core game-theoretic structure that this paper adopts. We synthesize these insights into a diagnostic framework specifying exactly which five conditions create coordination failures that resist standard solutions including repeated interaction, reputation, and institutional design. This enables systematic comparison across domains and reveals mathematical identity where previous work identified domain-specific phenomena.
2.2 Tragedies in Specific Domains
Multiple literatures have identified coordination failures in specific domains.
Hardin (1968) demonstrated how individual rationality leads to resource depletion when property rights are absent. Ostrom (1990) showed that under certain conditions—small groups, clear boundaries, monitoring capacity, graduated sanctions—communities can overcome commons tragedies through institutional design. However, her analysis also revealed that many commons lack these favorable conditions, making coordination structurally difficult.
Jervis (1978) formalized the security dilemma where states’ defensive preparations threaten other states, generating arms races that make all parties less secure. Mearsheimer (2001) extended this into offensive realism, arguing that international anarchy creates tragic dynamics where security competition becomes structurally difficult to avoid even when all states prefer peace. The key insight is that uncertainty about intentions, combined with the anarchic structure of international relations, makes cooperation on security matters extremely difficult.
Olson (1965) analyzed collective action problems in public goods provision, showing why rational individuals undercontribute to group benefits. Diamond and Dybvig (1983) modeled bank runs as coordination failures where individual rationality leads to systemic crisis. These analyses confirm that coordination failures pervade economic life, not merely environmental or security domains.
Our framework makes the underlying structure explicit, treating these as variations of the same underlying social dilemma.
2.3 Productivity and Working Hours
The literature on working hours exhibits puzzlement about why productivity gains don’t translate into leisure (Keynes, 1930).
Historical research documents the trajectory of working hours. Huberman and Minns (2007) show that annual working hours in industrialized economies declined from over 3,000 hours in 1870 to approximately 1,800 hours by 2000, but this decline has essentially stalled since the 1970s despite continued productivity growth. Pencavel (2018) documents the relationship between productivity and hours, finding that after initial declines, working hours have remained stubbornly stable.
Theoretical explanations focus primarily on preferences and culture. Becker (1965) developed time allocation theory showing how rising wages could increase or decrease labor supply depending on income and substitution effects. Schnaiberg (1980) introduced the “treadmill of production” concept, arguing that competitive pressures force continuous expansion. Alesina et al. (2005) examine US-Europe differences in working hours and conclude that European labor market regulations, advocated by unions in declining industries who argued “work less, work all,” explain the bulk of the difference, rather than cultural preferences or tax structures alone.
Recent work by Acemoglu and Restrepo (2020) and Acemoglu (2025) provides important insights into technology adoption, distinguishing between automation that displaces labor (“so-so automation”) and technology that creates new tasks or augments workers (“human-augmenting”). They argue that market distortions lead to socially excessive investment in displacement.
While these theories explain specific drivers—Schnaiberg (1980) on competitive intuition and Acemoglu and Restrepo (2020); Acemoglu (2025) on technology choice—the literature lacks a unified game-theoretic explanation for the coordination problem in allocation. We distinguish between which technologies firms adopt (Acemoglu’s focus) and how productivity gains are utilized (our focus). Also, while extensive research documents the secular decline in the labor share of income globally (Karabarbounis and Neiman, 2014; Elsby et al., 2013), a formal model explaining why competitive pressure systematically prevents even human-augmenting productivity gains from translating to worker welfare remains absent.
We provide the first formal game-theoretic model of productivity competition as an N-player social dilemma. Building on Schnaiberg (1980)’s treadmill intuition, we prove that firms face a structural coordination failure: each firm has a dominant strategy to expand output rather than reduce hours, yet universal restraint would Pareto-dominate universal expansion. We formalize the barriers to coordination that make this tragedy persist. This structural framework explains Schnaiberg (1980)’s empirical observation, complements Acemoglu and Restrepo (2020); Acemoglu (2025)’s analysis of adoption failures with a theory of allocation failures, and clarifies why the regulatory interventions documented by Alesina et al. (2005) require costly, persistent institutional support to partially override market dynamics.
2.4 AI Governance and Safety
The AI safety literature has grown rapidly in recent years, identifying multiple challenges in developing safe and beneficial AI systems.
Bostrom (2014) provided a comprehensive analysis of existential risks from superintelligent AI, emphasizing the orthogonality thesis (intelligence and goals are independent) and instrumental convergence (advanced AI systems will pursue convergent subgoals regardless of final objectives). Russell (2019) framed AI safety as an inverse reinforcement learning problem and proposed provably beneficial AI through value alignment. Amodei et al. (2016) identified concrete problems in AI safety: avoiding negative side effects, reward hacking, distributional shift, safe exploration, and robustness to distributional change. Dafoe (2018) analyzed AI governance challenges, identifying the need for international coordination on development and deployment. Bengio et al. (2025) provided a comprehensive synthesis of the current evidence on the capabilities, risks, and safety of advanced AI systems. Panigrahy and Sharan (2025) proved that under reasonable mathematical definitions of safety, trust, and artificial general intelligence (AGI), a safe and trusted AI system cannot be an AGI system. Armstrong et al. (2016) analyzed “racing to the precipice”, the concern that competitive pressure could drive unsafe AI development. Their analysis identifies the race dynamic but does not provide a formal structural model explaining why coordination fails.
While this literature thoroughly documents AI risks and governance challenges, it lacks a formal structural analysis of why coordination fails. What are the game-theoretic conditions that make AI governance difficult? Why do commitments to safety coexist with accelerating development? How does AI compare structurally to previous coordination challenges like NBC weapons?
3 A Unified Framework for Structural Tragedies
We now present a unified framework identifying necessary and sufficient conditions for structural tragedies. We formalize these conditions, prove they are necessary and sufficient to generate coordination failures, and validate the framework across canonical cases.
We use the term “Tragedy” in the strict sense defined by Whitehead (1925) and popularized by Hardin (1968): “the solemnity of the remorseless working of things.” A situation becomes a tragedy not through error or malice, but through inevitability. Specifically, it occurs when individual rationality compels defection even though actors collectively recognize cooperation as superior, while structural barriers prevent coordination. The tragedy lies in this structural trap: actors are not making mistakes—they are imprisoned by the logic of their own survival.
We formalize the payoff structure as follows. Let $s=(s_{1},s_{2},\ldots,s_{n})$ be a strategy profile for nn=N players, which is a vector specifying the chosen strategy $s_{i}\in\{C,D\}$ for every player $i$, where $C$ represents Cooperate (restraint) and $D$ represents Defect (advantage-seeking).
The utility (or payoff) for a specific player $i$ is given by a function $U_{i}$ that maps this complete strategy profile to a real-valued payoff:
$U_{i}(s)=U_{i}(s_{1},s_{2},\ldots,s_{n})$
For clarity, we will often separate the strategy of player $i$ from the strategies of all other players. We denote the strategy profile of all players except $i$ as $s_{-i}$. This allows us to write the utility function as $U_{i}(s_{i},s_{-i})$.
Using this formal notation, we can precisely define the five conditions for the tragedy:
3.1 The Five Conditions
A problem exhibits structural tragedy if and only if it satisfies the following five conditions:
Definition 1 (C1: N-Player Structure).
Multiple independent decision-makers ($n\geq 2$) each control their own choices. No single player can solve the problem unilaterally. Choices are made in a decentralized architecture without centralized coordination.
This condition specifies the basic game structure. If a single player controlled all decisions, coordination would be trivial. The tragedy emerges from decentralized choice architecture where multiple players must independently decide their strategies.
Definition 2 (C2: Binary Choice with Externalities).
Each player $i$ faces a binary choice between two strategies: Cooperate ($C$) representing restraint, or Defect ($D$) representing advantage-seeking. Individual choices create externalities affecting all players. Specifically, one player’s defection imposes negative externalities on others.
The binary structure simplifies the game while capturing the essential strategic tension. In practice, choices may appear continuous, but the strategic logic often reduces to a binary decision: cooperate or defect. The key is that individual choices impose externalities: one player’s defection makes cooperation more costly for others.
Definition 3 (C3: Dominance Property).
Defection yields higher payoff than cooperation regardless of other players’ choices. Formally,
$U_{i}(D,s_{-i})>U_{i}(C,s_{-i})\quad\text{for any possible }s_{-i}$
Dominance creates the dilemma. Each player has strict incentive to defect regardless of what others do. This is often driven by existential stakes. In high-stakes tragedies, survival instincts override social norms; if cooperating while others defect leads to elimination (e.g., death, bankruptcy), defection becomes an absolute imperative. When survival is at stake, the incentive to defect becomes overwhelming.
Definition 4 (C4: Pareto-Inefficiency).
Universal cooperation Pareto-dominates universal defection. Formally, let $s^{C}=(C,C,\ldots,C)$ be the Universal Cooperation profile and let $s^{D}=(D,D,\ldots,D)$ be the Universal Defection profile. This condition states that every player strictly prefers the universal cooperation outcome:
$U_{i}(s^{C})>U_{i}(s^{D})\quad\text{for all players }i.$
The equilibrium outcome is recognizably suboptimal: all players would prefer mutual cooperation to mutual defection if they could coordinate.
This condition ensures the tragedy is recognizable. Players understand that mutual defection makes everyone worse off than mutual cooperation. The tragedy is not hidden; it is visible, predictable, and collectively unwanted. Yet the structure makes it difficult to escape.
Definition 5 (C5: Enforcement Difficulty).
Structural barriers prevent effective coordination mechanisms. These barriers may include: (i) anarchy as in no central authority with enforcement power, (ii) verification impossibility as in actions cannot be reliably monitored, (iii) monitoring infeasibility as in costs of monitoring exceed benefits of coordination, or (iv) commitment problems as in players cannot credibly commit to future cooperation.
This condition distinguishes structural tragedies from simple coordination problems. Many Pareto-inefficient equilibria can be overcome through contracts, monitoring, or institutional design. Structural tragedies persist because enforcement mechanisms face fundamental barriers: not merely high costs but severe structural obstacles.
3.2 The Formal Result
We now prove these five conditions are both necessary and sufficient to generate structural tragedy.
Theorem 6 (Necessity and Sufficiency of Tragedy Conditions).
Consider an N-player game satisfying Conditions 1-2 (structure and externalities). Then:
  1. (i)
    Universal defection $s^{D}$ is the unique Nash equilibrium.
  2. (ii)
    This equilibrium is stable against unilateral deviations.
  3. (iii)
    This equilibrium is Pareto-dominated by universal cooperation $s^{C}$.
  4. (iv)
    This equilibrium persists absent radical structural change.
if and only if Conditions 3-5 (dominance, Pareto-inefficiency, and enforcement difficulty) also hold.
Proof.
We establish sufficiency and necessity separately, followed by the overall conclusion.
Part I: Sufficiency.
Assume Conditions 1-5 hold. We prove claims (i)-(iv).
Claim (i): Unique Nash Equilibrium.
Let $s=(s_{1},\ldots,s_{n})$ be a strategy profile where $s_{i}\in\{C,D\}$ for each player $i$.
First, we show $s^{D}=(D,\ldots,D)$ is a Nash equilibrium. For any player $i$, their payoff from defecting is $U_{i}(D,s^{D}_{-i})$, where $s^{D}_{-i}$ is the profile where all other players also defect. If player $i$ unilaterally deviates to cooperation, their payoff becomes $U_{i}(C,s^{D}_{-i})$. By the dominance property (Condition 3), $U_{i}(D,s^{D}_{-i})>U_{i}(C,s^{D}_{-i})$. Therefore, no player has an incentive to deviate, and $s^{D}$ is a Nash equilibrium.
Second, we show no other profile is a Nash equilibrium. Consider any profile $s$ where at least one player $j$ cooperates. This player’s payoff is $U_{j}(C,s_{-j})$, where $s_{-j}$ is the profile of all other players. If player $j$ deviates to defection, their payoff becomes $U_{j}(D,s_{-j})$. By the dominance property (Condition 3), $U_{j}(D,s_{-j})>U_{j}(C,s_{-j})$. Therefore, player $j$ has a strict incentive to deviate. Hence, no profile with any cooperation is a Nash equilibrium.
Therefore, $s^{D}$ is the unique Nash equilibrium.
Claim (ii): Stability.
Stability follows immediately from Claim (i). At the equilibrium $s^{D}$, no player has unilateral incentive to deviate by the dominance property. The equilibrium is thus stable against unilateral deviations.
Claim (iii): Pareto-Domination.
By Condition 4 (Pareto-inefficiency), $U_{i}(s^{C})>U_{i}(s^{D})$ for all players $i$. Therefore, universal cooperation strictly Pareto-dominates the Nash equilibrium.
Claim (iv): Persistence.
Even though all players recognize that $s^{C}$ Pareto-dominates $s^{D}$, they cannot coordinate to achieve it. By Condition 5, structural barriers prevent enforcement:
  • If anarchy prevails, no authority can compel cooperation
  • If verification is impossible, players cannot monitor compliance
  • If monitoring is infeasible, detection of defection comes too late
  • If commitment is problematic, promises to cooperate lack credibility
Without enforcement, any agreement to cooperate is unstable, as each player has a strict incentive to defect (from Claim (i)). The folk theorem, which allows for cooperation in repeated games, does not apply because its required conditions (perfect monitoring, sufficient patience, infinite repetition) fail.
Therefore the tragedy persists absent radical structural change.
Part II: Necessity.
We prove necessity by showing the contrapositive: if any of Conditions 1-5 fails to hold, then the tragedy structure breaks (at least one of claims (i)-(iv) fails). By contrapositive logic, this establishes that each condition is necessary for tragedy.
[…]
If there is a single decision-maker ($n=1$) or a central authority with binding power over all players, then there is no coordination problem. The single player or central authority simply implements the Pareto-optimal outcome $s^{C}$ directly. No strategic interaction exists, and no tragedy arises. Therefore C1 is necessary.
[…]
If individual choices create no externalities, each player’s payoff depends only on their own choice, not others’ choices: $U_{i}(s_{i},s_{-i})=U_{i}(s_{i})$. Each player independently optimizes their own choice without strategic consideration of others’ actions. There is no coordination problem because choices are independent. Strategic interaction disappears, eliminating the coordination failure. Therefore C2 is necessary.
[…]
If defection is not strictly dominant, then there exist some player $i$ and some opponent profile $s_{-i}$ such that $U_{i}(C,s_{-i})\geq U_{i}(D,s_{-i})$. In such cases, cooperation can be a best response to others cooperating. This creates the possibility of multiple Nash equilibria, including $s^{C}$ as a stable equilibrium. The tragedy is no longer structurally inevitable. Therefore C3 is necessary.
[…]
If the equilibrium $s^{D}$ is not Pareto-dominated by $s^{C}$, then for some player $i$ we have $U_{i}(s^{D})\geq U_{i}(s^{C})$, or universal defection is Pareto-optimal. Then the equilibrium outcome is efficient by definition: there exists no alternative outcome that makes everyone better off. Without Pareto-domination, there is no tragedy. The outcome may be a coordination problem, but it is not recognizably suboptimal. Therefore C4 is necessary.
[…]
Suppose structural barriers to enforcement are absent. Specifically, suppose any of the following holds:
  • A central authority exists with enforcement power, or
  • Actions are sufficiently observable that defection can be reliably detected, or
  • Monitoring is feasible at reasonable cost relative to coordination benefits, or
  • Players can make credible commitments to future cooperation
Under these conditions, the folk theorem for repeated games applies: cooperation can be sustained as a subgame perfect equilibrium through trigger strategies because defection is detectable and punishable. For the structural tragedy to hold, C5 must therefore imply barriers sufficient to preclude these solutions: specifically, monitoring must be too imperfect to detect defection, or patience too low for future consequences to matter. Absent such barriers, the problem dissolves from a structural tragedy to a solvable coordination challenge. Therefore C5 is necessary.
Conclusion.
[…]
This theorem formally distinguishes structural tragedies from solvable coordination problems. By establishing that the tragedy persists if and only if all five conditions hold, the result confirms that the equilibrium is structurally resilient against standard interventions. This provides precise diagnostic utility: any proposed solution that fails to break at least one condition is mathematically guaranteed to fail, while effective governance requires targeting the specific structural barriers identified by the framework.
3.3 Validation Across Canonical Cases
We now validate this framework by applying it to established coordination failures across multiple domains. Table 1 summarizes the analysis. We briefly discuss each case, demonstrating how it satisfies all five conditions.
Table 1: Five Conditions C1-C5 Across Canonical Coordination Failures, including the two new ones introduced in this paper.
Domain C1 C2 C3 C4 C5 Reference
Commons Hardin (1968)
Security Jervis (1978); Mearsheimer (2001)
Bank Runs Diamond and Dybvig (1983)
Space Debris Adilov et al. (2015)
Antibiotics Foster and Grundmann (2006)
Public Goods Olson (1965)
Hockey Helmets Schelling (1973)
Climate Change Ostrom (2010)
Productivity this paper
AI Governance this paper
Commons Tragedies
(Hardin, 1968): Multiple shepherds (N-player) each decide whether to add sheep to common pasture (binary choice creating negative externality through overgrazing). Each shepherd benefits from adding sheep regardless of others’ choices (dominance driven by individual profit maximization). Yet universal restraint would preserve the commons for all (Pareto-domination). Absent property rights or enforceable agreements (enforcement difficulty), the commons is depleted.
Security Dilemmas
(Jervis, 1978; Mearsheimer, 2001): Multiple states (N-player) each decide whether to arm or remain unarmed (binary choice where one state’s arms create security threats for others). Each state benefits from arming for security regardless of others’ choices (dominance driven by survival concerns in anarchy). Yet universal disarmament would provide security without arms costs (Pareto-domination). International anarchy prevents enforceable disarmament agreements (enforcement difficulty).
Bank Runs
(Diamond and Dybvig, 1983): Multiple depositors (N-player) each decide whether to withdraw or keep deposits in a bank (binary choice where withdrawals create panic). Each depositor benefits from withdrawing if others might withdraw (dominance driven by fear of losing deposits). Yet if no one panics, all maintain deposits safely (Pareto-domination). Absence of instant coordination during panics (enforcement difficulty) allows runs to occur. Deposit insurance (FDIC) broke this tragedy (for deposits up to the insurance limit) by eliminating existential stakes, removing Condition 3. Note that the wholesale funding and equities markets are not covered under the FDIC insurance so they may still be vulnerable to bank runs.
Space Debris
(Adilov et al., 2015): Multiple states and firms (N-player) each decide whether to limit launches or expand operations (binary choice where debris from launches threatens all satellites). Each actor benefits from launches regardless of others’ choices (dominance driven by commercial and security advantages). Yet universal restraint would preserve orbital space (Pareto-domination). International space law lacks enforcement mechanisms (enforcement difficulty).
Antibiotic Resistance
(Foster and Grundmann, 2006): Multiple actors including doctors, patients, farmers (N-player) each decide whether to limit or freely use antibiotics (binary choice where overuse breeds resistance affecting all). Each actor benefits from use regardless of others’ choices (dominance driven by immediate health or productivity gains). Yet universal restraint would preserve antibiotic effectiveness (Pareto-domination). Distributed usage across multiple sectors resists coordination (enforcement difficulty).
Public Goods
(Olson, 1965): Multiple individuals (N-player) each decide whether to contribute to a public good (Cooperate) or withhold contribution (Defect). Withholding allows an individual to save resources (free-ride) while still enjoying any good provided by others (externality). Each individual has a dominant incentive to free-ride, as their personal contribution is costly and has minimal impact on the total outcome (dominance). Yet, universal contribution would fund the public good, making everyone better off than if the good is not provided (Pareto-domination). Absent a coercive mechanism like taxation (enforcement difficulty), the public good is under-provided or not provided at all.
Hockey Helmets
(Schelling, 1973): Multiple players (N-player) each decide whether to wear a helmet (Cooperate/Restrain) or play without one (Defect/Advantage-seeking). Playing without a helmet provides a slight competitive advantage (e.g., better vision, ‘toughness’) but increases the overall risk of injury for all (negative externality). Each player benefits from the competitive edge of not wearing one, regardless of others’ choices (dominance). Yet, universal helmet use would neutralize the competitive effects and make everyone safer (Pareto-domination). Absent a league-wide, binding mandate (enforcement difficulty), players are incentivized to play without helmets despite the collective danger.
Climate Change
(Ostrom, 2010): Multiple nations and firms (N-player) each decide whether to reduce emissions (Cooperate) or pollute freely (Defect). Polluting provides immediate economic advantages while imposing global climate damage (negative externality). Each actor benefits from polluting to maintain economic competitiveness, regardless of others’ choices (dominance). Yet, universal restraint would prevent catastrophic climate costs, making all better off (Pareto-domination). International anarchy and sovereignty concerns prevent binding, enforceable agreements (enforcement difficulty).
Result:
Perfect verification across all canonical cases. Each tragedy satisfies the complete set of conditions. This suggests the framework captures the essential structure that creates coordination failures.
Notably, cases that do not satisfy all conditions either are not tragedies or have been successfully resolved. For example, bank runs violated Condition 3 (dominance) when deposit insurance eliminated existential stakes (up to the insurance limit) since depositors no longer benefit from withdrawing if others withdraw. This structural change broke the tragedy, validating our framework’s diagnostic power.
3.4 Implications of the Framework
This unified framework provides several analytical advantages:
Identification: We can identify new tragedies by checking whether a situation satisfies the framework. If it does, coordination will be structurally difficult. If any condition fails, standard solutions may suffice.
Prediction: We can predict difficulty of coordination by examining the intensity with which each condition is satisfied. Situations with extreme dominance (high existential stakes) or severe enforcement problems will resist coordination more strongly than situations where conditions are weakly satisfied.
Intervention Design: We can design interventions by targeting specific conditions. Breaking any single condition may resolve the tragedy.
3.5 Extension: Condition Intensity and Tragedy Severity
While the framework treats conditions as binary (present or absent) for analytical clarity, each condition varies in intensity in real-world applications. The severity of a structural tragedy increases with the intensity of each condition. This intensity dimension provides additional predictive and diagnostic power by distinguishing between manageable nuisances and existential threats.
3.5.1 C1: N-Player Structure - Number and Heterogeneity of Actors.
Intensity spectrum:
  • Low intensity: Few actors with stable identities (e.g., small groups like hockey leagues, bilateral security relationships)
  • Medium intensity: Moderate number of organized actors (e.g., dozens of states or firms in defined sectors like commons governance, public goods provision)
  • High intensity: Large number of heterogeneous actors across sectors (e.g., hundreds of nations plus countless firms in climate change, global productivity competition)
  • Extreme intensity: Massive number of individual-level decisions across decentralized populations (e.g., millions of depositors in bank runs, billions of patients and prescribers in antibiotics resistance)
Impact on tragedy severity:
More actors make coordination exponentially more difficult. Each actor’s individual contribution to the collective good diminishes as $n$ increases, while the temptation to defect remains constant. Additionally, heterogeneity among actors (different sizes, capabilities, preferences) further intensifies coordination difficulty.
3.5.2 C2: Externalities - Magnitude and Scope.
Intensity spectrum:
  • Low intensity: Small, localized negative externalities with delayed or uncertain effects (e.g., individual free-riding on public goods, minor competitive disadvantages in hockey helmets)
  • Medium intensity: Moderate externalities with gradual accumulation (e.g., commons depletion over years, antibiotic resistance developing gradually)
  • High intensity: Large, immediate externalities with rapid transmission (e.g., financial panic spreading within days, climate damages affecting all actors, AI capability advantages immediately threatening competitors)
  • Extreme intensity: Anticipatory externalities where actors respond to perceived threats before they materialize, creating self-reinforcing spirals (e.g., security dilemmas where intelligence about potential programs triggers immediate counter-programs)
Impact on tragedy severity:
Larger externalities create stronger feedback loops. When one actor’s defection severely harms others, it triggers more aggressive competitive responses, accelerating the race to the bottom. The speed and magnitude of externality transmission determines how rapidly the tragedy unfolds.
3.5.3 C3: Dominance - Strength of Existential Stakes.
Intensity spectrum:
  • Low intensity: Defection provides modest advantage; cooperation causes minor competitive disadvantage (e.g., visibility or comfort in hockey helmets, reduced returns in public goods)
  • Medium intensity: Defection provides significant competitive advantage; cooperation causes substantial market share loss but not immediate elimination (e.g., productivity competition where firms can survive with reduced hours, commons where individuals can survive with restraint, climate where nations face competitive disadvantage but can adapt)
  • High intensity: Defection necessary to avoid severe consequences; cooperation risks major losses (e.g., state conquest in security dilemmas, losing all deposits in bank runs)
  • Extreme intensity: Defection necessary to avoid permanent strategic obsolescence; restraint means irreversible loss with no possibility of recovery (e.g., transformative AI where falling behind creates permanent winner-take-all dynamics)
Impact on tragedy severity:
The intensity of existential stakes determines how strongly dominance holds. When cooperation means death or bankruptcy, actors will defect regardless of collective costs. Lower existential stakes create space for partial coordination or voluntary restraint.
3.5.4 C4: Pareto-Inefficiency - Magnitude of Collective Loss.
Intensity spectrum:
  • Low intensity: Universal cooperation slightly better than universal defection; modest welfare gains available (e.g., marginal safety improvements from universal helmet use)
  • Medium intensity: Universal cooperation moderately better; significant quality-of-life improvements but society continues functioning (e.g., commons preservation, public goods provision, productivity welfare gains, avoiding bank runs, maintaining antibiotic effectiveness)
  • High intensity: Universal cooperation vastly superior; defection leads to major catastrophic outcomes with potential for recovery (e.g., regional conflicts, economic depressions)
  • Extreme intensity: Universal cooperation necessary to avoid irreversible catastrophic outcomes with permanent losses (e.g., human extinction from nuclear war or AI misalignment, irreversible climate tipping points causing ecosystem collapse)
Impact on tragedy severity:
Greater Pareto-inefficiency means larger collective losses from coordination failure. This affects both the normative urgency (how important it is to solve) and practical difficulty (actors may invest more in solving severe tragedies). However, high inefficiency alone doesn’t make coordination easier; it only makes failure more costly.
3.5.5 C5: Enforcement Difficulty - Severity of Structural Barriers.
Intensity spectrum:
  • Low intensity: Enforcement absent or weak in current form but structurally feasible; clear authority could monitor and sanction violations with institutional development
  • Medium intensity: Enforcement difficult but possible with institutional investment; monitoring costly but achievable; commitment requires sustained effort (e.g., league rules for helmets, taxation for public goods)
  • High intensity: Enforcement very difficult due to international anarchy or distributed nature; verification problematic; commitment fragile; sanctions difficult but not impossible (e.g., commons without property rights, international labor standards, nuclear non-proliferation, climate agreements, bank run coordination)
  • Extreme intensity: Enforcement structurally impossible due to fundamental technical barriers (perfect hiding, instant copying, inseparable dual-use) combined with international anarchy and no feasible global authority (e.g., AI development where software is copyable, training concealable, capabilities unverifiable)
Impact on tragedy severity:
The impossibility of enforcement determines whether the tragedy is merely difficult or truly structural. When enforcement is difficult but possible (low to medium intensity), sufficient institutional investment can achieve coordination. When enforcement is structurally impossible (extreme intensity), no amount of effort suffices without changing the underlying structure.
Subcategories of enforcement difficulty:
  • Anarchy intensity: No authority $\rightarrow$ weak authority $\rightarrow$ strong authority
  • Verification impossibility: Perfect hiding $\rightarrow$ difficult detection $\rightarrow$ easy monitoring
  • Commitment problems: No credible commitment $\rightarrow$ costly commitment $\rightarrow$ binding commitment
3.6 Implications of Intensity Analysis
This intensity framework provides several analytical advantages:
Comparative Prediction:
We can predict relative difficulty of coordination across domains by comparing condition intensities. AI governance (for transformative breakthroughs) faces extreme intensity on three conditions simultaneously (C3, C4, C5), while severe historical cases like nuclear security dilemmas faced extreme intensity on two conditions (C2, C4) and climate change on one (C4). Current AI capabilities competition exhibits high but not extreme intensity, as evidenced by multiple viable competitors despite capability gaps.
Intervention Prioritization:
Interventions should target the highest-intensity conditions. For AI, the extreme intensity of C5 (enforcement impossibility due to verification) suggests compute governance (making verification partially possible) may be most promising despite remaining severe challenges.
Partial Coordination Explanation:
The European case shows lower intensity on some conditions (reduced existential stakes through wealthy welfare states, supranational governance reducing anarchy) enables partial coordination despite high intensity on others (global competition, verification difficulty).
Temporal Variation Explanation:
Condition intensities change over time. Productivity competition shows moderate base intensity (C2, C3 at medium due to gradual market dynamics and non-existential stakes), but globalization after 1980 increased N (more actors), strengthened externalities (faster competitive transmission), and weakened enforcement (capital mobility), making even this moderate tragedy harder to address. Climate change intensified as C4 shifted from high to extreme with growing evidence of irreversible tipping points.
Falsification Refinement:
The framework predicts that reducing condition intensity should reduce tragedy severity proportionally. Evidence of coordination success without reducing any condition’s intensity would falsify the framework. The European case confirms: partial reduction in intensity (some conditions) produces partial coordination (some dimensions).
3.7 Tragedy Index: Intensity Quantification
The intensity framework adds a continuous dimension to our binary framework, enabling more nuanced predictions while maintaining the core insight that all necessary and sufficient conditions must be satisfied (at non-zero intensity) for structural tragedy to occur.
The intensity framework also allows for two analytical advances:
  • Structural Fingerprinting: As illustrated in Figure 1, plotting intensities on a radar chart allows us to visualize the ‘shape’ of the coordination failure. This visualization reveals that AI Governance effectively ‘envelopes’ historical cases like Nuclear Security. While Nuclear Security spikes on specific dimensions (Externalities and Pareto-Inefficiency), AI Governance maintains high or extreme intensity across all five dimensions, creating a unique and formidable structural profile.
    Figure 1: The Structural Fingerprint of AI Governance. Comparison of condition intensities for four selected domains (logarithmic scale), chosen to illustrate diverse structural profiles and highlight AI Governance’s unique fingerprint. Note how AI Governance (red) envelopes other tragedies, exhibiting ‘Extreme’ intensity on three dimensions simultaneously (C3, C4, C5), whereas historical cases like Security or Climate Change spike only on specific dimensions. Overlapping areas (blended colors) indicate shared structural challenges.
  • The Tragedy Index: We propose a quantitative metric to compare severity across domains, acknowledging that coordination difficulty plausibly scales non-linearly with condition intensity.
We propose the Tragedy Index ($\iota$) as a heuristic quantification of coordination difficulty. Two methodological choices deserve acknowledgment:
Exponential Scaling:
We use base-10 exponential intensities ($I_{c}\in\{0,10,10^{2},10^{3},10^{4}\}$) because coordination difficulty plausibly scales non-linearly: a tragedy with ‘Extreme’ conditions on multiple dimensions may be orders of magnitude harder than one with ‘low’ conditions, not merely several times harder. The base-10 choice is somewhat arbitrary but enables intuitive order-of-magnitude interpretation.
Geometric Aggregation:
We combine intensities via geometric mean (the 5th root of the product):
$\iota=\left(\prod_{c=1}^{5}I_{c}\right)^{\frac{1}{5}}$
where $I_{c}$ corresponds to Solved/Absent (0), Low (10), Medium (100), High (1,000), or Extreme (10,000).
We employ the geometric mean because the five conditions are necessary: the tragedy structure requires the conjoint presence of all factors. Mathematically, necessity implies logical conjunction, which corresponds to multiplication rather than addition. If any condition intensity falls to zero, the index correctly yields zero: no tragedy exists. An arithmetic mean would incorrectly assign positive severity when a necessary condition is absent.
Additionally, the nth-root operation penalizes extreme conditions heavily: a single extreme-intensity condition (10,000) contributes less to the index than moderate intensity across multiple conditions, reflecting the disproportionate coordination difficulty when several conditions bind simultaneously. Alternative aggregations such as logarithmic sum would yield different absolute values but preserve ordinal rankings across domains.
The value of this index lies not in precise quantification but in systematic comparison. We interpret the resulting values as “Tragedy Magnitude”:
  • Magnitude 0: Solved (Stable)
  • Magnitude 10–100: Manageable Nuisance (Chronic)
  • Magnitude 100–1,000: Systemic Crisis (Acute)
  • Magnitude 1,000+: Existential/Terminal Threat
As shown in Table 2, under this heuristic measure, AI Governance ($\iota=3{,}981$) faces orders-of-magnitude greater coordination difficulty than other severe tragedies. Climate change, bank runs, and nuclear security dilemmas all face catastrophic stakes ($\iota=1{,}000$ each), but differ in their condition profiles: security achieves this through anticipatory dynamics and extinction-level risks despite few actors; climate through irreversibility; bank runs through mass-scale coordination failure. Productivity competition ($\iota=251$) and antibiotics resistance ($\iota=398$), despite their global scale and persistence, face lower intensity due to non-existential stakes and gradual dynamics. Manageable cases like public goods ($\iota=40$) have been successfully addressed through institutional solutions. This quantification reinforces the qualitative analysis: AI governance presents an unprecedented coordination challenge.
Table 2: Intensity Comparison of Conditions C1-C5 Across Domains, including the two new ones introduced in this paper. Intensities: L=Low (10), M=Medium (100), H=High (1,000), E=Extreme (10,000).a
Domain C1 C2 C3 C4 C5 Index ($\iota$)
Commons M M M M M 100
Securityb L E H E H 1,000
Bank Runs E H H M H 1,000
Space Debris M M M M H 158
Antibiotics E M M M H 398
Public Goods M L L M M 40
Hockey Helmets L L L L M 16
Climate Change H H M E H 1,000
Productivity H M M M H 251
AI Governancec H H E E E 3,981
aTable reflects intensity ratings after systematic review of condition specifications.
bSecurity dilemmas rated for nuclear-age scenarios where C2 reflects anticipatory responses to perceived threats and C4 reflects extinction-level risks.
cAI Governance ratings reflect transformative breakthroughs (AGI/superintelligence). Current AI capabilities competition would rate approximately (H, H, H, H, E) with Index $\approx$ 1,000, similar to climate change and bank runs, as multiple competitors survive capability gaps.
We now apply the basic framework (without the intensity extension) to productivity competition and AI development, demonstrating that both satisfy all necessary and sufficient conditions, and that both are structural tragedies.
3.8 Alternative Formalization: Game-Theoretic Measures
While the Tragedy Index provides intuitive comparison across domains through structural condition intensities, formal game-theoretic measures offer an alternative approach to quantifying tragedy severity. These measures focus on the payoff structure (Conditions 3 and 4) and provide mathematical precision about welfare losses and incentive misalignments.
Consider an N-player game satisfying Conditions 1-5. Let $U_{coop}$ denote global welfare under universal cooperation (profile $s^{C}$), $U_{nash}$ denote global welfare at the Nash equilibrium (profile $s^{D}$), and for symmetric games, let $T$ be the temptation payoff (defect while others cooperate), $R$ be the reward payoff (mutual cooperation), $P$ be the punishment payoff (mutual defection), and $S$ be the sucker payoff (cooperate while others defect). Structural tragedies satisfy the prisoner’s dilemma ordering: $T>R>P>S$.
The Normalized Rationality Gap:
(Nisan et al., 2007) measures the fraction of potential welfare lost due to coordination failure:
$\Delta_{norm}=\frac{U_{coop}-U_{nash}}{U_{coop}}=1-\frac{U_{nash}}{U_{coop}}$ (1)
This takes values in $[0,1]$, where $\Delta_{norm}\approx 0$ indicates minimal tragedy and $\Delta_{norm}\to 1$ indicates catastrophic welfare loss. For symmetric games, this simplifies to $\Delta_{norm}=(R-P)/R$.
The Price of Anarchy:
(Koutsoupias and Papadimitriou, 1999) measures the ratio between optimal social welfare and worst-case equilibrium welfare:
$\text{PoA}=\frac{U_{coop}}{U_{nash}}=\frac{R}{P}$ (2)
This multiplicative formulation naturally captures orders-of-magnitude differences: PoA $\approx$ 1 indicates near-efficiency, PoA $\in[10,100]$ indicates severe inefficiency, and PoA $\to\infty$ indicates existential tragedy where equilibrium approaches zero value.
The Temptation-Cooperation Ratio:
captures the tension between individual incentives and collective benefits (Rapoport and Chammah, 1965):
$\tau=\frac{T-R}{R-P}$ (3)
where $T$ is the temptation payoff (defect while others cooperate). This ratio measures how much individual gain from defection exceeds collective gain from cooperation: $\tau<1$ indicates coordination is structurally feasible, $\tau\approx 1$–5 indicates substantial coordination difficulty, and $\tau\to\infty$ indicates existential stakes make defection imperative.
These measures require detailed payoff specifications for each domain. For productivity competition, our formalization in Section 4 provides the foundation for such calculations. For AI governance, security dilemmas, and climate change, comprehensive payoff structures would require modeling uncertainty about catastrophic outcomes, discount rates, and long-term value functions—substantial undertakings beyond our current scope.
The Tragedy Index (Subsection 3.7) and game-theoretic measures serve complementary purposes: the Index enables systematic comparison across diverse domains through structural condition intensities without requiring precise payoff quantification, while game-theoretic measures provide rigorous welfare analysis when detailed payoff structures can be reliably estimated. Future research could apply these formal measures systematically across domains to validate and refine the Index’s ordinal rankings.
4 The Tragedy of Productivity
We now apply the unified framework to productivity competition, demonstrating that all five conditions are satisfied. This explains structurally why Keynes’s prediction failed: not from cultural preferences or policy failures, but from coordination barriers inherent in competitive markets.
Although worker welfare includes wages and benefits in addition to working hours, we will use working hours as a proxy for worker welfare as well as to simplify the formal analysis. Our welfare analysis focuses on firms and workers as the direct participants in labor market coordination. Productivity gains may generate broader societal effects (e.g., consumer surplus from lower prices), but these lie outside our scope; we note that workers and consumers substantially overlap, making the welfare comparison largely internal to the same population.
4.1 Mapping the Five Conditions to Firm Competition
4.1.1 C1: N-Player Structure.
Productivity competition involves multiple independent decision-makers:
  • Multiple firms compete in domestic and global markets
  • National governments face political pressure from firms and workers
  • No single actor can coordinate welfare improvements unilaterally
The decision architecture is fundamentally decentralized.
4.1.2 C2: Binary Choice with Externalities.
When productivity increases, each firm faces essentially a binary choice:
  • Strategy $C$ (Cooperate/Restrain): Reduce working hours proportionally to productivity gains while maintaining output
  • Strategy $D$ (Defect/Expand): Maintain or increase working hours to expand output and market share
Individual choices create externalities. When one firm expands output (defects), it increases competitive pressure on all rivals:
  • Expanded output captures market share from competitors
  • Increased competition reduces rivals’ profit margins
  • Rivals face pressure to expand output to avoid displacement
The externality is negative: one firm’s defection harms others by intensifying competition.
4.1.3 C3: Dominance Property.
Defection (output expansion) dominates cooperation (welfare improvements) regardless of competitors’ choices:
Case 1—If competitors reduce hours: Firm gains decisive market share advantage by maintaining output. Competitors who reduce hours produce less, creating market opportunity. The expanding firm captures this market share, increasing profits and market position.
Case 2—If competitors expand output: Firm must expand to avoid displacement. Competitors’ expansion increases supply and reduces prices. Reducing hours while competitors expand means producing less in a more competitive market—a recipe for bankruptcy.
This dominance is often driven by existential stakes. In the specific context of market competition, this is driven by the threat of insolvency. If a firm maintains high-cost working conditions while competitors expand output and lower prices, it risks losing market share to the point of bankruptcy. The survival instinct of the firm forces it to match the most aggressive competitor. As Alchian (1950) famously argued, market competition functions as an evolutionary selection mechanism; firms that fail to maximize profit risk elimination.
Historical evidence confirms firms consistently choose expansion over work reduction when facing competitive pressure. Manufacturing plants consistently respond to productivity improvements by expanding output rather than reducing labor input (Baily et al., 1992), and even during periods of declining work hours (1890s-1940s), reductions came primarily through labor movement pressure and legislation rather than voluntary firm decisions (Costa, 2000). The overwork pattern Schor (1992) documented reflects this systematic preference for expansion over welfare improvements.
4.1.4 C4: Pareto-Inefficiency.
Universal welfare improvements Pareto-dominate universal output expansion:
If all firms reduce hours proportionally to productivity gains:
  • Workers gain leisure time without income loss
  • Firms maintain relative market positions (no competitive disadvantage)
  • Total output remains sufficient for material abundance
  • Environmental externalities may decrease with reduced production
Compare this to mutual expansion (the equilibrium):
  • Workers work longer hours despite productivity gains
  • Firms engage in costly competition, reducing profit margins
  • Excess output may exceed demand, requiring artificial stimulus
  • Negative externalities from production increase
The equilibrium is recognizably suboptimal. This is precisely what Keynes observed, i.e., the capacity for leisure exists but remains unrealized. Workers could work less, firms could coordinate restraint, yet competition drives continued expansion.
4.1.5 C5: Enforcement Difficulty.
Multiple structural barriers prevent coordination:
International Anarchy: No global authority can enforce coordinated work reduction. National regulations face race-to-bottom dynamics as firms relocate to jurisdictions with longer hours or fewer restrictions. Capital mobility undermines any single nation’s ability to impose work reductions that competitors don’t match.
Verification Challenges: Even with agreements, compliance is difficult to verify. Firms can disguise overtime, create parallel structures, or use contractors. Workers facing job insecurity may voluntarily work longer hours to demonstrate commitment. Actual working hours may diverge substantially from official regulations.
Monitoring Infeasibility: The distributed nature of modern work, such as remote work, knowledge work, and, always-on digital connectivity, makes monitoring difficult. When work becomes cognitive rather than physical, and occurs outside traditional workplace boundaries, enforcement costs escalate.
Commitment Problems: Even if firms agree to coordinate, commitments lack credibility. Each firm has overwhelming incentive to defect once others reduce hours. Without binding enforcement, any agreement is unstable. First-mover disadvantage discourages unilateral action—the firm that reduces hours first suffers competitive disadvantage if others don’t follow.
4.1.6 Result — Structural Tragedy Confirmed:
Productivity competition satisfies all five conditions. Therefore, by Theorem 6, universal output expansion (long working hours) is the unique Nash equilibrium despite being Pareto-dominated by universal work reduction. This explains Keynes’s failed prediction structurally.
4.2 Formalization: The N-Firm Game
We now formalize the heuristic arguments from the previous section to show that the tragedy structure can be derived from first principles under standard assumptions.
Consider $n$ firms indexed by $i\in\{1,\ldots,n\}$ competing in a market. Firms are heterogeneous: firm $i$ has initial output $Q_{i}$ and working hours $H_{i}$ such that productivity is output per hour. We assume that working hours do not exceed the physiological threshold where marginal productivity turns negative.
We define two separate utility functions: one for the firm ($U_{F}$), driven by profit, and one for the workers ($U_{W}$), driven by welfare.
  • Firm Utility ($U_{F}$): This is pure profit: $\text{Revenue}-\text{Labor Cost}$, after ignoring other typical costs such as taxes and costs that may arise in mutual defection state such as additional advertising costs and price wars.
    • $\pi$ is the total market-wide profit.
    • $c$ is the cost per hour of labor.
  • Worker Utility ($U_{W}$): This is welfare: $\text{Wage}+\text{Leisure Value}$.
    • We assume wages are a component of $c$ and are maintained in both scenarios.
    • $\beta$ is the monetized utility value of one hour of leisure. (This is based on Akerlof’s partial gift exchange model (Akerlof, 1982)).
Both utility equations above ignore other components that stay constant between strategy changes.
When productivity increases by factor $\alpha>1$, each firm chooses between two strategies, $s_{i}\in\{C,D\}$:
  • $C$ (Cooperate): Reduce hours proportionally to $H_{i}/\alpha$, maintaining output $Q_{i}$
  • $D$ (Defect): Maintain hours $H_{i}$, expanding output to $\alpha Q_{i}$
Payoff Structure:
The utility for firm $i$ depends on its own strategy $s_{i}\in\{C,D\}$ and the strategy profile of its competitors, $s_{-i}$. In this specific market game, the impact of $s_{-i}$ is fully captured by the total output of all other firms, which we denote as $Q_{-i}(s_{-i})$.
The utility functions for firm $i$ and a worker in the firm are therefore:
  • If Cooperating ($s_{i}=C$): Output is $Q_{i}$, Hours are $H_{i}/\alpha$.
    • $U_{F,i}(C,s_{-i})=\pi\cdot\frac{Q_{i}}{Q_{i}+Q_{-i}(s_{-i})}-c\cdot\frac{H_{i}}{\alpha}$
    • $U_{W,i}(C)=\text{Wage}+\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha}$
  • If Defecting ($s_{i}=D$): Output is $\alpha Q_{i}$, Hours are $H_{i}$.
    • $U_{F,i}(D,s_{-i})=\pi\cdot\frac{\alpha Q_{i}}{\alpha Q_{i}+Q_{-i}(s_{-i})}-c\cdot H_{i}$
    • $U_{W,i}(D)=\text{Wage}$
Proposition 7 (Dominance in Productivity Competition).
For any firm $i$, competitor strategy profile $s_{-i}$ with total competitor output $Q_{-i}(s_{-i})>0$, and productivity factor $\alpha>1$:
  1. (i)
    Defection yields strictly higher market share than cooperation.
  2. (ii)
    Defection is the dominant strategy, i.e., $U_{i}(D,s_{-i})>U_{i}(C,s_{-i})$ for all $s_{-i}$, provided the marginal profit from captured market share exceeds the marginal value of foregone leisure.
Proof.
(i) The proof that market share for Defecting is strictly greater is unchanged and holds by $\alpha>1$. (ii) We find the net profit gain for the firm from defecting:
$U_{F,i}(D,s_{-i})-U_{F,i}(C,s_{-i})$ $=\left(\pi\cdot\frac{\alpha Q_{i}}{\alpha Q_{i}+Q_{-i}(s_{-i})}-c\cdot H_{i}\right)$
$\quad-\left(\pi\cdot\frac{Q_{i}}{Q_{i}+Q_{-i}(s_{-i})}-c\cdot\frac{H_{i}}{\alpha}\right)$
Grouping terms:
$U_{F,i}(D)-U_{F,i}(C)$ $=\underbrace{\pi\left(\frac{\alpha Q_{i}}{\alpha Q_{i}+Q_{-i}(s_{-i})}-\frac{Q_{i}}{Q_{i}+Q_{-i}(s_{-i})}\right)}_{\text{Market Share Gain (Always Positive)}}$
$\quad-\underbrace{\left(c\cdot H_{i}-c\cdot\frac{H_{i}}{\alpha}\right)}_{\text{Labor Cost Increase (Always Positive)}}$
Simplifying the cost term:
$U_{F,i}(D)-U_{F,i}(C)=\text{Market Share Gain}-\text{Labor Cost Increase}$
Defection is the dominant strategy (C3 is met) with the assumption that the profit from the Market Share Gain is greater than the Labor Cost Increase from not reducing hours.
Assumption Justification:
Historical evidence confirms this assumption holds in practice: firms consistently choose output expansion when productivity increases (Baily et al., 1992; Costa, 2000; Schor, 1992), demonstrating that competitive pressure overwhelms leisure preferences across diverse market structures and firm sizes.
Note that for workers, defection leads to a decrease in utility since
$U_{W,i}(D)-U_{W,i}(C)=(\text{Wage})-(\text{Wage}+\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha})=-\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha}$
This indicates a welfare opportunity loss for workers.
Proposition 8 (Pareto-Inefficiency in Productivity Competition).
Universal Cooperation ($s^{C}$) strictly Pareto-dominates Universal Defection ($s^{D}$).
Proof.
We must show that $U_{i}(s^{C})>U_{i}(s^{D})$ for all actors (both firms and workers). The key insight is that when all firms scale output by the same factor, relative market shares remain unchanged.
  • For Firms ($U_{F}$):
    • $U_{F,i}(s^{C})=\pi\cdot\frac{Q_{i}}{\sum_{j}Q_{j}}-c\cdot\frac{H_{i}}{\alpha}$
    • $U_{F,i}(s^{D})=\pi\cdot\frac{\alpha Q_{i}}{\sum_{j}\alpha Q_{j}}-c\cdot H_{i}$
The difference is:
$U_{F,i}(s^{C})-U_{F,i}(s^{D})$ $=\left(-c\cdot\frac{H_{i}}{\alpha}\right)-\left(-c\cdot H_{i}\right)$
$=\underbrace{c\cdot\frac{(\alpha-1)H_{i}}{\alpha}}_{\text{Labor Cost Savings}}$
This value is strictly positive (assuming $c>0$). Firms are better off under $s^{C}$.
  • For Workers ($U_{W}$):
    • $U_{W,i}(s^{C})=\text{Wage}+\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha}$
    • $U_{W,i}(s^{D})=\text{Wage}$
The difference is:
$U_{W,i}(s^{C})-U_{W,i}(s^{D})=\beta\cdot\frac{(\alpha-1)H_{i}}{\alpha}$
This value is strictly positive (assuming $\beta>0$). Workers are better off under $s^{C}$.
Since both firms and workers are strictly better off, $s^{C}$ Pareto-dominates $s^{D}$. The tragedy holds: firms’ individually rational dominant strategy (Proposition 7) leads to a unique Nash equilibrium ($s^{D}$) that is worse for everyone.
Crucially, this result holds for all firms simultaneously regardless of their sizes, initial hours, or market positions. Large and small firms, those with long and short hours, all benefit from coordinated restraint. The Pareto-domination is genuine as there exists no firm that prefers mutual defection to mutual cooperation. ∎
Note on Market Profit:
The proof assumes total market profit $\pi$ remains constant for analytical clarity. In practice, universal output expansion would likely reduce industry profit through price competition, meaning the actual welfare loss from mutual defection exceeds our estimate. Our formalization therefore provides a conservative lower bound on the tragedy’s severity.
Result — Structural Tragedy Proven:
Combined with the N-player structure, externalities, and enforcement difficulty established in Section 4, productivity competition satisfies the complete set of conditions for structural tragedy. The coordination failure emerges from the fundamental competitive structure of markets facing productivity shocks, not from special assumptions about firm homogeneity or market structure.
4.3 Formalization: Extensions, Robustness, and Limitations
Having established the core coordination failure for working hours under uniform productivity gains, we now examine extensions to other welfare dimensions, robustness to relaxed assumptions, and limitations of the analysis.
4.3.1 Extension to Multiple Welfare Dimensions
While we formalize the coordination failure for working hours, the same structural logic applies to other dimensions of worker welfare. When productivity increases, firms face analogous binary choices:
  • Wages: Raise proportionally (C) vs. maintain while expanding output (D)
  • Benefits: Enhance (C) vs. maintain while expanding output (D)
  • Job security: Stabilize employment (C) vs. maintain flexibility while expanding (D)
Each dimension exhibits the same five conditions for structural tragedy. Which dimension manifests the coordination failure depends on institutional constraints. Hour regulations shift pressure to wages; wage floors shift pressure to employment; strong employment protections shift pressure to benefits. The declining labor share of income globally (Karabarbounis and Neiman, 2014; Elsby et al., 2013) provides evidence that productivity gains systematically flow away from worker welfare across all dimensions, though the specific manifestation varies by institutional context.
4.3.2 Relaxing the Uniform Productivity Assumption
The baseline analysis assumes all firms experience the same productivity gain $\alpha$. This simplification is most realistic for economy-wide technological shifts (e.g., electrification, computerization, internet) that eventually diffuse broadly, or for long-run analysis after differential adoption rates have equilibrated through competitive selection.
Dominance Under Heterogeneity:
With heterogeneous productivity gains $\alpha_{i}$ varying across firms, the dominance result (Proposition 7) remains fully valid: each firm $i$ has incentive to expand given its own $\alpha_{i}>1$, regardless of others’ productivity gains. The proof is identical, replacing $\alpha$ with $\alpha_{i}$ throughout. Therefore the core mechanism driving the tragedy—individual incentive to defect regardless of what others do—persists under heterogeneous productivity shocks.
Pareto-inefficiency Under Heterogeneity:
The Pareto-inefficiency result (Proposition 8) assumes comparable productivity gains. With heterogeneous $\alpha_{i}$, workers remain strictly better off under $s^{C}$, but the outcome for firms depends on the distribution of productivity gains.
When heterogeneity is extreme, high-productivity firms may prefer universal defection to capture market share from rivals. However, such extreme heterogeneity leads to rapid market consolidation through the exit of low-productivity firms (Melitz, 2003). The surviving firms then face the tragedy among themselves: once similar-productivity firms remain in the market, Pareto-inefficiency re-emerges among the competitive set. Thus, extreme heterogeneity does not eliminate the tragedy but rather localizes it to the relevant competitive cohort.
For the empirically common case of moderate heterogeneity within industries, Pareto-inefficiency holds: the productivity distribution of surviving firms exhibits substantial but not extreme dispersion (Syverson, 2011), and all firms benefit from coordinated restraint relative to the racing equilibrium. The dominance property (Condition 3) ensures that even when C4 holds weakly, coordination remains structurally difficult because every firm prefers unilateral defection.
4.3.3 Generality and Robustness
The formalization establishes that productivity competition exhibits structural tragedy under general conditions:
Beyond Binary Choices:
The dominance property is not an artifact of the binary choice model. It extends to continuous strategy spaces where firms choose specific hours $h_{i}\in[0,H_{max}]$. In such settings, the marginal incentive to increase output (to capture market share) persists at every level of $h_{i}$ provided the competitive elasticity is sufficiently high.
Risk Aversion Strengthens the Tragedy:
Since the utility loss from losing market share (existential risk) outweighs the utility gain from leisure, risk-averse firms face even stronger pressures to defect than risk-neutral ones.
Generality Across Market Structures:
The tragedy holds for:
  • Firms of any size distribution (concentrated or competitive markets)
  • Any initial working hours across firms
  • Any productivity gain $\alpha>1$
  • Any configuration of competitor choices
  • Heterogeneous productivity gains $\alpha_{i}$ (within the surviving competitive cohort)
4.3.4 Welfare Cost Increases with Productivity
Sensitivity analysis of the utility functions reveals a counter-intuitive dynamic: the welfare cost of the coordination failure increases with technological progress. We define the “cooperative surplus” as the difference between welfare in the cooperative outcome ($s^{C}$) and the Nash equilibrium ($s^{D}$). As the productivity factor $\alpha$ rises, this surplus grows asymptotically toward the total labor cost, meaning the gap between potential leisure and actual leisure widens.
This implies that highly productive economies sacrifice significantly more potential welfare than low-productivity ones. Technological abundance amplifies rather than resolves the structural waste. Modern economies with eightfold productivity gains since Keynes could provide far more leisure than 1930s economies, yet the coordination failure prevents realizing this potential.
4.4 Empirical Validation: The European Case
The European experience provides crucial empirical validation. If our structural analysis is correct, even under maximally favorable institutional conditions, we should observe: (1) partial mitigation achieving some welfare improvements but not full Keynesian leisure, (2) high costs requiring substantial institutional apparatus, and (3) erosion over time as competitive pressure gradually undermines coordination.
Favorable Conditions:
Europe represents the most favorable case for coordination: wealthy democracies with strong institutions, supranational governance (European Union) providing coordination capacity, cultural consensus supporting work-life balance, strong labor unions with political influence, and legal frameworks mandating maximum hours and minimum vacation. If coordination is possible anywhere, it should succeed in Europe.
Partial Mitigation:
European workers have achieved shorter hours than American workers—approximately 1,350 hours annually in Germany versus 1,800 in the US (OECD, 2025). This represents real achievement: a reduction of approximately 450 hours compared to the United States. However, even German workers work far more than Keynes’s predicted 780 hours annually. The coordination is partial, not complete.
Moreover, productivity-welfare decoupling manifests differently across institutional contexts. While European workers gained leisure relative to US counterparts, Europe also exhibits declining labor share of income since 1980, slower wage growth relative to productivity, persistent pressure on social benefits, and higher unemployment in several major economies. The tragedy is not escaped but redistributed: European workers gained leisure at the cost of wage growth and employment security, while US workers experienced stagnant wages despite longer hours. Both regions show productivity-welfare decoupling driven by competitive pressure, manifesting through whichever dimension institutional constraints leave exposed.
High Costs:
This coordination requires substantial costs.
Economic costs: European GDP growth has lagged the United States over recent decades, with significant divergence since 2008 (World Bank, 2024). While this divergence is driven by multiple factors, shorter working hours impose a mechanical ceiling on total output.
Regulatory costs: Maintaining shorter hours requires extensive regulation: France’s 35-hour workweek involves complex rules about overtime, exceptions, and compliance monitoring (L&E Global, 2024), while Germany’s system requires workplace councils, sector-level bargaining, and intricate labor law.
Competitive pressure: European firms face competitive disadvantages against firms from regions with longer working hours, generating continuous political pressure to loosen regulations.
Erosion Over Time:
Most tellingly, coordination shows signs of erosion. France’s 35-hour law contains numerous exceptions and workarounds; actual hours often exceed nominal limits (L&E Global, 2024). Germany’s Hartz reforms increased flexibility and effectively lengthened working hours in some sectors (IAB-Forum, 2023). European working hours have largely stabilized since the 1980s despite continued productivity growth (Huberman and Minns, 2007); the decline observed from 1870-1980 has stalled, suggesting coordination has reached its limits under persistent competitive pressure.
Interpretation:
The European case demonstrates that institutional intervention can achieve partial, costly mitigation—not elimination of the tragedy. Institutions shift the equilibrium toward Pareto-optimality but cannot reach it absent structural change to competitive conditions themselves. Even under maximally favorable conditions, coordination remains partial (hours far exceed technological capacity), costly (requiring extensive institutional apparatus), and fragile (showing continuous erosion). The underlying tragic structure persists even when temporarily overridden.
4.5 Why Standard Solutions Face Structural Barriers?
Our framework explains why commonly proposed solutions to the productivity-welfare decoupling face structural barriers. Each proposed solution fails to address one or more of the five necessary conditions.
Solution 1 — Individual Firm Action:
Encourage firms to voluntarily reduce working hours or increase wages proportional to productivity gains.
Problem: Violates dominance condition (C3). Any firm reducing hours unilaterally suffers competitive disadvantage while competitors capture its market share. The logic of collective action (Olson, 1965) explains why voluntary restraint by individual firms cannot succeed when defection yields higher payoffs regardless of competitors’ strategies. First-mover disadvantage is severe in competitive markets.
Solution 2 — Industry-Level Voluntary Agreements:
Create voluntary agreements among firms to coordinate on reduced hours or higher wages.
Problem: Violates enforcement condition (C5). As Stigler (1964) demonstrated for cartel stability, voluntary agreements among competitors are inherently unstable because each firm has a dominant incentive to defect by secretly expanding hours to capture market share while others comply. While reputational mechanisms or industry associations might provide weak enforcement, these cannot overcome the profit advantages from defection. Without government enforcement imposing credible sanctions, such agreements inevitably collapse.
Solution 3 — National Regulation:
Implement national labor standards requiring reduced working hours or minimum wage increases tied to productivity.
Problem: Faces international competition and capital mobility. Böckerman and Maliranta (2012) provide micro-level evidence from Finland: despite strong labor market institutions and centralized wage bargaining, exposure to international trade systematically drove exits of high-labor-share plants, reducing the aggregate labor share by reallocating production toward more competitive establishments. This validates the race-to-the-bottom dynamic (Rodrik, 1997), demonstrating that even wealthy democracies with strong institutions face structural pressure from international competition. National regulation without global coordination merely shifts production to less regulated jurisdictions.
Solution 4 — International Labor Agreements:
Establish international treaties or agreements on labor standards, including maximum working hours and minimum wages tied to productivity.
Problem: Violates enforcement condition (C5). Flanagan (2003) demonstrates that international labor standards face severe enforcement barriers: the International Labour Organization lacks sanctioning power, compliance is voluntary, and nations face strong incentives to free-ride by endorsing standards publicly while maintaining lax enforcement to attract capital investment. As Ostrom (1990) established, effective governance requires graduated sanctions and low-cost monitoring, both absent at the international level due to sovereignty concerns and international anarchy.
Solution 5 — Technological Unemployment:
Assume automation will eventually force shorter working weeks as human labor becomes unnecessary.
Problem: Misidentifies the mechanism. As Autor (2015) documents, despite dramatic technological progress, employment has remained robust because technology creates complementary tasks even as it automates existing ones. More fundamentally, automation increases productivity, which intensifies rather than resolves the coordination failure. The dominance condition (C3) ensures that firms use productivity gains to expand market share rather than reduce hours, unless coordination mechanisms force otherwise. Historical evidence supports this: productivity increased eightfold since Keynes’s prediction, yet working hours declined far less than proportionally.
Solution 6 — Cultural Change:
Encourage societal preference shifts toward valuing leisure over consumption.
Problem: Insufficient against structural pressure. Even if workers prefer leisure, competitive dynamics compel long hours. Schor (1992) documents that despite expressed preferences for more leisure time, American workers face structural pressures that override these preferences. The dominance condition (C3) ensures that cultural preferences alone cannot overcome the structural necessity of maintaining competitive position when others work longer hours. Preferences must align with incentives, which the competitive structure prevents.
These solutions face structural barriers because they fail to address the complete set of conditions. Partial solutions that leave some conditions unsatisfied achieve at most temporary, costly, and eroding coordination, exactly as observed in Europe.
4.6 What Would Work?
Our framework identifies interventions that could resolve the productivity tragedy by breaking one or more of the five necessary conditions. We emphasize that these interventions are structural requirements, not policy recommendations, to clarify the magnitude of barriers any solution must overcome.
Break Condition 1 (N-Player Structure):
Consolidate decision-makers into smaller groups, making coordination feasible through reduced free-rider problems. As Olson (1965) established, smaller groups ($n$) dramatically lower coordination costs.
However, achieving this through industrial consolidation faces severe barriers. Antitrust law in most jurisdictions prohibits anticompetitive coordination, and for good reason: consolidation that enables coordination on labor practices would simultaneously enable price-fixing and output restriction, trading one market failure (coordination tragedy) for another (monopoly power) that harms consumer welfare. Alternatively, international consolidation of labor regulation authority could reduce the number of competing jurisdictions, but this faces political barriers from sovereignty concerns and competing national interests.
Break Condition 2 (Binary Choice with Externalities):
Eliminate the competitive externality where one firm’s expansion harms competitors’ market positions.
However, this is structurally impossible since competitive pressure is inherent to market economies. As Schumpeter (1942) argued, competition through creative destruction—where successful firms displace less efficient ones—is capitalism’s fundamental mechanism for driving productivity growth. Removing this externality would eliminate the competitive process itself. The externality is reciprocal: protecting compliant firms from market share loss necessitates restricting expanding firms, which contradicts the productivity-enhancing function of market competition.
Break Condition 3 (Dominance Property):
Reduce the existential stakes of restraint so that cooperation doesn’t threaten firm or worker survival.
For workers, mechanisms include Universal Basic Income (UBI), stronger unemployment insurance, or other social safety nets that provide unconditional material security. As Van Parijs (1995) argues, guaranteeing an unconditional material baseline grants workers “real freedom” to refuse exploitative employment terms, breaking the dominance of survival pressure that compels acceptance of maximum-intensity work arrangements.
For firms, mechanisms would include subsidies, tax incentives, or market structures that reward restraint rather than penalizing it; for example, government contracts preferentially awarded to firms with reduced hours, or subsidies compensating for reduced output during hour reductions.
However, both face substantial implementation barriers. For worker-side mechanisms, fiscally adequate UBI levels require substantial tax increases or redistribution that face political opposition across diverse political systems (Martinelli, 2017). Even where safety nets exist, they may be insufficient to overcome competitive pressure when rivals maintain longer hours; workers with basic income still face relative income competition and status concerns that perpetuate long hours (Frank, 1999).
For firm-side mechanisms, subsidies reducing bankruptcy risk from reduced output face severe moral hazard concerns: firms might claim hour reductions while maintaining actual hours, or exploit subsidies without genuine coordination. Additionally, such programs would require ongoing substantial public expense and create competitive distortions favoring subsidized over non-subsidized firms. Most fundamentally, any mechanism rewarding restraint must overcome the problem that market structures inherently reward expansion—changing this requires fundamentally restructuring competitive dynamics, which returns to the impossibility of breaking Condition 2 (eliminating competitive externalities).
Break Condition 4 (Pareto-Inefficiency):
Accept the equilibrium as efficient rather than seeking coordination. Some economists, notably Prescott (2004), argue that transatlantic differences in working hours reflect rational responses to marginal tax rate differences rather than coordination failures, with Americans optimally choosing longer hours given higher after-tax wages.
However, as Pencavel (2018) demonstrates, observed hours often do not reflect worker preferences; instead, institutional and market constraints significantly determine working time, and workers cannot simply choose their preferred hours. Additionally, the tax-incentive explanation cannot account for the secular decline in labor share within countries over time, including during periods of tax stability. The micro-level restructuring documented by Böckerman and Maliranta (2012)—where high-labor-share plants systematically exit under competitive pressure—provides direct evidence of coordination failure operating alongside any tax or preference effects. If workers genuinely prefer current hours and the equilibrium is Pareto-efficient, no coordination problem exists. Our framework and the empirical evidence suggest otherwise.
Break Condition 5 (Enforcement Difficulty):
Solve the enforcement problem through effective global labor standards with comprehensive monitoring and binding sanctions. This requires international agreement on maximum hours and minimum wage-productivity linkages, verification mechanisms for compliance across jurisdictions, sanctions sufficient to deter defection, and sufficient participation to prevent race-to-bottom through non-compliant jurisdictions.
However, effective global enforcement faces formidable barriers. As Ostrom (1990) identified, successful commons governance requires graduated sanctions and low-cost monitoring—both absent in global labor markets due to international anarchy, sovereignty concerns, and capital mobility that enables regulatory arbitrage. The International Labour Organization establishes standards but lacks enforcement power; achieving binding global labor standards would require unprecedented coordination among competing nations, effectively solving a higher-order coordination problem to address the first-order productivity tragedy.
Comprehensive Solution:
The productivity tragedy can be solved, but only through interventions that address its structural foundations. Given the difficulty of breaking any single condition completely, the most feasible approach likely combines partial interventions across multiple conditions: strengthening social safety nets (C3), enhancing coordinated regional labor standards (C5) among economically integrated blocs, and accepting modest efficiency costs from reduced competitive pressure. However, as European experience demonstrates, even such comprehensive efforts face continuous erosion from global competitive pressure. Incremental reforms that leave the five conditions substantially intact will achieve at most partial, costly, and eroding coordination.
5 The Tragedy of AI Governance
We now apply the unified framework to AI development, demonstrating that AI governance faces the same five conditions but with amplified intensity. This makes coordination structurally more difficult than for historical arms control cases including nuclear, biological, and chemical (NBC) weapons.
Scope of Analysis:
Our analysis focuses on the governance challenge posed by transformative AI breakthroughs: systems approaching or exceeding human-level general intelligence (AGI) or superintelligence. We assume such systems pose non-excludable existential risks if developed without adequate safety measures (Bostrom 2014, Russell 2019, Bengio 2025). We examine the implications of relaxing this assumption after establishing Pareto-inefficiency (C4).
5.1 Mapping the Five Conditions to AI Development
5.1.1 Condition 1: N-Player Structure
AI development involves numerous independent actors: AI labs, tech companies, universities and research institutions, nations, and open-source communities.
The number of actors is large and growing. No single actor can coordinate AI development unilaterally. The architecture is fundamentally decentralized across private sector, public sector, and international actors.
5.1.2 Condition 2: Binary Choice with Externalities
When AI capabilities advance, each actor faces essentially a binary choice:
  • Strategy $C$ (Cooperate/Restrain): Slow development to ensure safety, conduct thorough testing, wait for governance mechanisms
  • Strategy $D$ (Defect/Advance): Accelerate capabilities development to maintain competitive position
Individual choices create negative externalities. When one actor accelerates development:
  • Competitors face pressure to accelerate or lose market/strategic position
  • Safety margins narrow as capabilities race ahead of governance
  • Risks from powerful AI systems affect everyone regardless of who develops them
  • First-mover advantages create winner-take-all dynamics
The externality is severe since one actor’s defection substantially increases risks for all.
5.1.3 Condition 3: Dominance Property
Acceleration dominates restraint regardless of competitors’ choices:
Case 1—If competitors restrain: Actor gains decisive first-mover advantage by accelerating. In AI markets characterized by network effects, data advantages, and talent concentration, first-mover advantages are enormous. The actor that reaches advanced capabilities first may capture the entire market.
Case 2—If competitors accelerate: Actor must accelerate to avoid strategic elimination. Falling behind in AI capabilities may mean permanent market displacement for firms or strategic vulnerability for nations. When competitors accelerate, restraint means obsolescence.
The dominance is driven by existential stakes for firms, nations, and researchers. For example, a firm who cannot raise capital for massive compute (or acceleration) risk obsolescence. This structural pressure forced even the leading non-profit actor to restructure first into a ‘capped-profit’ entity (OpenAI, 2019) and eventually into a Public Benefit Corporation to remove fundraising limits (OpenAI, 2025).
5.1.4 Condition 4: Pareto-Inefficiency
Universal restraint Pareto-dominates universal acceleration:
If all actors coordinate to slow development and prioritize safety:
  • Risks from advanced AI are reduced through careful development
  • Relative competitive positions are preserved (no disadvantage from restraint)
  • Time is available for governance mechanisms to develop
  • Benefits from AI still accrue but with better risk management
Compare this to mutual acceleration (the equilibrium):
  • Risks accumulate faster than governance mechanisms develop
  • Safety testing is compressed or skipped under time pressure
  • First-mover races waste resources on competitive positioning
  • Potential for catastrophic outcomes that benefit no one
Leading AI researchers explicitly recognize this Pareto-domination. Surveys show substantial minorities estimating non-trivial probabilities of catastrophic outcomes (Bengio et al., 2025; Grace et al., 2018). The AI safety community argues current trajectories are collectively harmful (Askell et al., 2019). Yet development accelerates.
Asymmetric Gains and the Winner-Take-All Problem:
AI development creates potential for extremely asymmetric outcomes. Unlike productivity competition where firms maintain roughly proportional positions, AI capabilities races may produce winner-take-all dynamics where the first to achieve advanced AI captures decisive advantages, potentially controlling entire markets, achieving strategic dominance, or even technological singularity.
This asymmetry threatens the Pareto-inefficiency condition. If an actor believes they will be the winner, they might rationally prefer the racing outcome (where they capture enormous gains) over a cooperative outcome (where advantages are constrained or shared). For prospective winners, the situation appears not as a tragedy but as a rational pursuit of dominance.
However, three factors restore the tragedy structure even for prospective winners:
First, existential risk is non-excludable.
The catastrophic risks from misaligned or insufficiently robust advanced AI systems affect all actors regardless of who develops them first (Bostrom, 2014). A winner who achieves advanced AI through hasty development faces the same risks of loss of control, unintended consequences, or catastrophic failure as any other actor. The instrumental convergence thesis (Bostrom, 2014) suggests that advanced AI systems will pursue subgoals (self-preservation, resource acquisition, goal preservation) that may conflict with human interests regardless of which actor deploys them. Winners and losers alike face potential extinction or permanent disempowerment.
Crucially, the tragedy persists because the rate of capability gains in a race condition naturally outpaces the rate of safety verification. Consequently, the winner is statistically likely to deploy an unsafe system before establishing control, rendering the expected utility of ‘winning’ negative even for the victor. This conclusion is reinforced by empirical data: Bengio et al. (2025); Grace et al. (2018) found that substantial minorities of AI researchers estimate non-negligible probabilities of extremely bad outcomes including human extinction. If these probability estimates are even roughly accurate, the astronomical loss associated with an unsafe deployment outweighs even enormous competitive gains, making the race irrational for all participants.
Second, winner identity is uncertain ex ante.
While racing, no actor knows with certainty they will be the winner. Multiple actors have credible capabilities. Geopolitical and corporate competition involves numerous sophisticated actors, any of whom might achieve breakthroughs. The “DeepSeek Shock” of early 2025 serves as a potent empirical validator of this uncertainty. By achieving frontier-level reasoning capabilities with orders-of-magnitude greater training efficiency, a resource-constrained actor demonstrated that massive capital and hardware ‘moats’ are not impregnable (DeepSeek-AI, 2025). Crucially, while this lowers barriers to entry, it does not reduce the ‘winner-take-all’ prize of the final breakthrough; instead, it increases the volatility of the race by proving that current leaders cannot rely on resource dominance to secure their future position. The identity of the ultimate winner remains radically uncertain. From behind this veil of uncertainty, all actors—even current leaders—face expected losses from an uncoordinated race.
This parallels arms races where no state knows ex ante whether it will achieve dominance. Even if one state would retrospectively prefer the outcome where it achieved hegemony, all states prefer coordinated restraint to a race with uncertain winners and certain risks.
Third, even decisive technological leads may be temporary.
Unlike traditional military conquest where territory captured remains under control, AI capabilities can potentially be copied, stolen, or independently rediscovered. Winner-take-all assumptions may prove mistaken if multiple actors achieve advanced AI within relatively short timeframes, or if AI capabilities prove impossible to monopolize due to espionage, independent discovery, or proliferation.
Therefore, even with extremely asymmetric potential gains, the Pareto-inefficiency condition holds when existential risks are incorporated into rational calculation. The racing equilibrium threatens all actors including prospective winners, making coordinated restraint collectively preferable despite competitive temptations.
Critical caveat:
This defense of Pareto-inefficiency relies on two empirical claims: (1) catastrophic risks from advanced AI are sufficiently probable and severe to dominate even winner-take-all competitive gains, and (2) actors incorporate these risks into their decision-making rather than pure output or power maximization. The first claim is defended by substantial technical literature on AI safety (Bostrom, 2014; Russell, 2019; Amodei et al., 2016). The second claim is more tenuous: if actors systematically discount existential risks or believe they can control advanced AI despite expert warnings, the existential dimension of the tragedy weakens.
However, even under systematic discounting of existential risk, the tragedy structure persists at reduced intensity. Coordination failure would manifest through wasteful resource competition: duplicated R&D spending across labs, unsustainable talent bidding wars, massive capital obligations against uncertain returns, and infrastructure investments vulnerable to algorithmic breakthroughs that devalue hardware moats-as DeepSeek’s efficiency gains demonstrated. The resulting “war of attrition” remains Pareto-inefficient (all participants would prefer coordinated investment), lowering C3 and C4 from extreme to high intensity. The enforcement difficulty (C5) remains at extreme intensity regardless of risk estimates, as the technical barriers to verification are unchanged. The tragedy persists; only its stakes diminish.
5.1.5 Condition 5: Enforcement Difficulty
Multiple structural barriers prevent coordination:
International Anarchy: No global authority can enforce coordinated AI governance. Nations compete for AI advantages. International agreements lack binding force. AI is dual-use, making arms control agreements difficult to negotiate.
Verification Impossibility: AI development occurs in software, making verification extremely difficult. Model weights can be copied and modified. Training runs can occur covertly.
Monitoring Infeasibility: Even with agreements, monitoring compliance faces severe challenges. What constitutes ‘dangerous’ AI capabilities may be ambiguous until after development. The distributed nature of AI research across many institutions resists centralized monitoring.
Commitment Problems: Even if actors agree to coordinate, commitments lack credibility. Each actor has overwhelming incentive to defect and gain first-mover advantages once others restrain. Secret defection is technologically feasible. First-mover disadvantage from restraint is severe.
5.1.6 Result: Structural Tragedy Confirmed
AI development satisfies the complete set of conditions with high intensity when catastrophic risks are incorporated into rational assessment. The Pareto-inefficiency condition is satisfied for all actors including prospective winners because existential risks are non-excludable, winner identity is uncertain ex ante, and even decisive leads may prove temporary. Therefore, by Theorem 6, universal acceleration is the unique Nash equilibrium despite being Pareto-dominated by coordinated restraint.
As noted in the critical caveat above, actors who systematically discount existential risks would still face a structural tragedy at reduced intensity (Index $\approx 1,000$ rather than $\approx 3,981$), comparable to climate change and bank runs. The empirical evidence suggests leading actors do recognize catastrophic risks even while competitive pressure compels continued development, precisely the tragedy structure our framework predicts.
Robustness of the Racing Dynamic:
Just as with productivity, this tragedy structure is robust to relaxed assumptions. The dominance of acceleration holds in continuous strategy spaces: if actors choose a level of investment or compute $k_{i}\in[0,K_{max}]$, the marginal strategic advantage of increasing $k_{i}$ remains positive as long as competitors are advancing. Consequently, ‘slowing down’ is not a stable equilibrium; the Nash equilibrium collapses to maximum feasible acceleration ($K_{max}$). Furthermore, strategic risk aversion (fear of falling behind) amplifies the race. Because the loss of sovereignty or market relevance is viewed as a certain, immediate catastrophe, while coordination failure consequences (whether existential or economic) are viewed as probabilistic and future catastrophe, risk-averse actors rationally prioritize avoiding the immediate deterministic loss, thereby intensifying the acceleration.
5.2 AI Governance vs. Historical Arms Control: An Eight-Dimension Comparison
To assess whether AI governance is more or less difficult than previous coordination challenges, we systematically compare AI to successful arms control cases. Nuclear, biological, and chemical (NBC) weapons, all achieved substantial (though imperfect) international coordination. How does AI compare?
Table 3 summarizes the comparison across eight crucial dimensions. We now discuss each dimension in detail.
Table 3: AI Governance vs. Historical Arms Control Across Eight Dimensions. AI represents a structural outlier, facing the most unfavorable condition for coordination across all eight dimensions simultaneously, whereas historical cases benefitted from mitigating factors (e.g., observability, separability) in key areas.
Dimension Chemical / Bio. Nuclear AI
Strategic Value Military only, limited Military, high Military + Economic, extreme
Dual-Use Separable (some) Separable (enrichment) Inseparable
Verification Observable (facilities) Observable (materials) Unobservable (software)
Reputational Costs High (taboo) High (stigma) Low (prestige)
Entry Barriers Moderate Very High Low
Substitutes Yes (conventional) Yes (conventional) None
Time Scale Decades warning Decades warning Years warning
Convergence Military only Military only Military + Economic
5.2.1 Dimension 1: Strategic Value
Chemical and biological weapons have limited strategic value. They are unreliable, difficult to control, and often militarily ineffective compared to conventional weapons. This limited value reduced incentives for development and made restraint easier.
Nuclear weapons have high strategic value for deterrence, making complete elimination difficult. However, their value is purely military. Economic incentives for nuclear development are weak.
AI has extreme strategic value in both military and economic domains. Military applications include autonomous weapons, intelligence analysis, cyber operations, and logistics optimization. Economic applications include productivity enhancement, automation, market advantage, and platform dominance. This dual-domain value creates reinforcing incentives far stronger than for any previous arms control case.
5.2.2 Dimension 2: Dual-Use Inseparability
Chemical and biological agents have some separability between civilian and military applications. Industrial chemicals differ from weaponized agents. Civilian biology can be distinguished from weaponization programs with moderate effort.
Nuclear programs have some separability. Civilian nuclear power uses low-enriched uranium; weapons require highly enriched uranium or plutonium. Monitoring enrichment levels provides some verification capability.
AI research and development is completely inseparable between civilian and military applications. The same models, training techniques, and hardware serve both purposes. Foundational research benefits both domains simultaneously. Attempting to restrict military AI while permitting civilian AI is structurally impossible—they are identical technologies with different deployment contexts.
This inseparability creates severe enforcement problems. Any agreement to restrict military AI would face immediate problems: civilian AI companies would argue their research is purely civilian, verification would be impossible, and the distinction itself would be meaningless.
5.2.3 Dimension 3: Verification
Chemical and biological weapons require observable facilities leaving physical traces. Nuclear weapons require large enrichment facilities and detectable materials, enabling moderate verification success. AI development occurs in software and is fundamentally unverifiable—model weights can be copied instantly, training runs can occur covertly, and testing for dangerous capabilities requires developing the very systems one seeks to restrict.
5.2.4 Dimension 4: Reputational Costs
Chemical and biological weapons face powerful taboos with strong international norms (Price, 1995; Organisation for the Prohibition of Chemical Weapons (2025), OPCW; United Nations Office for Disarmament Affairs, 2025). Nuclear weapons face mixed stigma and prestige. AI development faces inverse reputational dynamics, leading development confers prestige rather than stigma, making restraint appear as weakness rather than responsibility.
5.2.5 Dimension 5: Entry Barriers
Nuclear weapons face very high barriers (billions of dollars, advanced technology, years of development) that concentrated capabilities in few hands, enabling coordination. AI faces low and declining entry barriers: frontier development costs hundreds of millions but is accessible to many actors and declining with algorithmic progress, multiplying the coordination problem.
5.2.6 Dimension 6: Substitutes
Chemical, biological, and nuclear weapons have conventional substitutes for most military objectives. AI has no substitutes for either economic or military applications—actors who restrain fall irreversibly behind, amplifying the dominance condition.
5.2.7 Dimension 7: Time Scale
Chemical and nuclear weapons had decades between emergence and arms control (1925-1993 for chemical, 1942-1968 for nuclear). AI governance faces compressed timelines: modern deep learning emerged around 2012, with expert surveys estimating a 50% chance of human-level AI within roughly two decades (Grace et al., 2018), leaving far less time for institutional development.
5.2.8 Dimension 8: Economic-Military Convergence
Chemical and biological weapons involved only military considerations. Nations developed and deployed them (when they did) for military advantage, making the coordination problem purely security-focused.
Nuclear weapons similarly involved only military considerations, with the exception of some civilian nuclear power development which proved separable.
AI represents complete convergence of economic and military imperatives. Every advance in civilian AI capabilities immediately enhances military capabilities. Every AI laboratory contributes simultaneously to economic competitiveness and military potential. Nations cannot choose to prioritize economic development over military security or vice versa—AI advancement serves both goals identically.
This convergence creates reinforcing pressures absent in previous coordination challenges. Even if nations could overcome military competition (as partial nuclear arms control achieved), economic competition would independently drive AI development. Even if firms could coordinate to slow capabilities races for safety, military imperatives would independently accelerate development. The dual reinforcement makes coordination structurally more difficult than previous single-domain challenges.
5.2.9 Summary Across Dimensions
AI faces unfavorable conditions on all eight dimensions compared to successful historical arms control:
  • Higher strategic value (dual-domain)
  • Complete dual-use inseparability
  • Fundamental verification impossibility
  • Reversed reputational incentives
  • Lower entry barriers multiplying actors
  • No substitutes increasing opportunity costs
  • Compressed timelines preventing institutional development
  • Economic-military convergence creating reinforcing pressures
This systematic comparison establishes that AI governance faces coordination challenges structurally worse than any previous arms control case. AI faces unfavorable conditions on every dimension simultaneously. The inseparability of economic and military AI (the dual-use dilemma) is the decisive structural barrier: stopping military AI requires stopping economic AI, which is suicidal for a nation’s productivity.
5.3 Empirical Validation: The Russia-Ukraine Drone War
The Russia-Ukraine conflict validates the framework’s structural predictions for competitive dynamics under existential stakes. When Russia invaded Ukraine in February 2022, both sides had minimal drone capabilities. Within two years, deployment escalated from dozens to thousands of drones monthly (Center for Strategic and International Studies, 2024; Watling and Reynolds, 2023), with Ukraine expected to produce 4.5 million drones in 2025 (Nieczypor and Matuszak, 2025). This exponential growth occurred despite years of prior international dialogue on weapons governance.
The escalation demonstrates our framework’s core dynamic. Each side faced overwhelming incentive to match the other’s deployment (C3: dominance driven by existential stakes). Drone deployment by one side created immediate pressure for counter-deployment (C2: negative externalities). No international authority could enforce restraint during active conflict (C5: enforcement impossibility). Pre-war governance discussions proved irrelevant once competitive pressure intensified.
Current drones remain predominantly remotely piloted: physically observable, with detectable manufacturing and verifiable deployment. Yet governance still failed. Now consider autonomous drones: more lethal, requiring fewer operators, capable of faster response, but also more prone to errors without human oversight. The governance challenge steepens: verifying autonomous compliance requires inspecting software logic, yet strategic incentives preclude sharing targeting code that would reveal vulnerabilities to adversaries. The transition from remote-controlled to autonomous systems amplifies every dimension of the coordination failure while adding verification impossibility.
If governance fails for remotely piloted drones, a relatively governable technology, it will fail more decisively for autonomous drones. And if autonomous drone governance proves intractable, consider AI integrated across the full spectrum of military systems: autonomous vehicles, cyber operations, logistics, intelligence analysis, command and control. Each integration multiplies the coordination challenge our framework identifies. The drone war thus serves as both confirmation of the framework and a conservative lower bound for AI governance challenges ahead.
5.4 Why Standard Solutions Face Structural Barriers?
Our framework explains why commonly proposed AI governance solutions face structural barriers. Each solution fails to adequately address one or more of the five necessary conditions.
Solution 1 — Voluntary Industry Self-Regulation:
Encourage AI labs to voluntarily adopt safety standards and slow development.
Problem: Violates dominance condition (C3). As Maas (2019) argues in the context of military AI arms control, strategic stability concerns make voluntary restraint inherently unstable—any actor that restrains development while competitors advance faces strategic elimination. The 2023 open letter calling for an AI development pause produced no actual pause despite thousands of signatures (Future of Life Institute, 2023), demonstrating that even widespread recognition of risks cannot overcome competitive pressure. Even safety-conscious organizations face overwhelming pressure to accelerate when rivals make breakthroughs. OpenAI’s trajectory exemplifies this: founded as a non-profit to ensure safe AI development, competitive necessity forced restructuring first into a “capped-profit” entity (OpenAI, 2019) and eventually into a Public Benefit Corporation to remove fundraising limits (OpenAI, 2025). Voluntary restraint is unstable under winner-take-all competitive dynamics.
Solution 2 — International Treaties:
Establish binding international agreements limiting AI capabilities development, analogous to nuclear non-proliferation treaties.
Problem: Violates enforcement condition (C5). As our eight-dimension comparison (Table 3) demonstrates, AI faces more severe verification challenges than biological or chemical weapons. Koblentz (2009) documents how verification failures have rendered the Biological Weapons Convention largely ineffective despite physical detectability of biological weapons programs. AI faces even steeper challenges: software can be developed covertly, training can be concealed, and capabilities cannot be reliably detected without access to proprietary systems. International anarchy means no global authority can enforce compliance, and verification impossibility precludes credible monitoring.
Solution 3 — AI Safety Research:
Solve the technical alignment problem, making advanced AI systems safe by design.
Problem: Insufficient to resolve coordination failure even if technically successful. Safety research addresses the risk of misaligned AI but cannot eliminate competitive pressure from aligned AI. As discussed in Condition 4 in Section 5.1, even perfectly aligned AI systems create winner-take-all dynamics through economic and strategic advantages. The coordination failure arises from competitive pressure to be first, which safety research alone cannot eliminate. Moreover, safety research itself faces the racing dynamic: labs may cut corners on safety validation when competitors approach breakthroughs (Armstrong et al., 2016).
Solution 4 — National Regulation:
Implement comprehensive safety requirements and capability restrictions within individual nations.
Problem: Violates enforcement condition (C5) through regulatory arbitrage. As Armstrong et al. (2016) formally model, unilateral safety regulations create competitive disadvantage, causing development to migrate to less cautious jurisdictions—merely shifting rather than reducing existential risk. The decentralized, software-based nature of AI development enables rapid relocation compared to physical industries. Without global coordination, national regulation faces the same race-to-bottom dynamics as labor standards under globalization.
Solution 5 — Compute Governance:
Restrict access to advanced computing hardware required for frontier AI development, as proposed by Anderljung et al. (2023).
Problem: Faces erosion over time despite initial promise. While compute restrictions could theoretically create entry barriers through hardware chokepoints, Sastry et al. (2024) document how algorithmic efficiency improvements systematically lower hardware requirements, enabling actors to achieve frontier capabilities with previously sub-frontier hardware. The DeepSeek breakthrough of early 2025 exemplifies this dynamic: achieving frontier reasoning capabilities with orders-of-magnitude greater training efficiency effectively undermines static compute-based thresholds. Over time, what requires cutting-edge datacenters today becomes achievable on commodity hardware, eroding the enforcement mechanism. Additionally, compute governance requires international cooperation that faces standard enforcement barriers from sovereignty concerns and verification challenges.
Solution 6 — Energy Monitoring:
Monitor data center energy consumption as a verification mechanism for frontier AI development, since large-scale training requires massive electricity usage. While Sastry et al. (2024) note that energy consumption correlates with compute usage for large training runs, they acknowledge this creates at best an indirect signal of AI development.
Problem: Energy monitoring faces fundamental dual-use problems. Data center energy consumption is observable and growing rapidly, but nations must invest in energy infrastructure for reasons entirely independent of AI—grid modernization, electrification, industrial policy, and general cloud computing demand. Restricting energy investment to limit AI development is economically and politically infeasible given these legitimate alternative uses. Any agreement to monitor energy for AI governance purposes would be trivially circumvented by dual-use justifications: large energy consumption could be attributed to cryptocurrency mining, scientific computing, commercial cloud services, or other legitimate activities. Unlike specialized AI chips that have limited alternative uses, energy infrastructure is inherently general-purpose, making verification through energy monitoring structurally impossible.
These solutions face structural barriers because they fail to address the complete set of conditions simultaneously. The analysis suggests that AI governance requires interventions of unprecedented scope and coordination—qualitatively beyond what successful arms control achieved—precisely because AI development satisfies all five tragedy conditions with extreme intensity.
5.5 What Would Work?
Our framework identifies interventions that could theoretically resolve the AI governance tragedy by breaking one or more of the five necessary conditions. For AI governance, this is structurally more difficult than for productivity due to amplified intensity across all conditions. We also emphasize that these interventions represent structural requirements, not policy recommendations, to clarify the magnitude of barriers any solution must overcome.
Break Condition 1 (N-Player Structure):
Consolidate all transformative AI development into a single global project—a heavily regulated monopoly or international consortium. Bostrom (2014) terms this a “Singleton”: a global coordination mechanism with sufficient authority to prevent competitive AI development. Such an entity could eliminate the racing dynamic by becoming the sole developer.
However, this solution creates severe risks that may exceed those it solves. Concentrating transformative AI development in a single entity enables permanent power consolidation, potential totalitarian control over transformative technology, and elimination of competitive safeguards against misuse. The cure might be worse than the disease, trading coordination tragedy for concentration risk. Additionally, achieving global agreement to establish such an entity faces formidable obstacles given competing national interests, sovereignty concerns, and deep mistrust among major powers regarding who would control such an entity.
Break Condition 2 (Binary Choice with Externalities):
Develop AI systems that inherently do not create strategic advantages, eliminating competitive pressure.
However, this is structurally difficult because AI capabilities inherently create strategic advantages since economic productivity, military capability, or scientific advancement cannot be developed in ways that do not affect competitive positions. This reflects the Collingridge (1980) dilemma: the strategic implications and competitive externalities of transformative AI only become fully apparent once the technology is developed, at which point competitive dependencies preclude coordinated restraint. Early restraint fails from uncertainty about implications; late restraint fails from path dependency and lock-in to competitive dynamics. The externality is intrinsic to the technology’s transformative potential.
Break Condition 3 (Dominance Property):
Reduce winner-take-all dynamics and existential stakes making defection dominant. Mechanisms include: (1) credible international agreements reducing security competition between major powers, (2) economic restructuring through profit-sharing mechanisms such as the ‘Windfall Clause’ proposed by O’Keefe et al. (2020), where AI developers pre-commit to distributing exceptional profits (exceeding 1% of world GDP) broadly, and (3) demonstrated proof that advanced AI poses catastrophic risks even to prospective winners, undermining the perceived advantage of racing.
However, all three mechanisms face substantial barriers. International security agreements face severe commitment credibility problems: nations cannot reliably forego military advantages from transformative AI when adversaries might defect, creating overwhelming pressure to maintain competitive development. Economic restructuring through windfall clauses requires binding pre-commitment before competitive pressures emerge, making voluntary adoption difficult, and cannot address non-economic strategic advantages—military capability and geopolitical influence remain winner-take-all even with profit-sharing. Demonstrated proof of catastrophic risks faces an epistemic catch-22: definitive proof requires observing catastrophic outcomes, at which point it is too late. As Bostrom (2014) notes, the difficulty of credibly demonstrating novel existential risks before they materialize is a fundamental challenge, and even acknowledged risks may be insufficient when individual actors face dominant incentives to proceed if they believe others will do so regardless.
Break Condition 4 (Pareto-Inefficiency):
Accept the racing equilibrium as optimal, rejecting the premise that coordination would improve outcomes. Some argue that rapid AI deployment’s benefits outweigh coordination costs, effectively denying the Pareto-inefficiency condition.
However, this position faces severe challenges. As Bostrom (2014) and Russell (2019) document, the existential risks from misaligned superintelligence are non-excludable since even prospective winners cannot escape consequences of systems that exceed human control. The astronomical disvalue of existential catastrophe dominates even enormous competitive gains. This position requires either underestimating these non-excludable risks or assuming they are negligible, both of which the substantial AI safety research literature finds unconvincing. As discussed in Section 5.1, even prospective winners face negative expected utility from racing when probability of catastrophic outcomes is considered.
Break Condition 5 (Enforcement Difficulty):
Implement hardware-based verification using advanced semiconductor manufacturing as a chokepoint. Anderljung et al. (2023) argue that ‘compute governance’, e.g., monitoring and controlling access to high-end GPU clusters required for frontier AI training, offers the most viable enforcement mechanism, as specialized hardware is more detectable than software development. A global regime tracking chip production, data center construction, and computing cluster deployment could theoretically enforce development limits.
However, this advantage faces systematic erosion. As Sastry et al. (2024) document, algorithmic efficiency improvements systematically lower hardware requirements, allowing actors to achieve frontier capabilities with previously sub-frontier hardware—as demonstrated by DeepSeek’s early 2025 breakthrough achieving frontier results with dramatically reduced compute. Over multi-year timescales, hardware governance faces obsolescence as efficiency improvements enable powerful AI development on widely available hardware. Additionally, effective governance requires unprecedented international cooperation among competing powers, including adversarial nations with strong incentives to defect for strategic advantage. Verification of compliance faces challenges from concealed computing clusters and clandestine development programs.
Comprehensive Solution:
Resolving the AI governance tragedy requires interventions that simultaneously address multiple conditions: reducing the number of independent actors (C1), implementing credible verification mechanisms (C5), and reducing winner-take-all dynamics (C3). The most feasible approach likely combines:
  • International consortium among leading AI powers (partial C1 consolidation)
  • Compute governance with hardware tracking (partial C5 enforcement)
  • Windfall clauses and benefit-sharing mechanisms (partial C3 reduction)
  • Continued investment in technical AI safety (risk mitigation if governance fails)
However, our framework reveals the structural difficulty: each intervention faces severe implementation barriers, and partial solutions leave the tragedy substantially intact. The gap between theoretical solutions and practical feasibility measures the unprecedented nature of the coordination challenge. Unlike nuclear weapons, where physical constraints enabled partial success through non-proliferation regimes, AI’s software nature, dual-use characteristics, and low entry barriers create coordination requirements beyond historical precedent. The framework suggests we face a structurally harder problem than any successfully managed global coordination challenge to date.
5.6 Reconciliation with Conservative Productivity Estimates
Acemoglu (2025) provides conservative estimates of AI’s productivity impact (approximately 0.55% annual growth over ten years), attributing this limited potential to the prevalence of “so-so automation” that displaces labor without maximizing output. However, the structural tragedy operates independently of both magnitude and technological type. Even if AI offers only modest, displacement-heavy advantages, the dominance condition (C3) holds: firms cannot afford to forego even marginal efficiencies when competitors adopt them. The gap between exponential growth rates widens over time regardless of the base rate, meaning economic and military advantages compound. Conservative estimates still show AI confers sufficient advantages that actors cannot afford restraint while competitors advance.
Uncertainty amplifies this dynamic. Consistent with Prospect Theory (Kahneman and Tversky, 1979), actors facing the perceived certainty of strategic obsolescence will gamble on risky acceleration rather than accept certain disadvantage.
6 Implications and Falsifiability
Having established that both productivity competition and AI development satisfy the all necessary and sufficient conditions for structural tragedy, we now discuss implications, falsifiability, and the crucial distinction between structural pressure and metaphysical necessity.
6.1 What the Framework Predicts
Our framework generates strong predictions testable against empirical evidence:
For Productivity:
  1. 1.
    Working hours will not decline substantially in competitive market economies absent structural intervention
  2. 2.
    Productivity gains will primarily flow to output expansion rather than welfare improvements
  3. 3.
    Attempts to coordinate work reduction through voluntary agreements will fail
  4. 4.
    National regulations reducing hours will face competitive pressure and erosion
  5. 5.
    Cultural shifts favoring leisure will be insufficient absent structural change
For AI:
  1. 1.
    Voluntary pauses on capabilities development will fail to materialize or will collapse quickly
  2. 2.
    Safety commitments will erode under competitive pressure as capabilities approach advanced levels
  3. 3.
    Racing dynamics will intensify as capabilities approach human-level performance
  4. 4.
    National regulations alone will be insufficient and will face circumvention
  5. 5.
    International coordination attempts will face structural barriers absent enforcement mechanisms
Falsifiability:
The framework is falsified if we observe:
  • Working hours declining substantially despite continued competition
  • Sustained and verifiable voluntary pause on AI capabilities development lasting 12+ months with major actors participating
  • Verifiable global cap on compute or capabilities enforced without defection
  • Coordination achieving substantial progress on any of the five conditions without new enforcement mechanisms
  • Removal of any one of the five conditions through policy or institutional change
These predictions are specific enough to be empirically testable. The framework does not claim coordination is metaphysically impossible but only structurally difficult. Evidence of successful coordination that overcomes structural barriers would falsify our analysis.
6.2 Scope and Limitations
This framework identifies structural conditions that create coordination difficulties, but we do not claim to provide comprehensive causal explanations for all historical variation in productivity-welfare relationships or coordination outcomes across domains.
Temporal Variation:
The intensification of the tragedy after 1980 coincides with globalization, union decline, and increased capital mobility strengthening the conditions our framework identifies. However, detailed historical analysis is needed to establish definitive causal mechanisms for these temporal shifts.
Cross-national Variation:
Different institutional arrangements produce different manifestations of the productivity tragedy. Some countries exhibit the tragedy through stagnant wages despite productivity growth, others through stable hours but higher unemployment, and still others through pressure on benefits and job security. Our framework predicts that productivity gains will not flow fully to workers when competitive conditions are strong, but does not predict which specific dimension will bear the adjustment cost in each institutional context. Analysis of these institutional details and their interaction with competitive pressure is beyond our scope.
Condition Intensity:
As we have shown with the tragedy intensity extension, real-world conditions vary in intensity. When conditions hold weakly rather than strongly, partial coordination may be more feasible and sustainable. For instance, European labor market coordination succeeds in maintaining lower working hours than the United States, suggesting that weaker global competitive pressure (in non-tradable sectors) or stronger institutions can attenuate (though not eliminate) the tragedy. Future research should investigate systematically how condition intensity relates to the degree of coordination difficulty and the sustainability of partial coordination arrangements.
Empirical Validation:
We provide initial empirical evidence through declining labor share of income, productivity-wage decoupling, the European case for productivity, and the Russia-Ukraine drone war for AI development. However, comprehensive testing across historical periods, countries, sectors, and coordination attempts is necessary to fully validate our framework and identify boundary conditions. We view this paper as establishing a diagnostic framework and providing suggestive evidence, with systematic empirical validation remaining an important direction for future research.
7 Conclusion
7.1 Summary of Contributions
We have established four contributions:
First, we synthesized a unified diagnostic framework from established game theory (the social dilemma) by identifying five necessary and sufficient conditions. Second, we extended this framework to quantify tragedy severity through condition intensities, introducing a Tragedy Index that enables systematic comparison of coordination difficulty across domains, revealing that transformative AI breakthroughs face orders-of-magnitude greater coordination challenges than climate change, nuclear weapons, or bank runs. Third, we applied this framework to productivity, providing a formal game-theoretic model (building on (Schnaiberg, 1980)) that proves the coordination failure compelling firms to expand output is a structural tragedy, explaining Keynes’s failed prediction. The European case shows the tragedy is not escaped but merely redistributed at high cost. Fourth, we applied this framework to AI governance, demonstrating that transformative breakthroughs face the same structure but with amplified intensity. Our systematic comparison showed such governance is structurally more difficult than historical arms control. The drone war (on the way towards full autonomy) validates this dynamic.
7.2 Theoretical and Practical Significance
This structural framework enables several advances:
Prediction and Comparison: The framework enables systematic prediction of coordination difficulty and comparison across domains.
Intervention Design: Understanding which conditions create tragedy clarifies intervention targets. Only interventions targeting the five structural conditions have prospects for success.
Falsifiability: The framework generates testable predictions about which coordination attempts will succeed or fail based on structural conditions.
Clear Diagnosis: The framework explains why standard solutions fail and provides realistic assessment of coordination difficulty.
7.3 Future Research
This framework enables several research directions:
Identification of New Tragedies:
By checking the five conditions, can new coordination failures be identified in emerging domains? Do technologies like quantum computing, synthetic biology, or geoengineering exhibit the structural conditions that predict coordination difficulty? Could applying the framework to existing problems where solutions have failed reveal structural barriers previously overlooked, suggesting interventions that target the underlying conditions rather than symptoms?
Intensity Measurement:
What is the best way to quantify condition intensity? Is our Tragedy Index formulation a useful starting point, or would alternative aggregation methods better capture coordination difficulty? How does measured intensity correlate with observed governance outcomes across historical cases? Could systematic application of game-theoretic measures (e.g., Price of Anarchy) across domains validate the Index’s ordinal rankings and reveal whether structural and payoff-based approaches converge?
Formalization Refinement:
Our framework applies game-theoretic concepts largely heuristically. Could dynamic models with continuous strategies and empirically calibrated parameters confirm the tragedy persists under more realistic assumptions? Would computational simulations of multi-agent systems validate equilibrium predictions?
Productivity Model Extensions:
Our productivity formalization uses simplified utility functions and assumes comparable productivity gains. Could more realistic utility specifications incorporating risk aversion, time preferences, and worker-firm bargaining better capture the mechanisms? How does the tragedy structure change under persistent heterogeneous productivity shocks where market consolidation is incomplete or ongoing?
Cross-domain Dynamics:
How do tragedies in different domains interact? Does the productivity tragedy accelerate AI development by providing resources and competitive pressure for the capabilities race? Can success in coordinating one domain create spillovers that facilitate coordination in structurally similar domains?
Comparative Governance Analysis:
Would systematic comparison of governance attempts across domains reveal which structural conditions most resist coordination? Which interventions most effectively target specific conditions, and under what circumstances do partial solutions succeed or fail?
Historical Validation:
Could analysis of past coordination successes and failures through the framework reveal whether structural changes preceded successful coordination? Do cases like the Montreal Protocol, nuclear test bans, or labor regulations show condition-breaking interventions, or merely temporary overrides that eventually eroded?
Military AI and Conflict Escalation:
Does autonomous weapons development create a distinct tragedy where AI-enabled military systems lower the threshold for initiating conflict? If decision-making speed and reduced human casualties make war less costly for attackers, does this break the mutual restraint that deterrence theory assumes? Could the framework predict which military applications of AI face the most severe coordination barriers?
Institutional Resilience:
The European case shows partial coordination under favorable conditions, but with signs of erosion. What determines the sustainability of costly institutional overrides? Can we predict which coordination arrangements will persist versus collapse under competitive pressure?
Implications for Political Science and Economics:
Could the structural tragedy framework unify phenomena currently treated separately across disciplines? Would reframing problems like arms races, trade wars, regulatory competition, or financial contagion through the five-condition lens reveal common structural solutions? Can the framework inform institutional design in international relations and political economy?